### v1

# Stability and Decay for the 2D Anisotropic Navier-Stokes Equations with Fractional Horizontal Dissipation on $\mathbb{R}^{2}$

Abstract
The stability problem for the 2D Navier-Stokes equations with dissipation in only one direction on $\mathbb{R}^{2}$ is not fully understood. This dissipation is in the intermediate regime between the fully dissipative Navier-Stokes and the inviscid Euler. Navier-Stokes solutions in $\mathbb{R}^{2}$ decay algebraically in time while Euler solutions can grow rather rapidly in time. This paper solves the fundamental stability and large-time behavior problem on the anisotropic Navier-Stokes with fractional dissipation $\Lambda_{1}^{2s}$ for all $0\leq s<1$. The case $s=1$ corresponds to the standard one directional dissipation $\partial_{1}^{2}$. Different techniques are developed to treat different ranges of fractional exponents: $0\leq s\leq\frac{3}{4}$, $\frac{3}{4}<s<\frac{11}{12}$, and $\frac{11}{12}\leq s<1$. The final range is the most difficult case, for which we introduce the spatial polynomial $A_{2}$ weights and exploit the boundedness of Riesz transforms on weighted $L^{2}$-spaces.
Keywords: 2D anisotropic Navier-Stokes equations; Fractional horizontal dissipation; Stability; Decay estimate; $A_{p}$ weight.
Mathematics Subject Classification (2020): 35Q30; 35B35; 35B40; 76D05; 76E09.
1 Introduction
This paper studies the two-dimensional (2D) incompressible Navier-Stokes equations with fractional dissipation acting only in one spatial direction, given by
$\left\{\begin{aligned} &\partial_{t}u+\nu\Lambda_{1}^{2s}u+u\cdot\nabla u=-\nabla p,\qquad(x_{1},x_{2})\in\mathbb{R}^{2},\ t>0,\\ &\nabla\cdot u=0,\\ &u|_{t=0}=u_{0},\end{aligned}\right$. (1.1)
where $u(t,x)=(u_{1}(t,x_{1},x_{2}),u_{2}(t,x_{1},x_{2}))$ denotes the velocity field, $p(t,x)$ is the pressure, and $\nu>0$ is the viscosity coefficient. The operator $\Lambda_{1}^{2s}$ denotes the fractional derivative in the $x_{1}$-direction, defined via the Fourier transform by
$\widehat{\Lambda_{1}^{2s}f}(\xi)=|\xi_{1}|^{2s}\widehat{f}(\xi),\qquad 0\leq s\leq 1.$
When $s=1$, equation (1.1) reduces to the standard 2D Navier-Stokes equations with horizontal dissipation:
$\left\{\begin{aligned} &\partial_{t}u-\nu\partial_{1}^{\,2}u+u\cdot\nabla u=-\nabla p,\qquad(x_{1},x_{2})\in\mathbb{R}^{2},\ t>0,\\ &\nabla\cdot u=0.\end{aligned}\right$. (1.2)
Such anisotropic dissipative models arise naturally in fluid dynamics when there is a significant disparity between horizontal and vertical length scales. In geophysical flows, thin-layer fluids, and boundary layer regimes, the dominant dissipative mechanisms may be effectively confined to a preferred direction after appropriate rescaling. Classical examples of this phenomenon include reduced models derived from the Navier-Stokes equations under anisotropic scaling limits, such as the Prandtl-type systems [18, 19, 11]. Fractional dissipation in a single direction can also be viewed as an effective description of nonlocal mixing or anomalous diffusion along that direction, and thus has some meaningful physical interpretation.
From a mathematical perspective, anisotropic dissipation places the system in an intermediate regime between the fully dissipative Navier-Stokes equations and the inviscid Euler equations. For the standard 2D Navier-Stokes equations in the whole space $\mathbb{R}^{2}$ with full Laplacian dissipation, it is well known that solutions exhibit algebraic decay in time. Such decay can be derived using the Fourier splitting method introduced by Schonbek [20], which shows that the system is asymptotically stable and that solutions decay as time grows. In particular, in two dimensions one has the decay estimates
$\|u(\cdot,t)\|_{L^{q}}\leq Ct^{-1+1/q},\quad 2\leq q\leq\infty\quad\mbox{and}\quad\|\nabla u(\cdot,t)\|_{L^{q}}\leq Ct^{-3/2+1/q},\quad 2\leq q<\infty,$
as established in [20]. These estimates reflect the strong stabilizing effect of isotropic viscosity on the large-time dynamics of the flow.
In contrast, the 2D Euler equations, namely (1.1) with $\nu=0$, exhibit fundamentally different behavior. While Yudovich’s classical theory guarantees global existence and uniqueness of solutions with bounded vorticity, higher Sobolev norms of solutions typically grow in time. In particular, the gradient of the vorticity may experience substantial growth, and determining its precise growth rate has been a central problem in the analysis of inviscid flows. Seminal works by Sverák and Kiselev constructed smooth initial vorticity supported in a disk for which the vorticity gradient grows double exponentially in time [14]. In periodic domains, exponential-in-time growth of higher Sobolev norms has been established [6, 7, 8, 25]. In the whole space $\mathbb{R}^{2}$, linear-in-time growth was obtained by Choi and Jeong [5], as well as by Drivas, Elgindi, and Jeong [10], and more recently, superlinear growth examples were constructed by Jeong, Yao, and Zhou [12]. These results demonstrate that, despite global well-posedness, the Euler dynamics can exhibit strong instability at the level of higher regularity.
The model considered in this paper, with dissipation acting only in one direction, lies precisely between these two classical regimes. On the one hand, by Yudovich-type arguments, the system admits a unique global solution for sufficiently regular initial data. On the other hand, the dissipation is too weak to immediately imply decay properties comparable to those of the fully dissipative Navier-Stokes equations. This raises a fundamental question concerning stability and large-time behavior: does the solution behave more like Navier-Stokes, exhibiting decay and stabilization, or more like Euler, allowing growing high-frequency activity?
This question is not trivial. The absence of dissipation in one spatial direction creates substantial difficulties in closing energy estimates at the level of higher derivatives. Although dissipation in one direction is present, it fails to directly control all components of the nonlinear interactions. As a consequence, the best available a priori bounds suggest that certain Sobolev norms of solutions may grow exponentially in time.
To illustrate this difficulty, consider the vorticity formulation of (1.1),
$\partial_{t}\omega+u\cdot\nabla\omega+\Lambda_{1}^{2s}\omega=0,$
where the vorticity $\omega\overset{\mathrm{def}}{=}\nabla\times u$. It is well known that the $L^{1}\cap L^{\infty}$ norm of $\omega$ remains uniformly bounded. However, when one applies the energy method to estimate $\nabla\omega$, one obtains
$\frac{d}{dt}\|\nabla\omega(t)\|_{L^{2}}^{2}+2\nu\|\Lambda_{1}^{s}\nabla\omega(t)\|_{L^{2}}^{2}=$ $-2\int_{\mathbb{R}^{2}}\nabla u\cdot\nabla\omega\cdot\nabla\omega\,dx$
$=$ $-2\int_{\mathbb{R}^{2}}\partial_{1}u_{1}\,(\partial_{1}\omega)^{2}\,dx-2\int_{\mathbb{R}^{2}}\partial_{1}u_{2}\,\partial_{2}\omega\,\partial_{1}\omega\,dx$
$\quad-2\int_{\mathbb{R}^{2}}\partial_{2}u_{1}\,\partial_{1}\omega\,\partial_{2}\omega\,dx-2\int_{\mathbb{R}^{2}}\partial_{2}u_{2}\,(\partial_{2}\omega)^{2}\,dx.$ (1.3)
When $s=1$, the last two terms on the right-hand side of (1) do not admit time-integrable upper bounds. As a result, the dissipation is insufficient to prevent potential growth of $\|\nabla\omega(t)\|_{L^{2}}$. Understanding whether this apparent growth reflects a genuine instability or can be suppressed by more delicate anisotropic mechanisms is a central motivation of this paper.
It is also worth emphasizing that the stability problem here depends crucially on the spatial domain. When the spatial domain is $\mathbb{T}\times\mathbb{R}$, the work of Dong, Wu, Xu, and Zhu [9] established the stability of solutions to the 2D Navier-Stokes equations with dissipation acting only in one direction. Their analysis relies on decomposing the velocity field into its horizontal average and oscillatory components, and they showed that the $H^{1}$-norm of the oscillatory part decays exponentially in time. However, this approach fails in the whole-space setting $\mathbb{R}^{2}$. In contrast to the partially periodic case, one can no longer separate the zero horizontal Fourier mode from the oscillatory component and no Poincaré type inequality can be applied in the whole space $\mathbb{R}^{2}$.
This paper resolves the stability problem for (1.1) for the full range $0\leq s<1$. Different analytical techniques are developed for different ranges of the fractional dissipation exponent $s$. For $0\leq s\leq\tfrac{3}{4}$, we observe that although a larger value of $s$ strengthens the dissipation and is advantageous for global regularity, the situation is more subtle for stability problems. In this regime, a smaller value of $s$ actually reduces the number of derivatives required to establish the time integrability of nonlinear terms. As a consequence, a suitable redistribution of derivatives leads to improved time-integrability properties. This mechanism explains why weaker fractional dissipation can nevertheless be sufficient to establish stability in this range.
To extend stability beyond $s>\tfrac{3}{4}$, we combine stability estimates with large-time decay analysis. This naturally requires additional assumptions on the initial perturbation, specifically that it belongs to a negative-order Sobolev space in the horizontal direction, namely
$\Lambda_{1}^{-\sigma}u_{0}\in L^{2},\qquad 0<\sigma<\tfrac{1}{2}.$
The restriction $\sigma<\tfrac{1}{2}$ is essential in order to propagate the negative regularity along the flow. By coupling decay estimates with stability arguments, we are able to establish stability of (1.1) for
$\tfrac{3}{4}<s<\tfrac{11}{12}.$
Finally, to treat the remaining regime $\tfrac{11}{12}\leq s<1$ (in fact, this approach also applies to all $s>\tfrac{3}{4}$), we introduce a new strategy based on weighted energy estimates. More precisely, we employ a weight function in the vertical variable $x_{2}$ that grows polynomially of order $\gamma$ outside the unit disk and is truncated to $1$ inside it. The analysis in this regime makes essential use of the theory of $A_{p}$ weights and the boundedness of Riesz transforms on the corresponding weighted $L^{2}$ spaces. These tools allow us to control nonlocal terms arising from the Biot–Savart law in the presence of spatial weights, which is crucial for closing the weighted energy estimates. Under the assumption that the initial data and its derivatives exhibit suitable decay in the $x_{2}$-direction, we prove that solutions to (1.1) are stable and enjoy optimal decay rates in a neighborhood of the trivial solution.
We now summarize our main results, organized according to different ranges of the fractional dissipation exponent $s$. We first consider the range $0\leq s\leq\tfrac{3}{4}$. In this regime, small initial data in $H^{k}(\mathbb{R}^{2})$ with $k\geq 3$ is sufficient to guarantee global stability.
Theorem 1.1.
Consider the anisotropic Navier-Stokes system (1.1) with viscosity coefficient $\nu>0$. Let $0\leq s\leq\tfrac{3}{4}$ and assume the initial data $u_{0}\in H^{k}(\mathbb{R}^{2})$, $k\geq 3$, satisfies $\nabla\cdot u_{0}=0$. Then there exists a sufficiently small constant $\varepsilon>0$ such that if $\|u_{0}\|_{H^{k}}\leq\varepsilon$, the system admits a unique global solution
$u\in C([0,\infty);H^{k}(\mathbb{R}^{2})),$
which satisfies the uniform bound
$\|u(t)\|_{H^{k}}+\nu^{\frac{1}{2}}\left(\int_{0}^{t}\|\Lambda_{1}^{s}u(\tau)\|_{H^{k}}^{2}\,d\tau\right)^{\frac{1}{2}}\leq C\varepsilon,\qquad\forall\,t\geq 0.$
When the dissipation exponent satisfies $s>\tfrac{3}{4}$, energy estimates alone are no longer sufficient to control the nonlinear terms. In this regime, stability must be coupled with decay estimates, which in turn requires additional assumptions on the initial perturbation in the horizontal direction. The following theorem establishes the nonlinear stability and large-time decay of solutions for dissipation exponents in the range
$\tfrac{3}{4}<s<\tfrac{11}{12}.$
Theorem 1.2.
Let
$\tfrac{3}{4}<s<\tfrac{5}{12}+\sigma,\qquad\tfrac{1}{3}<\sigma<\tfrac{1}{2},$
and assume the initial data $u_{0}\in H^{k}(\mathbb{R}^{2})$ with
$k>\max\left\{\frac{2\sigma}{1-2\sigma}+1,\;9\right\}$
is divergence-free. Suppose that $u_{0}$ admits a stream function $\psi_{0}$ such that
$u_{0}=(u_{0,1},u_{0,2})=\nabla^{\perp}\psi_{0}:=(-\partial_{2}\psi_{0},\partial_{1}\psi_{0}),$
and that there exists a sufficiently small constant $\varepsilon>0$ such that
$\|u_{0}\|_{H^{k}}+\|\Lambda_{1}^{-\sigma}\psi_{0}\|_{L^{2}}+\|\Lambda_{1}^{-\sigma}u_{0,1}\|_{L^{2}}+\|\Lambda_{1}^{-\sigma}\partial_{2}u_{0,1}\|_{L^{2}}\leq\varepsilon.$ (1.4)
Then the system admits a unique global-in-time solution
$u\in C([0,\infty);H^{k}(\mathbb{R}^{2})),$
satisfying
$\|\Lambda_{1}^{-\sigma}u(t)\|_{H^{k}}+\nu^{\frac{1}{2}}\left(\int_{0}^{t}\|\Lambda_{1}^{s-\sigma}u(\tau)\|_{H^{k}}^{2}\,d\tau\right)^{\frac{1}{2}}\leq C\varepsilon,\qquad\forall\,t\geq 0,$
together with the decay estimates
$\|u_{1}(t)\|_{L^{2}}$ $\leq C_{0}\varepsilon(1+t)^{-\sigma/2s},$
$\|u_{2}(t)\|_{L^{2}}$ $\leq C_{0}\varepsilon(1+t)^{-(\sigma+\frac{2}{3})/2s},$
$\|\partial_{2}u_{1}(t)\|_{L^{2}}$ $\leq C_{0}\varepsilon(1+t)^{-\sigma/2s},$
$\|\partial_{1}u_{1}(t)\|_{L^{2}}$ $\leq C_{0}\varepsilon(1+t)^{-(\sigma+1)/2s},$
$\|\partial_{1}u_{2}(t)\|_{L^{2}}$ $\leq C_{0}\varepsilon(1+t)^{-2\sigma/s},$
where $C_{0}>0$ is a universal constant.
A special feature of this result is that the second velocity component $u_{2}$ decays faster than the first component $u_{1}$. This anisotropic decay is revealed through the use of the stream function and is a key ingredient in the proof of the theorem. The enhanced decay of $u_{2}$ plays a crucial role in closing the nonlinear estimates. Moreover, since the Navier-Stokes equations do not propagate negative Sobolev regularity of the form $\Lambda^{-\sigma}u$ for $\sigma\geq\tfrac{1}{2}$, the restriction $\sigma<\tfrac{1}{2}$ is necessary.
Remark 1.3.
We further remark that, even for fractional dissipation in the range $0<s<\tfrac{3}{4}$, under assumptions analogous to those of Theorem 1.2, we can also establish the following decay estimates:
$\|u_{1}(t)\|_{L^{2}}$ $\leq C_{0}\varepsilon(1+t)^{-\sigma/2s},$
$\|u_{2}(t)\|_{L^{2}}$ $\leq C_{0}\varepsilon(1+t)^{-(4\sigma+1)/4s},$ (1.5)
$\|\partial_{2}u_{1}(t)\|_{L^{2}}$ $\leq C_{0}\varepsilon(1+t)^{-\sigma/2s},$
$\|\partial_{1}u_{1}(t)\|_{L^{2}}$ $\leq C_{0}\varepsilon(1+t)^{-(\sigma+1)/2s},$
$\|\partial_{1}u_{2}(t)\|_{L^{2}}$ $\leq C_{0}\varepsilon(1+t)^{-(2\sigma+1)/2s}.$
The derivation of the above estimates parallels that of Theorem 1.2, we only present the estimates for $u_{2}$, which differs slightly, in Appendix B.
Once again, these estimates shows that the second velocity component $u_{2}$ exhibits a faster decay rate than $u_{1}$, reflecting the underlying anisotropic structure of the system.
The decay rate stated in Theorem 1.2 is not sufficiently fast to treat the range $s>\tfrac{11}{12}$. This limitation is partly due to the appearance of the Riesz transform in the representation of $u_{2}$, which prevents obtaining faster decay for this component. This difficulty motivates us to introduce spatial weights and develop a weighted energy approach, which allows us to extend the global existence and regularity results to the full range $0<s<1$. We note that Riesz transforms are bounded on weighted $L^{2}$ spaces associated with $A_{2}$ weights, which makes this approach possible.
More precisely, we introduce a weight function for $x_{2}$ and solve the problem for $0<s<1$. This type of weight function grows at a polynomial rate of order $\gamma$ outside the unit disk and is truncated to 1 inside it. We prove that if the initial data and its derivative decay in the $x_{2}$-direction with some rate, then the solution of (1.1) is stable and have an optimal decay rate near the trivial solution.
Definition 1.4 (Truncated Power Weight).
We define $x]^{\gamma}$ on $\mathbb{R}$ as:
$[x]^{\gamma}:=\begin{cases}1,&\ |x|\leq 1,\\ |x|^{\gamma},&\ |x|>1.\end{cases}$
Theorem 1.5.
Let the parameters $s,\sigma$ and $\gamma$ satisfy
$\frac{3}{4}<s<\frac{1}{2}+\sigma,\qquad\frac{1}{3}<\sigma<\frac{1}{2},\qquad 0<\gamma<\frac{3}{10}.$
Assume that the initial data $u_{0}\in H^{k}(\mathbb{R}^{2})$ with
$k>\frac{2\sigma+2}{1-2\sigma}+1$
is divergence-free and admits a stream function $\psi_{0}$ such that $u_{0}=(u_{0,1},u_{0,2})=\nabla^{\perp}\psi_{0}$. Suppose furthermore that there exists a small constant $\varepsilon>0$ such that
$\|u_{0}\|_{H^{k}}+\|\Lambda_{1}^{-\sigma}[x_{2}]^{\gamma}\psi_{0}\|_{L^{2}}+\|\Lambda_{1}^{-\sigma}[x_{2}]^{\frac{3\gamma+4}{7}}u_{0}\|_{L^{2}}+\|\Lambda_{1}^{-\sigma}[x_{2}]^{\frac{5\gamma+2}{7}}\partial_{2}u_{0,1}\|_{L^{2}}$
$\quad+\|[x_{2}]^{\frac{3\gamma+4}{7}}u_{0}\|_{L^{2}}+\|[x_{2}]^{\frac{5\gamma+2}{7}}\partial_{2}u_{0,1}\|_{L^{2}}+\|[x_{2}]^{\frac{3\gamma+4}{7}}\partial_{1}u_{0}\|_{L^{2}}\leq\varepsilon.$ (1.6)
Then the system admits a unique global-in-time solution
$u\in C([0,\infty);H^{k}(\mathbb{R}^{2})),$
satisfying
$\|\Lambda_{1}^{-\sigma}u(t)\|_{H^{k}}+\nu^{1/2}\left(\int_{0}^{t}\|\Lambda_{1}^{s-\sigma}u(\tau)\|_{H^{k}}^{2}\,d\tau\right)^{1/2}<\varepsilon,\qquad\forall\,t\geq 0.$
Moreover, the solution enjoys the following weighted decay estimates:
$\|[x_{2}]^{\frac{3\gamma+4}{7}}u(t)\|_{L^{2}}$ $\leq C_{0}\varepsilon(1+t)^{-\sigma/2s},$ (1.7)
$\|[x_{2}]^{\frac{3\gamma+4}{7}}\partial_{1}u(t)\|_{L^{2}}$ $\leq C_{0}\varepsilon(1+t)^{-(\sigma+1)/2s},$ (1.8)
$\|[x_{2}]^{\frac{5\gamma+2}{7}}\partial_{2}u_{1}(t)\|_{L^{2}}$ $\leq C_{0}\varepsilon(1+t)^{-\sigma/2s}.$ (1.9)
In addition, we have a stronger weighted decay estimate for $u_{2}$:
$\|[x_{2}]^{\gamma}u_{2}(t)\|_{L^{2}}$ $\leq C_{0}\varepsilon(1+t)^{-(\sigma+1)/2s},$ (1.10)
$\|[x_{2}]^{\gamma}\partial_{1}u_{2}(t)\|_{L^{2}}$ $\leq C_{0}\varepsilon(1+t)^{-(2\sigma+1)/2s}.$ (1.11)
where $C_{0}>0$ is a universal constant.
By introducing polynomial weights in the vertical variable $x_{2}$, we are able to compensate for this loss and recover additional decay for $u_{2}$. The specific fractional exponents in the weights are carefully chosen to close the nonlinear estimates, and the restriction $\gamma>0$ ensures the integrability of weighted nonlinear interactions. The boundedness of Riesz transforms on weighted $L^{2}$ spaces with $A_{2}$ weights guarantees that the weighted energy estimates close and allows the extension of the global existence and decay theory to the full range $0<s<1$. It is noteworthy that for some $\gamma$, the weight of $u_{1}$ and its derivatives are not an $A_{2}$ weight, but we can still prove that it allows us to eliminate the Riesz transform in the pressure $p$, thereby establishing the boundedness of the Leray projection $\mathbb{P}$ under this weighted norm. The details will be addressed in Section 5.
Remark 1.6.
Indeed, the upper bound $\gamma<\frac{3}{10}$ is not necessary due to the monotone property of the weight function (see Lemma 2.9).
Remark 1.7.
The decay rates in (1.7)–(1.11) are optimal.
Remark 1.8.
The decay estimate (1.7)–(1.11) are also true without weight.
Notations. In the $d$-dimensional Euclidean space $\mathbb{R}^{d}$, $L^{p}$ denotes the standard Lebesgue space. Given two normed spaces $N_{1}$ and $N_{2}$ on measure spaces $(X_{1},\mu_{1})$ and $(X_{2},\mu_{2})$, respectively, the anisotropic space $N_{1}(X_{1},\mu_{1})\,N_{2}(X_{2},\mu_{2})$ is defined by first taking the $N_{1}$-norm in the $X_{1}$ variable and then the $N_{2}$-norm in the $X_{2}$ variable. More precisely, for a function $f$ defined on the product space $X_{1}\times X_{2}$, we set
$\|f\|_{N_{1}(X_{1},\mu_{1})\,N_{2}(X_{2},\mu_{2})}:=\bigl\|\,\|f(x_{1},x_{2})\|_{N_{1}(X_{1},\mu_{1})}\,\bigr\|_{N_{2}(X_{2},\mu_{2})}.$
The Sobolev space $H^{k}$ consists of functions whose weak derivatives up to order $k$ belong to $L^{2}$, while the homogeneous Sobolev space $\dot{H}^{k}$ is defined by requiring only the highest-order weak derivatives to belong to $L^{2}$.
The paper is organized as follows. Section 2 collects preliminary lemmas that are used throughout the paper. In Section 3, we prove Theorem 1.1, which establishes stability for $0\leq s\leq\tfrac{3}{4}$. Section 4 is devoted to the proof of Theorem 1.2, covering the range $\tfrac{3}{4}<s<\tfrac{11}{12}$, where the analysis requires a delicate coupling with decay estimates. In Section 5, we prove Theorem 1.5, which extends the stability result to the full range $\tfrac{3}{4}<s<1$ and establishes optimal decay estimates for $u$ in weighted spaces. Finally, the appendix contains the proofs of several technical lemmas, as well as detailed estimates used in Remark 1.3.
2 Anisotropic Estimates and Tools
This section gathers several analytic tools needed for the stability and decay analysis. These include anisotropic product estimates, Sobolev inequalities, decay estimates for the fractional heat semigroup, and weighted inequalities associated with truncated power weights. We also recall the $A_{p}$-weight theory and present the boundedness of Riesz transforms in weighted $L^{2}$ spaces, which plays a key role in the weighted energy estimates developed later.
Due to the anisotropic nature of the system, we first introduce the following two anisotropic product estimates to handle the nonlinear terms. Estimates of this type have been widely used in the study of regularity and stability problems for anisotropic partial differential equations; see, for example, [3, 4, 15, 23, 24].
Lemma 2.1.
Let $f,g\in H^{1}(\mathbb{R}^{2})$. Then there exists a constant $C>0$ such that
$\|fg\|_{L_{x_{1}}^{1}L_{x_{2}}^{2}}\leq C\,\|f\|_{L^{2}}^{1/2}\|\partial_{2}f\|_{L^{2}}^{1/2}\,\|g\|_{L^{2}}.$
Proof.
By Hölder’s inequality in the $x_{1}$ variable, we obtain
$\|fg\|_{L_{x_{1}}^{1}L_{x_{2}}^{2}}=\bigl\|\|fg\|_{L_{x_{1}}^{1}}\bigr\|_{L_{x_{2}}^{2}}\leq\bigl\|\|f\|_{L_{x_{1}}^{2}}\|g\|_{L_{x_{1}}^{2}}\bigr\|_{L_{x_{2}}^{2}}\leq\bigl\|\|f\|_{L_{x_{1}}^{2}}\bigr\|_{L_{x_{2}}^{\infty}}\|g\|_{L^{2}}.$
Applying Minkowski’s inequality and the one-dimensional Gagliardo–Nirenberg inequality in the $x_{2}$ variable yields
$\bigl\|\|f\|_{L_{x_{1}}^{2}}\bigr\|_{L_{x_{2}}^{\infty}}\leq C\,\|f\|_{L^{2}}^{1/2}\|\partial_{2}f\|_{L^{2}}^{1/2}.$
Combining the above estimates gives the desired inequality. ∎
Lemma 2.2.
Let $f,g\in H^{1}(\mathbb{R}^{2})$. Then there exists a constant $C>0$ such that
$\|fg\|_{L^{2}}\leq C\|f\|_{L^{2}}^{1/2}\|\partial_{1}f\|_{L^{2}}^{1/2}\|g\|_{L^{2}}^{1/2}\|\partial_{2}g\|_{L^{2}}^{1/2}.$
Proof.
By Hölder’s inequality, we obtain
$\|fg\|_{L^{2}}\leq C\,\|f\|_{L_{x_{1}}^{\infty}L_{x_{2}}^{2}}\,\|g\|_{L_{x_{1}}^{2}L_{x_{2}}^{\infty}}.$
Using Minkowski’s inequality together with the one-dimensional Gagliardo–Nirenberg inequality in the $x_{1}$ and $x_{2}$ variables, respectively, we have
$\|f\|_{L_{x_{1}}^{\infty}L_{x_{2}}^{2}}\leq C\,\|f\|_{L^{2}}^{1/2}\|\partial_{1}f\|_{L^{2}}^{1/2},\qquad\|g\|_{L_{x_{1}}^{2}L_{x_{2}}^{\infty}}\leq C\,\|g\|_{L^{2}}^{1/2}\|\partial_{2}g\|_{L^{2}}^{1/2}.$
Combining these estimates yields the desired inequality. ∎
The following lemma gives a product law in homogeneous Sobolev spaces. Its proof can be found in [1].
Lemma 2.3.
Let $f$ and $g$ be smooth functions. Assume that $s_{1},s_{2}<\frac{d}{2}$, $s_{1}+s_{2}>0$, and
$s+\frac{d}{2}=s_{1}+s_{2}.$
Then there exists a constant $C>0$, depending only on $d$, $s_{1}$, and $s_{2}$, such that
$\|\Lambda^{s}(fg)\|_{L^{2}}\leq C\,\|\Lambda^{s_{1}}f\|_{L^{2}}\,\|\Lambda^{s_{2}}g\|_{L^{2}}.$
The next lemma provides a related product estimate that will be useful in our analysis. Its proof is deferred to Appendix A.
Lemma 2.4.
Let $m>0$. There exists a constant $C=C(m)>0$ such that for all $f,g\in H^{m}(\mathbb{R}^{d})$,
$\|\Lambda^{m}(fg)\|_{L^{2}}\leq C\|(\Lambda^{m}f)g\|_{L^{2}}+C\|f(\Lambda^{m}g)\|_{L^{2}}.$ (2.1)
More generally, for $m_{1}\in\mathbb{R}$ and $m_{2}>0$, one has
$\|\Lambda^{m_{1}}\Lambda^{m_{2}}(fg)\|_{L^{2}}\leq C\left\|\Lambda^{m_{1}}\big[(\Lambda^{m_{2}}f)g\big]\right\|_{L^{2}}+C\left\|\Lambda^{m_{1}}\big[f(\Lambda^{m_{2}}g)\big]\right\|_{L^{2}}.$ (2.2)
We also recall the classical Hardy-Littlewood-Sobolev inequality, the proof can be found in [2].
Lemma 2.5 (Hardy-Littlewood-Sobolev inequality).
There holds
$\|\Lambda^{-\rho}f\|_{L^{q}(\mathbb{R}^{d})}\leq C\|f\|_{L^{p}(\mathbb{R}^{d})},$
with
$\frac{1}{q}+\frac{\rho}{d}=\frac{1}{p},\quad 0<\rho<d,\quad 1<p<q<\infty.$
To estimate the quadratic term involving a negative derivative, we introduce the following anisotropic product estimate. The idea of the proof is inspired by [13], but the argument is adapted and refined to suit the present setting.
Lemma 2.6.
Let $f,g$ be two smooth functions on $\mathbb{R}^{2}$ and let $\frac{1}{4}<\sigma<\frac{1}{2}$. Then
$\|\Lambda_{1}^{-\sigma}(fg)\|_{L^{2}}\leq C\|\partial_{2}f\|_{L^{2}}^{\frac{1}{2}}\|f\|_{L^{2}}^{\sigma}\|\partial_{1}f\|_{L^{2}}^{\frac{1}{2}-\sigma}\|g\|_{L^{2}}.$
Proof.
The proof relies on anisotropic Sobolev embeddings and interpolation estimates and is specific to the mixed regularity structure of the problem. By Lemma 2.5, we have
$\left\|\Lambda_{1}^{-\sigma}(fg)\right\|_{L^{2}}=\left\|\left\|\Lambda_{1}^{-\sigma}(fg)\right\|_{L_{x_{1}}^{2}}\right\|_{L_{x_{2}}^{2}}\leq\left\|\left\|fg\right\|_{L_{x_{1}}^{q}}\right\|_{L_{x_{2}}^{2}},$ (2.3)
where $\frac{1}{q}=\frac{1}{2}+\sigma$. Since $q<2$, by the Minkowski and Hölder inequalities, we obtain
$\left\|\left\|fg\right\|_{L_{x_{1}}^{q}}\right\|_{L_{x_{2}}^{2}}$ $\leq\left\|\left\|fg\right\|_{L_{x_{2}}^{2}}\right\|_{L_{x_{1}}^{q}}\leq\left\|\left\|f\right\|_{L_{x_{2}}^{\infty}}\left\|g\right\|_{L_{x_{2}}^{2}}\right\|_{L_{x_{1}}^{q}}\leq\left\|f\right\|_{L_{x_{1}}^{\frac{1}{\sigma}}L_{x_{2}}^{\infty}}\left\|g\right\|_{L^{2}}.$ (2.4)
Using Minkowski’s inequality and Gagliardo-Nirenberg’s inequality in $x_{2}$-direction, we have
$\left\|f\right\|_{L_{x_{1}}^{\frac{1}{\sigma}}L_{x_{2}}^{\infty}}\leq C\left\|\left\|f\right\|_{L_{x_{2}}^{2}}^{\frac{1}{2}}\left\|\partial_{2}f\right\|_{L_{x_{2}}^{2}}^{\frac{1}{2}}\right\|_{L_{x_{1}}^{\frac{1}{\sigma}}}.$
Applying Hölder’s inequality in $x_{1}$-variable, we have
$\left\|\left\|f\right\|_{L_{x_{2}}^{2}}^{\frac{1}{2}}\left\|\partial_{2}f\right\|_{L_{x_{2}}^{2}}^{\frac{1}{2}}\right\|_{L_{x_{1}}^{\frac{1}{\sigma}}}\leq C\left\|\left\|f\right\|_{L_{x_{2}}^{2}}^{\frac{1}{2}}\right\|_{L_{x_{1}}^{\frac{4}{4\sigma-1}}}\left\|\left\|\partial_{2}f\right\|_{L_{x_{2}}^{2}}^{\frac{1}{2}}\right\|_{L_{x_{1}}^{4}}=C\left\|f\right\|_{L_{x_{2}}^{2}L_{x_{1}}^{\frac{2}{4\sigma-1}}}^{\frac{1}{2}}\left\|\partial_{2}f\right\|_{L^{2}}^{\frac{1}{2}}.$
Using Minkowski’s inequality and the Sobolev embedding in $x_{1}$-direction, we deduce
$\left\|f\right\|_{L_{x_{2}}^{2}L_{x_{1}}^{\frac{2}{4\sigma-1}}}^{\frac{1}{2}}\left\|\partial_{2}f\right\|_{L^{2}}^{\frac{1}{2}}\leq\left\|f\right\|_{L_{x_{1}}^{\frac{2}{4\sigma-1}}L_{x_{2}}^{2}}^{\frac{1}{2}}\left\|\partial_{2}f\right\|_{L^{2}}^{\frac{1}{2}}\leq C\left\|\Lambda_{1}^{1-2\sigma}f\right\|_{L^{2}}^{\frac{1}{2}}\left\|\partial_{2}f\right\|_{L^{2}}^{\frac{1}{2}}.$
By interpolation, we have
$\left\|\Lambda_{1}^{1-2\sigma}f\right\|_{L^{2}}^{\frac{1}{2}}\left\|\partial_{2}f\right\|_{L^{2}}^{\frac{1}{2}}$ $\leq C\left\|\partial_{2}f\right\|_{L^{2}}^{\frac{1}{2}}\left(\left\|f\right\|_{L^{2}}^{2\sigma}\left\|\partial_{1}f\right\|_{L^{2}}^{1-2\sigma}\right)^{\frac{1}{2}}$
$\leq C\left\|\partial_{2}f\right\|_{L^{2}}^{\frac{1}{2}}\left\|f\right\|_{L^{2}}^{\sigma}\left\|\partial_{1}f\right\|_{L^{2}}^{\frac{1}{2}-\sigma}.$
Thus we conclude
$\left\|f\right\|_{L_{x_{1}}^{\frac{1}{\sigma}}L_{x_{2}}^{\infty}}\leq C\left\|\partial_{2}f\right\|_{L^{2}}^{\frac{1}{2}}\left\|f\right\|_{L^{2}}^{\sigma}\left\|\partial_{1}f\right\|_{L^{2}}^{\frac{1}{2}-\sigma}.$ (2.5)
Combining estimates (2.3), (2.4), and (2.5), we obtain
$\|\Lambda_{1}^{-\sigma}(fg)\|_{L^{2}}\leq C\|\partial_{2}f\|_{L^{2}}^{\frac{1}{2}}\|f\|_{L^{2}}^{\sigma}\|\partial_{1}f\|_{L^{2}}^{\frac{1}{2}-\sigma}\|g\|_{L^{2}},$
which completes the proof. ∎
The fractional heat semigroup estimate is frequently used in the following text, the proof of which can be found in [16].
Lemma 2.7.
Let $\sigma\geq 0$, $\alpha>0$, $\nu>0$, $1\leq p\leq q\leq\infty$. Then
$\|\Lambda^{\sigma}e^{-\nu(-\Delta)^{\alpha}t}f\|_{L^{q}\left(\mathbb{R}^{d}\right)}\leq Ct^{-\frac{\sigma}{2\alpha}-\frac{d}{2\alpha}\left(\frac{1}{p}-\frac{1}{q}\right)}\|f\|_{L^{p}\left(\mathbb{R}^{d}\right)}.$
The following lemma provides a decay estimate for a convolution-type integral. Its proof is given in Appendix A.
Lemma 2.8.
Assume $\alpha\geq 1$, $\beta>1$. Then, for some constant $C>0$ and $t>1$, we have
[…]
Assume $\alpha<1$. Then, for some constant $C>0$, we have
$\int_{0}^{t}(t-\tau)^{-\alpha}(1+\tau)^{-\beta}\>d\tau\leq\begin{cases}C(1+t)^{-\alpha},\ &\beta>1,\\ C(1+t)^{-\alpha}\ln(1+t),\ &\beta=1,\\ C(1+t)^{1-\alpha-\beta},\ &\beta<1.\end{cases}$ (2.7)
The following lemmas describe several basic properties of the truncated power weight function. The first lemma shows that the weighted $L^{2}$-norm is monotone with respect to the power of the weight.
Lemma 2.9 (Monotonicity of weights).
For any $\gamma_{1}\leq\gamma_{2}$, one has
$\|[x]^{\gamma_{1}}f\|_{L^{2}}\leq\|[x]^{\gamma_{2}}f\|_{L^{2}}.$
The second lemma provides a Poincaré-Friedrichs type inequality involving truncated power weights.
Lemma 2.10.
Let $0<\gamma\leq 1$. Then the following weighted Poincaré-Friedrichs inequality holds:
$\|[x]^{\gamma}f\|_{L^{2}(\mathbb{R})}\leq C\|[x]^{\gamma+1}f^{\prime}\|_{L^{2}(\mathbb{R})}.$
The third lemma establishes a weighted Gagliardo-Nirenberg type inequality.
Lemma 2.11.
Let $\zeta>0$ and $\vartheta\in\mathbb{R}$. Then
$\|[x_{2}]^{\zeta}f\|_{L^{\infty}(\mathbb{R})}\leq C\|[x_{2}]^{\zeta-\frac{1}{2}}f\|_{L^{2}(\mathbb{R})}+C\|[x_{2}]^{\zeta-\vartheta}f\|_{L^{2}(\mathbb{R})}^{\frac{1}{2}}\|[x_{2}]^{\zeta+\vartheta}f^{\prime}\|_{L^{2}(\mathbb{R})}^{\frac{1}{2}}.$
The proof of Lemma 2.9 is straightforward. The proofs of Lemmas 2.10 and 2.11 are deferred to Appendix A.
To handle the Riesz operators in weighted $L^{2}$ spaces, we make use of the $A_{p}$ weight theory. We first recall the definition of $A_{p}$ weights.
Definition 2.12 ($A_{p}$ weights).
A weight function $d\colon\mathbb{R}^{d}\to[0,\infty)$ is said to belong to the $A_{p}$ class if:
  1. 1.
    $d\in L_{\mathrm{loc}}^{1}(\mathbb{R}^{d})$ and $d(x)>0$ almost everywhere;
  2. 2.
    there exists a constant $C>0$ such that for every cube $Q\subset\mathbb{R}^{d}$, the $A_{p}$ condition
    $\left(\frac{1}{|Q|}\int_{Q}d(y)\,dy\right)\left(\frac{1}{|Q|}\int_{Q}d(y)^{-\frac{1}{p-1}}\,dy\right)^{p-1}\leq C,\qquad 1<p<\infty,$
    holds, where $C$ is independent of $Q$.
The following lemma gives the boundedness of Riesz operators in weighted $L^{2}$ spaces. Its proof is deferred to Appendix A.
Lemma 2.13.
Let $\mathcal{R}_{1}=\partial_{1}\Lambda^{-1}$ and $\mathcal{R}_{2}=\partial_{2}\Lambda^{-1}$ be the Riesz transforms. Let $T$ be any one of the operators $\mathcal{R}_{1}^{2}$, $\mathcal{R}_{1}\mathcal{R}_{2}$, or $\mathcal{R}_{2}^{2}$. Then, for any $1<\kappa<1$, there exists a constant $C>0$ such that
$\int_{\mathbb{R}^{2}}|Tf(x)|^{2}\,[x_{2}]^{\kappa}\,dx\leq C\int_{\mathbb{R}^{2}}|f(x)|^{2}\,[x_{2}]^{\kappa}\,dx.$
3 Proof of Theorem 1.1
This section is devoted to the proof of Theorem 1.1. We first note that the 2D Navier-Stokes equations with fractional horizontal dissipation are locally well posed in the Sobolev space $H^{k}(\mathbb{R}^{2})$ for any integer $k\geq 3$. The local existence and uniqueness can be established by standard arguments similar to those used for the classical Navier-Stokes or Euler equations; see, for example, [17]. We therefore omit the details of the local theory.
Our main effort in this section is to derive global-in-time a priori bounds, which allow the local solutions to be extended globally. These bounds are obtained through careful energy estimates. A crucial ingredient is a suitable upper bound on the nonlinear term, which compensates for the lack of vertical dissipation. The key estimate is summarized in the following lemma. The proof of this lemma is nontrivial and is based on a case-by-case analysis in the ranges $0\leq s\leq\tfrac{1}{4}$, $\tfrac{1}{4}\leq s\leq\tfrac{1}{2}$, and $\tfrac{1}{2}\leq s\leq\tfrac{3}{4}$. The proof is given after the proof of Theorem 1.1.
Lemma 3.1.
Let $k\geq 3$ be an integer, as in Theorem 1.1. For any integer $1\leq\beta\leq k$, $\ell\in\{1,2\}$, and $0\leq s\leq\frac{3}{4}$, the following estimate holds:
$\sum_{\beta=1}^{k}\int_{\mathbb{R}^{2}}\partial_{\ell}^{\beta}f\,\partial_{1}\partial_{\ell}^{k-\beta}f\,\partial_{\ell}^{k}f\,dx\leq C_{0}\|f\|_{H^{k}}\,\|\Lambda_{1}^{\,s}f\|_{H^{k}}^{2},$ (3.1)
where $C_{0}>0$ is a constant independent of $f$.
3.1 Proof of Theorem 1.1
Proof of Theorem 1.1:.
The proof is devoted to deriving global-in-time a priori bounds on solutions of (1.1). By virtue of the norm equivalence
$\|f\|_{H^{k}}^{2}\sim\|f\|_{L^{2}}^{2}+\sum_{m=1}^{2}\|\partial_{m}^{k}f\|_{L^{2}}^{2},$
it suffices to derive energy estimates for the $L^{2}$ norm and the highest-order derivatives. By a standard energy argument, we obtain the $L^{2}$ energy estimate for $u$,
$\|u(t)\|_{L^{2}}^{2}+2\nu\int_{0}^{t}\|\Lambda_{1}^{s}u(\tau)\|_{L^{2}}^{2}\,d\tau=\|u_{0}\|_{L^{2}}^{2}.$ (3.2)
We now turn to the estimate for the highest-order derivatives. Applying $\partial_{m}^{k}$ to the equation (1.1) and taking the $L^{2}$ inner product with $\partial_{m}^{k}u$ yields
$\sum_{m=1}^{2}\frac{d}{dt}\left\|\partial_{m}^{k}u\right\|_{L^{2}}^{2}+2\nu\sum_{m=1}^{2}\left\|\Lambda_{1}^{s}\partial_{m}^{k}u\right\|_{L^{2}}^{2}=-2\sum_{m=1}^{2}\big(\partial_{m}^{k}(u\cdot\nabla u),\partial_{m}^{k}u\big)_{L^{2}}\overset{\text{def}}{=}N.$ (3.3)
To estimate the nonlinear term $N$, we rewrite it as
$N=-2\sum_{m,i,j=1}^{2}\sum_{\alpha=1}^{k}\binom{k}{\alpha}\big(\partial_{m}^{\alpha}u_{j}\,\partial_{j}\partial_{m}^{k-\alpha}u_{i},\,\partial_{m}^{k}u_{i}\big)_{L^{2}},$ (3.4)
where $\binom{k}{\alpha}=\frac{k!}{\alpha!(k-\alpha)!}$ is the binomial coefficient. Here we have used the fact that the term corresponding to $\alpha=0$ vanishes due to the divergence-free condition $\nabla\cdot u=0$. We decompose $N$ into three parts:
$N=N_{1}+N_{2}+N_{3},$
where $N_{1}$ corresponds to the case $m=1$ in (3.4),
$N_{1}:=-2\sum_{i,j=1}^{2}\sum_{\alpha=1}^{k}\binom{k}{\alpha}\int_{\mathbb{R}^{2}}\partial_{1}^{\alpha}u_{j}\,\partial_{j}\partial_{1}^{k-\alpha}u_{i}\,\partial_{1}^{k}u_{i}\,dx,$
$N_{2}$ corresponds to the case $m=2$ and $j=1$,
$N_{2}:=-2\sum_{i=1}^{2}\sum_{\alpha=1}^{k}\binom{k}{\alpha}\int_{\mathbb{R}^{2}}\partial_{2}^{\alpha}u_{1}\,\partial_{1}\partial_{2}^{k-\alpha}u_{i}\,\partial_{2}^{k}u_{i}\,dx,$
and $N_{3}$ corresponds to the case $m=2$ and $j=2$,
$N_{3}:=$ $-2\sum_{i=1}^{2}\sum_{\alpha=1}^{k}\binom{k}{\alpha}\int_{\mathbb{R}^{2}}\partial_{2}^{\alpha}u_{2}\,\partial_{2}\partial_{2}^{k-\alpha}u_{i}\,\partial_{2}^{k}u_{i}\,dx$
$=$  $\;2\sum_{i=1}^{2}\sum_{\alpha=1}^{k}\binom{k}{\alpha}\int_{\mathbb{R}^{2}}\partial_{1}\partial_{2}^{\alpha-1}u_{1}\,\partial_{2}\partial_{2}^{k-\alpha}u_{i}\,\partial_{2}^{k}u_{i}\,dx,$
where, in the last step, we have used the incompressibility condition $\partial_{2}u_{2}=-\partial_{1}u_{1}$.
We now estimate $N_{1}$, $N_{2}$, and $N_{3}$ using Lemma 3.1. Although these terms involve different derivative combinations, they all share a common structure covered by Lemma 3.1. Consequently, we obtain
$N=N_{1}+N_{2}+N_{3}\leq C_{0}\|u\|_{H^{k}}\,\|\Lambda_{1}^{s}u\|_{H^{k}}^{2}.$ (3.5)
Inserting (3.5) into (3.3), and integrating over time, we obtain
$\sum_{m=1}^{2}\left\|\partial_{m}^{k}u\right\|_{L^{2}}^{2}+2\nu\int_{0}^{t}\sum_{m=1}^{2}\left\|\Lambda_{1}^{s}\partial_{m}^{k}u\right\|_{L^{2}}^{2}d\tau$
$\leq$ $\sum_{m=1}^{2}\left\|\partial_{m}^{k}u_{0}\right\|_{L^{2}}^{2}+C_{0}\int_{0}^{t}\left\|\Lambda_{1}^{s}u(\tau)\right\|_{H^{k}}^{2}\|u(\tau)\|_{H^{k}}d\tau.$ (3.6)
Combining (3.6) with the $L^{2}$ energy estimate (3.2), we deduce
$\|u(t)\|_{H^{k}}^{2}+2\nu\int_{0}^{t}\|\Lambda_{1}^{s}u(\tau)\|_{H^{k}}^{2}\,d\tau\leq\|u_{0}\|_{H^{k}}^{2}+C_{0}\int_{0}^{t}\|\Lambda_{1}^{s}u(\tau)\|_{H^{k}}^{2}\|u(\tau)\|_{H^{k}}\,d\tau.$ (3.7)
If $\|u_{0}\|_{H^{k}}\leq\varepsilon$ with $\varepsilon<\nu/C_{0}$, then a standard bootstrap argument applied to (3.7) yields
$\|u(t)\|_{H^{k}}^{2}+\nu\int_{0}^{t}\|\Lambda_{1}^{s}u(\tau)\|_{H^{k}}^{2}\,d\tau\leq C\|u_{0}\|_{H^{k}}^{2}\leq C\varepsilon^{2},\qquad\text{for all }t>0.$
This completes the proof of Theorem 1.1. ∎
3.2 Proof of Lemma 3.1
This subsection is devoted to the proof of Lemma 3.1.
Proof of Lemma 3.1.
The proof is divided into three cases according to the range of the fractional exponent $s$.
Case I: $0\leq s\leq\tfrac{1}{4}$.
For $\ell\in\{1,2\}$, we decompose the nonlinear term
$\sum_{\beta=1}^{k}\int_{\mathbb{R}^{2}}\partial_{\ell}^{\beta}f\,\partial_{1}\partial_{\ell}^{k-\beta}f\,\partial_{\ell}^{k}f\,dx$
into three parts:
$\sum_{\beta=1}^{k}\int_{\mathbb{R}^{2}}\partial_{\ell}^{\beta}f\,\partial_{1}\partial_{\ell}^{k-\beta}f\,\partial_{\ell}^{k}f\,dx={}$ $\sum_{\beta=1}^{k-2}\int_{\mathbb{R}^{2}}\partial_{\ell}^{\beta}f\,\partial_{1}\partial_{\ell}^{k-\beta}f\,\partial_{\ell}^{k}f\,dx$
$+\int_{\mathbb{R}^{2}}\partial_{\ell}^{k-1}f\,\partial_{1}\partial_{\ell}f\,\partial_{\ell}^{k}f\,dx+\int_{\mathbb{R}^{2}}\partial_{\ell}^{k}f\,\partial_{1}f\,\partial_{\ell}^{k}f\,dx$
$=:{}$ $F_{1}^{I}+F_{2}^{I}+F_{3}^{I}.$
It then follows from Hölder’s inequality that
$F^{I}_{1}$ $\leq C\sum_{\beta=1}^{k-2}\|\partial^{\beta}_{\ell}f\|_{L^{\frac{1}{2s}}_{x_{1}}L_{x_{2}}^{\infty}}\|\partial_{1}\partial^{k-\beta}_{\ell}f\|_{L^{\frac{2}{1-2s}}_{x_{1}}L_{x_{2}}^{2}}\|\partial^{k}_{\ell}f\|_{L^{\frac{2}{1-2s}}_{x_{1}}L_{x_{2}}^{2}}.$
By Sobolev embedding,
$F^{I}_{1}$ $\leq C\sum_{\beta=1}^{k-2}\|\Lambda_{1}^{\frac{1}{2}-2s}\partial^{\beta}_{\ell}f\|_{L_{x_{1}}^{2}H^{1}_{x_{2}}}\|\Lambda_{1}^{s}\partial_{1}\partial^{k-\beta}_{\ell}f\|_{L^{2}}\|\Lambda_{1}^{s}\partial^{k}_{\ell}f\|_{L^{2}}.$
Since $1\leq\beta\leq k-2$, we have
$F^{I}_{1}$ $\leq C\|f\|_{H^{k}}\|\Lambda_{1}^{s}f\|_{H^{k}}^{2}.$
$F^{I}_{2}$ and $F^{I}_{3}$ can be estimated similarly,
$F^{I}_{2}$ $\leq C\|\partial^{k-1}_{\ell}f\|_{L^{\frac{2}{1-2s}}_{x_{1}}L_{x_{2}}^{\infty}}\|\partial_{1}\partial_{\ell}f\|_{L^{\frac{1}{2s}}_{x_{1}}L_{x_{2}}^{2}}\|\partial^{k}_{\ell}f\|_{L^{\frac{2}{1-2s}}_{x_{1}}L_{x_{2}}^{2}}$
$\leq C\|\Lambda_{1}^{s}\partial^{k-1}_{\ell}f\|_{L_{x_{1}}^{2}H^{1}_{x_{2}}}\|\Lambda_{1}^{\frac{1}{2}-2s}\partial_{1}\partial_{\ell}f\|_{L^{2}}\|\Lambda_{1}^{s}\partial^{k}_{\ell}f\|_{L^{2}}$
$\leq C\|f\|_{H^{k}}\|\Lambda_{1}^{s}f\|_{H^{k}}^{2},$
and
$F^{I}_{3}$ $\leq C\|\partial^{k}_{\ell}f\|_{L^{\frac{2}{1-2s}}_{x_{1}}L_{x_{2}}^{2}}\|\partial_{1}f\|_{L^{\frac{1}{2s}}_{x_{1}}L_{x_{2}}^{\infty}}\|\partial^{k}_{\ell}f\|_{L^{\frac{2}{1-2s}}_{x_{1}}L_{x_{2}}^{2}}$
$\leq C\|\Lambda_{1}^{s}\partial^{k}_{\ell}f\|_{L^{2}}\|\Lambda_{1}^{\frac{1}{2}-2s}\partial_{1}f\|_{L_{x_{1}}^{2}H^{1}_{x_{2}}}\|\Lambda_{1}^{s}\partial^{k}_{\ell}f\|_{L^{2}}$
$\leq C\|f\|_{H^{k}}\|\Lambda_{1}^{s}f\|_{H^{k}}^{2}.$
In fact, when $0\leq s<\tfrac{1}{4}$, the horizontal derivative $\partial_{1}$ on the left-hand side of (3.1) is not essential, and the energy inequality can be closed using only the $\partial_{1}$-regularity provided by Sobolev embedding. However, once $s\geq\tfrac{1}{4}$, the Sobolev embedding
$\dot{H}^{\frac{1}{2}-2s}\hookrightarrow L^{\frac{1}{2s}}$
no longer holds, and a different treatment is required.
Case II: $\frac{1}{4}\leq s<\frac{1}{2}$.
For $\ell\in\{1,2\}$, we decompose
$\sum_{\beta=1}^{k}\int_{\mathbb{R}^{2}}\partial_{\ell}^{\beta}f\,\partial_{1}\partial_{\ell}^{k-\beta}f\,\partial_{\ell}^{k}f\,dx$
into two parts:
$\sum_{\beta=1}^{k}\int_{\mathbb{R}^{2}}\partial_{\ell}^{\beta}f\,\partial_{1}\partial_{\ell}^{k-\beta}f\,\partial_{\ell}^{k}f\,dx={}$ $\sum_{\beta=1}^{k-1}\int_{\mathbb{R}^{2}}\partial_{\ell}^{\beta}f\,\partial_{1}\partial_{\ell}^{k-\beta}f\,\partial_{\ell}^{k}f\,dx$
$+\int_{\mathbb{R}^{2}}\partial_{\ell}^{k}f\,\partial_{1}f\,\partial_{\ell}^{k}f\,dx$
$=:{}$ $F_{1}^{II}+F_{2}^{II}.$
For the term $F_{1}^{II}$, by Hölder’s inequality and Sobolev embedding,
$F^{II}_{1}\leq$ $C\sum_{\beta=1}^{k-1}\|\partial^{\beta}_{\ell}f\|_{L^{\frac{2}{1-2s}}_{x_{1}}L_{x_{2}}^{\infty}}\|\partial_{1}\partial^{k-\beta}_{\ell}f\|_{L^{\frac{1}{s}}_{x_{1}}L_{x_{2}}^{2}}\|\partial^{k}_{\ell}f\|_{L^{2}}$
$\leq$ $C\sum_{\beta=1}^{k-1}\|\Lambda_{1}^{s}\partial^{\beta}_{\ell}f\|_{L_{x_{1}}^{2}H^{1}_{x_{2}}}\|\Lambda_{1}^{\frac{1-2s}{2}}\partial_{1}\partial^{k-\beta}_{\ell}f\|_{L^{2}}\|\partial^{k}_{\ell}f\|_{L^{2}}$
$\leq$ $C\sum_{\beta=1}^{k-1}\|\Lambda_{1}^{s}\partial^{\beta}_{\ell}f\|_{L_{x_{1}}^{2}H^{1}_{x_{2}}}\|\Lambda^{s}_{1}\Lambda_{1}^{1-\frac{4s-1}{2}}\partial^{k-\beta}_{\ell}f\|_{L^{2}}\|\partial^{k}_{\ell}f\|_{L^{2}}.$
By the ranges of $s$ and $\beta$, we obtain
$\|\Lambda_{1}^{s}\partial_{\ell}^{\beta}f\|_{L_{x_{1}}^{2}H_{x_{2}}^{1}}\leq\|\Lambda_{1}^{s}f\|_{H^{k}},\qquad\|\Lambda_{1}^{s}\Lambda_{1}^{1-\frac{4s-1}{2}}\partial_{\ell}^{k-\beta}f\|_{L^{2}}\leq\|\Lambda_{1}^{s}f\|_{H^{k}}.$
Consequently, we deduce
$F_{1}^{II}\leq C\,\|\Lambda_{1}^{s}f\|_{H^{k}}^{2}\,\|f\|_{H^{k}}.$
Similarly, the term $F_{2}^{II}$ can be estimated by applying Hölder’s inequality and Sobolev embedding,
$F^{II}_{2}\leq$ $C\|\partial^{k}_{\ell}f\|_{L^{\frac{2}{1-2s}}_{x_{1}}L_{x_{2}}^{2}}\|\partial_{1}f\|_{L^{\frac{1}{s}}_{x_{1}}L_{x_{2}}^{\infty}}\|\partial^{k}_{\ell}f\|_{L^{2}}$
$\leq$ $C\|\Lambda_{1}^{s}\partial^{k}_{\ell}f\|_{L^{2}}\|\Lambda_{1}^{\frac{1-2s}{2}}\partial_{1}f\|_{L_{x_{1}}^{2}H^{1}_{x_{2}}}\|\partial^{k}_{\ell}f\|_{L^{2}}$
$\leq$ $C\|\Lambda_{1}^{s}\partial^{k}_{\ell}f\|_{L^{2}}\|\Lambda_{1}^{s}\Lambda_{1}^{1-\frac{4s-1}{2}}f\|_{L_{x_{1}}^{2}H^{1}_{x_{2}}}\|\partial^{k}_{\ell}f\|_{L^{2}}$
$\leq$ $C\|\Lambda_{1}^{s}f\|_{H^{k}}^{2}\|f\|_{H^{k}}.$
Case III: $\frac{1}{2}\leq s\leq\frac{3}{4}$.
For $\ell\in\{1,2\}$, by taking Fourier transform with respect to $x_{1}$, we have
$\sum_{\beta=1}^{k}\int_{\mathbb{R}^{2}}\partial^{\beta}_{\ell}f\partial_{1}\partial^{k-\beta}_{\ell}f\partial^{k}_{\ell}fdx=$ $\frac{1}{2\pi}\sum_{\beta=1}^{k}\int_{\mathbb{R}^{2}}\mathcal{F}_{x_{1}}\big(\partial^{\beta}_{\ell}f\partial^{k}_{\ell}f\big)(\xi_{1},x_{2})\mathcal{F}_{x_{1}}\big(\partial_{1}\partial^{k-\beta}_{\ell}f\big)(\xi_{1},x_{2})d\xi_{1}dx_{2}.$
Then we introduce the following notations:
$\mathcal{L}^{h}f:=\mathcal{F}_{x_{1}}^{-1}\left(\chi(\xi_{1})\mathcal{F}_{x_{1}}f\right),\quad\mathcal{H}^{h}f:=\mathcal{F}_{x_{1}}^{-1}\left(\varphi(\xi_{1})\mathcal{F}_{x_{1}}f\right),$
where $(\chi,\varphi)$ are cutoff functions which are defined as
$\chi(\xi_{1}):=\begin{cases}1&\text{if }|\xi_{1}|\leq 1,\\ 0&\text{if }|\xi_{1}|>1,\end{cases}\quad\quad\quad\varphi(\xi_{1}):=\begin{cases}0&\text{if }|\xi_{1}|\leq 1,\\ 1&\text{if }|\xi_{1}|>1.\end{cases}$
Then for each $\ell\in\{1,2\}$, the integral can be decomposed as
$\sum_{\beta=1}^{k}\int_{\mathbb{R}^{2}}\partial^{\beta}_{\ell}f\partial_{1}\partial^{k-\beta}_{\ell}f\partial^{k}_{\ell}fdx$
$=$ $\frac{1}{2\pi}\sum_{\beta=1}^{k}\int_{\mathbb{R}^{2}}\chi(\xi_{1})\mathcal{F}_{x_{1}}\big(\partial^{\beta}_{\ell}f\partial^{k}_{\ell}f\big)(\xi_{1},x_{2})\mathcal{F}_{x_{1}}\big(\partial_{1}\partial^{k-\beta}_{\ell}f\big)(\xi_{1},x_{2})d\xi_{1}dx_{2}$
$+\frac{1}{2\pi}\sum_{\beta=1}^{k}\int_{\mathbb{R}^{2}}\varphi(\xi_{1})\mathcal{F}_{x_{1}}\big(\partial^{\beta}_{\ell}f\partial^{k}_{\ell}f\big)(\xi_{1},x_{2})\mathcal{F}_{x_{1}}\big(\partial_{1}\partial^{k-\beta}_{\ell}f\big)(\xi_{1},x_{2})d\xi_{1}dx_{2}$
$=$ $F^{III}_{L}+F^{III}_{H}.$
Estimate for the low-frequency part:
For $F^{III}_{L}$, by Hölder inequality, we have
$F^{III}_{L}\leq$ $C\sum_{\beta=1}^{k}\big\||\xi_{1}|^{\frac{3}{2}-s}\chi(\xi_{1})\mathcal{F}_{x_{1}}\big(\partial^{k-\beta}_{\ell}f\big)\big\|_{L_{\xi_{1}}^{2}L_{x_{2}}^{\infty}}\big\||\xi_{1}|^{s-\frac{1}{2}}\mathcal{F}_{x_{1}}\big(\partial^{\beta}_{\ell}f\partial^{k}_{\ell}f\big)\big\|_{L_{\xi_{1}}^{2}L_{x_{2}}^{1}}.$ (3.8)
By Plancherel’s theorem in the $x_{1}$-direction, we can pass to the physical space:
$F^{III}_{L}\leq$ $C\sum_{\beta=1}^{k}\big\|\Lambda_{1}^{s}\mathcal{L}^{h}\big(\partial^{k-\beta}_{\ell}f\big)\big\|_{L_{x_{1}}^{2}L_{x_{2}}^{\infty}}\big\|\Lambda_{1}^{s-\frac{1}{2}}(\partial^{\beta}_{\ell}f\partial^{k}_{\ell}f)\big\|_{L_{x_{1}}^{2}L_{x_{2}}^{1}}.$ (3.9)
In the previous step, we used the inequality
$|\xi_{1}|^{\frac{3}{2}-s}\chi(\xi_{1})\leq C|\xi_{1}|^{s}\chi(\xi_{1}),$
which holds for $s\leq\tfrac{3}{4}$. Applying Lemma 2.4, we obtain
$\big\|\Lambda_{1}^{s-\frac{1}{2}}(\partial_{\ell}^{\beta}f\,\partial_{\ell}^{k}f)\big\|_{L_{x_{1}}^{2}L_{x_{2}}^{1}}\leq{}C\big\|(\Lambda_{1}^{s-\frac{1}{2}}\partial_{\ell}^{\beta}f)\,(\partial_{\ell}^{k}f)\big\|_{L_{x_{1}}^{2}L_{x_{2}}^{1}}+C\big\|(\partial_{\ell}^{\beta}f)\,(\Lambda_{1}^{s-\frac{1}{2}}\partial_{\ell}^{k}f)\big\|_{L_{x_{1}}^{2}L_{x_{2}}^{1}}.$ (3.10)
Substituting (3.10) into (3.9), we deduce
$F^{III}_{L}\leq{}$ $C\sum_{\beta=1}^{k}\big\|\Lambda_{1}^{s}\mathcal{L}^{h}(\partial_{\ell}^{k-\beta}f)\big\|_{L_{x_{1}}^{2}L_{x_{2}}^{\infty}}\big\|(\Lambda_{1}^{s-\frac{1}{2}}\partial_{\ell}^{\beta}f)\,(\partial_{\ell}^{k}f)\big\|_{L_{x_{1}}^{2}L_{x_{2}}^{1}}$
$+C\sum_{\beta=1}^{k}\big\|\Lambda_{1}^{s}\mathcal{L}^{h}(\partial_{\ell}^{k-\beta}f)\big\|_{L_{x_{1}}^{2}L_{x_{2}}^{\infty}}\big\|(\partial_{\ell}^{\beta}f)\,(\Lambda_{1}^{s-\frac{1}{2}}\partial_{\ell}^{k}f)\big\|_{L_{x_{1}}^{2}L_{x_{2}}^{1}}.$ (3.11)
By Hölder’s inequality and Sobolev embedding,
$\|(\Lambda_{1}^{s-\frac{1}{2}}\partial^{\beta}_{\ell}f)~(\partial^{k}_{\ell}f)\|_{L_{x_{1}}^{2}L_{x_{2}}^{1}}\leq$ $\|\Lambda_{1}^{s-\frac{1}{2}}\partial^{\beta}_{\ell}f\|_{L_{x_{1}}^{4}L_{x_{2}}^{2}}\|\partial^{k}_{\ell}f\|_{L_{x_{1}}^{4}L_{x_{2}}^{2}}$
$\leq$ $\|\Lambda_{1}^{s-\frac{1}{4}}\partial^{\beta}_{\ell}f\|_{L^{2}}\|\Lambda_{1}^{\frac{1}{4}}\partial^{k}_{\ell}f\|_{L^{2}},$ (3.12)
and
$\|(\partial^{\beta}_{\ell}f)~(\Lambda_{1}^{s-\frac{1}{2}}\partial^{k}_{\ell}f)\|_{L_{x_{1}}^{2}L_{x_{2}}^{1}}\leq$ $\|\partial^{\beta}_{\ell}f\|_{L_{x_{1}}^{4}L_{x_{2}}^{2}}\|\Lambda_{1}^{s-\frac{1}{2}}\partial^{k}_{\ell}f\|_{L_{x_{1}}^{4}L_{x_{2}}^{2}}$
$\leq$ $\|\Lambda_{1}^{\frac{1}{4}}\partial^{\beta}_{\ell}f\|_{L^{2}}\|\Lambda_{1}^{s-\frac{1}{4}}\partial^{k}_{\ell}f\|_{L^{2}}.$ (3.13)
By interpolation, we have
$\|\Lambda_{1}^{s-\frac{1}{4}}\partial^{\beta}_{\ell}f\|_{L^{2}}\|\Lambda_{1}^{\frac{1}{4}}\partial^{k}_{\ell}f\|_{L^{2}}\leq\|\partial^{\beta}_{\ell}f\|_{L^{2}}^{\frac{1}{4s}}\|\Lambda_{1}^{s}\partial^{\beta}_{\ell}f\|_{L^{2}}^{1-\frac{1}{4s}}\|\partial^{k}_{\ell}f\|_{L^{2}}^{1-\frac{1}{4s}}\|\Lambda_{1}^{s}\partial^{k}_{\ell}f\|_{L^{2}}^{\frac{1}{4s}},$ (3.14)
and
$\|\Lambda_{1}^{\frac{1}{4}}\partial^{\beta}_{\ell}f\|_{L^{2}}\|\Lambda_{1}^{s-\frac{1}{4}}\partial^{k}_{\ell}f\|_{L^{2}}$ $\leq\|\partial^{\beta}_{\ell}f\|_{L^{2}}^{1-\frac{1}{4s}}\|\Lambda_{1}^{s}\partial^{\beta}_{\ell}f\|_{L^{2}}^{\frac{1}{4s}}\|\partial^{k}_{\ell}f\|_{L^{2}}^{\frac{1}{4s}}\|\Lambda_{1}^{s}\partial^{k}_{\ell}f\|_{L^{2}}^{1-\frac{1}{4s}}.$ (3.15)
Then substituting (3.12), (3.13), (3.14) and (3.15) into (3.11), using Sobolev embedding $H^{1}_{x_{2}}\hookrightarrow L_{x_{2}}^{\infty}$, we have
$F^{III}_{L}\leq$ $C\sum_{\beta=1}^{k}\|\Lambda_{1}^{s}\partial^{k-\beta}_{\ell}f\|_{L_{x_{1}}^{2}H^{1}_{x_{2}}}\|\partial^{\beta}_{\ell}f\|_{L^{2}}^{\frac{1}{4s}}\|\Lambda_{1}^{s}\partial^{\beta}_{\ell}f\|_{L^{2}}^{1-\frac{1}{4s}}\|\partial^{k}_{\ell}f\|_{L^{2}}^{1-\frac{1}{4s}}\|\Lambda_{1}^{s}\partial^{k}_{\ell}f\|_{L^{2}}^{\frac{1}{4s}}$
$+C\sum_{\beta=1}^{k}\|\Lambda_{1}^{s}\partial^{k-\beta}_{\ell}f\|_{L_{x_{1}}^{2}H^{1}_{x_{2}}}\|\partial^{\beta}_{\ell}f\|_{L^{2}}^{1-\frac{1}{4s}}\|\Lambda_{1}^{s}\partial^{\beta}_{\ell}f\|_{L^{2}}^{\frac{1}{4s}}\|\partial^{k}_{\ell}f\|_{L^{2}}^{\frac{1}{4s}}\|\Lambda_{1}^{s}\partial^{k}_{\ell}f\|_{L^{2}}^{1-\frac{1}{4s}}.$
Since $1\leq\beta\leq k$, we have
$F^{III}_{L}$ $\leq C\|\Lambda_{1}^{s}f\|_{H^{k}}^{2}\|f\|_{H^{k}}.$
Estimate for the high-frequency part:
For each $\ell\in\{1,2\}$, we decompose the analysis of $F^{III}_{H}$ into two parts.
$F^{III}_{H}=$ $\sum_{\beta=1}^{k-1}\int_{\mathbb{R}^{2}}\varphi(\xi_{1})\mathcal{F}_{x_{1}}\big(\partial^{\beta}_{\ell}f\partial^{k}_{\ell}f\big)(\xi_{1},x_{2})\mathcal{F}_{x_{1}}\big(\partial_{1}\partial^{k-\beta}_{\ell}f\big)(\xi_{1},x_{2})d\xi_{1}dx_{2}$
$+\sum_{\beta=1}^{k-1}\int_{\mathbb{R}^{2}}\varphi(\xi_{1})\mathcal{F}_{x_{1}}\big(\partial^{k}_{\ell}f\partial^{k}_{\ell}f\big)(\xi_{1},x_{2})\mathcal{F}_{x_{1}}\big(\partial_{1}f\big)(\xi_{1},x_{2})d\xi_{1}dx_{2}$
$:=$ $F^{III}_{H1}+F^{III}_{H2}.$
For the term $F^{III}_{H1}$, using Hölder’s inequality together with the properties of the cutoff function $\varphi$, and noting that $s>\tfrac{1}{2}$, we obtain
$F^{III}_{H1}\leq$ $C\sum_{\beta=1}^{k-1}\big\|\varphi(\xi_{1})\mathcal{F}_{x_{1}}\big(\partial^{\beta}_{\ell}f\partial^{k}_{\ell}f\big)\big\|_{L^{2}_{\xi_{1}}L_{x_{2}}^{2}}\big\|\mathcal{F}_{x_{1}}(\partial_{1}\partial^{k-\beta}_{\ell}f)\|_{L^{2}_{\xi_{1}}L_{x_{2}}^{2}}$
$\leq$ $C\sum_{\beta=1}^{k-1}\big\||\xi_{1}|^{s-\frac{1}{2}}\varphi(\xi_{1})\mathcal{F}_{x_{1}}\big(\partial^{\beta}_{\ell}f\partial^{k}_{\ell}f\big)\big\|_{L^{2}_{\xi_{1}}L_{x_{2}}^{2}}\big\|\mathcal{F}_{x_{1}}(\partial_{1}\partial^{k-\beta}_{\ell}f)\|_{L^{2}_{\xi_{1}}L_{x_{2}}^{2}}.$
By Plancherel’s theorem and Lemma 2.3,
$\big\||\xi_{1}|^{s-\frac{1}{2}}\varphi(\xi_{1})\mathcal{F}_{x_{1}}\big(\partial^{\beta}_{\ell}f\partial^{k}_{\ell}f\big)\big\|_{L^{2}_{\xi_{1}}L_{x_{2}}^{2}}\leq$ $C\Big\|\big\|\Lambda_{1}^{s-\frac{1}{2}}\big(\partial^{\beta}_{\ell}f\partial^{k}_{\ell}f\big)\big\|_{L_{x_{1}}^{2}}\Big\|_{L_{x_{2}}^{2}}$
$\leq$ $C\Big\|\big\|\Lambda_{1}^{\frac{s}{2}}\partial^{\beta}_{\ell}f\big\|_{L_{x_{1}}^{2}}\|\Lambda_{1}^{\frac{s}{2}}\partial^{k}_{\ell}f\big\|_{L_{x_{1}}^{2}}\Big\|_{L_{x_{2}}^{2}}$
$\leq$ $C\big\|\Lambda_{1}^{\frac{s}{2}}\partial^{\beta}_{\ell}f\big\|_{L_{x_{1}}^{2}L_{x_{2}}^{\infty}}\|\Lambda_{1}^{\frac{s}{2}}\partial^{k}_{\ell}f\big\|_{L^{2}}.$
By interpolation and Sobolev embedding, we have
$F^{III}_{H1}\leq$ $C\sum_{\beta=1}^{k-1}\|\partial^{\beta}_{\ell}f\|_{L_{x_{1}}^{2}H^{1}_{x_{2}}}^{\frac{1}{2}}\|\Lambda_{1}^{s}\partial^{\beta}_{\ell}f\|_{L_{x_{1}}^{2}H^{1}_{x_{2}}}^{\frac{1}{2}}\|\partial^{k}_{\ell}f\|_{L^{2}}^{\frac{1}{2}}\|\Lambda_{1}^{s}\partial^{k}_{\ell}f\|_{L^{2}}^{\frac{1}{2}}\big\|\Lambda_{1}^{s}\Lambda_{1}^{1-s}\partial^{k-\beta}_{\ell}f\|_{L^{2}}.$
Since $1\leq\beta\leq k-1$, we have
$F^{III}_{H1}\leq C\|f\|_{H^{k}}\|\Lambda_{1}^{s}f\|_{H^{k}}^{2}.$
Similarly, by Hölder’s inequality and Lemma 2.3, we obtain
$F^{III}_{H2}\leq C\|\Lambda_{1}^{\frac{s}{2}}\partial^{k}_{\ell}f\|_{L^{2}}\|\Lambda_{1}^{\frac{s}{2}}\partial^{k}_{\ell}f\|_{L^{2}}\|\partial_{1}f\|_{L_{x_{1}}^{2}L_{x_{2}}^{\infty}}.$
Then, by interpolation and Sobolev embedding, we obtain
$F^{III}_{H2}\leq C\|\partial^{k}_{\ell}f\|_{L^{2}}\|\Lambda_{1}^{s}\partial^{k}_{\ell}f\|_{L^{2}}\|\Lambda_{1}^{s}\Lambda_{1}^{1-s}f\|_{L_{x_{1}}^{2}H^{1}_{x_{2}}}\leq C\|f\|_{H^{k}}\|\Lambda_{1}^{s}f\|_{H^{k}}^{2}.$
This completes the proof of Lemma 3.1. ∎
4 Proof of Theorem 1.2
In this section, we prove Theorem 1.2. The argument is based on a bootstrap approach combined with refined energy estimates that yield global-in-time decay and stability.
Bootstrap assumptions. We assume that there exists a constant $C_{0}>0$ such that, for all $t\in[0,T)$, the solution satisfies the following uniform bounds:
$\|\Lambda_{1}^{-\sigma}u(t)\|_{H^{k}}$ $\leq C_{0}\varepsilon,$ (4.1)
$\|u_{1}(t)\|_{L^{2}}$ $\leq C_{0}\varepsilon(1+t)^{-\sigma/2s},$ (4.2)
$\|u_{2}(t)\|_{L^{2}}$ $\leq C_{0}\varepsilon(1+t)^{-(\sigma+\frac{2}{3})/2s},$ (4.3)
$\|\partial_{2}u_{1}(t)\|_{L^{2}}$ $\leq C_{0}\varepsilon(1+t)^{-\sigma/2s},$ (4.4)
$\|\partial_{1}u_{1}(t)\|_{L^{2}}$ $\leq C_{0}\varepsilon(1+t)^{-(\sigma+1)/2s},$ (4.5)
$\|\partial_{1}u_{2}(t)\|_{L^{2}}$ $\leq C_{0}\varepsilon(1+t)^{-2\sigma/s}.$ (4.6)
Improved estimates. Based on the initial condition (1.4) and the above a priori bounds, we will establish the improved estimates
$\|\Lambda_{1}^{-\sigma}u(t)\|_{H^{k}}$ $\leq\tfrac{1}{2}C_{0}\varepsilon,$ (4.7)
$\|u_{1}(t)\|_{L^{2}}$ $\leq\tfrac{1}{2}C_{0}\varepsilon(1+t)^{-\sigma/2s},$ (4.8)
$\|u_{2}(t)\|_{L^{2}}$ $\leq\tfrac{1}{2}C_{0}\varepsilon(1+t)^{-(\sigma+\frac{2}{3})/2s},$ (4.9)
$\|\partial_{2}u_{1}(t)\|_{L^{2}}$ $\leq\tfrac{1}{2}C_{0}\varepsilon(1+t)^{-\sigma/2s},$ (4.10)
$\|\partial_{1}u_{1}(t)\|_{L^{2}}$ $\leq\tfrac{1}{2}C_{0}\varepsilon(1+t)^{-(\sigma+1)/2s},$ (4.11)
$\|\partial_{1}u_{2}(t)\|_{L^{2}}$ $\leq\tfrac{1}{2}C_{0}\varepsilon(1+t)^{-2\sigma/s}.$ (4.12)
Once (4.7)–(4.12) are established, a standard bootstrap argument implies that the maximal existence time satisfies $T=+\infty$, and hence all decay estimates hold globally in time.
4.1 Estimate for $\|\Lambda_{1}^{-\sigma}u\|_{L^{2}}$
We establish the estimate for $\|\Lambda_{1}^{-\sigma}u\|_{L^{2}}$ through the following energy argument. Applying $\Lambda_{1}^{-\sigma}$ to the equation and taking the $L^{2}$ inner product with $\Lambda_{1}^{-\sigma}u$, we obtain
$\frac{1}{2}\frac{d}{dt}\|\Lambda_{1}^{-\sigma}u\|_{L^{2}}^{2}+\nu\|\Lambda_{1}^{s-\sigma}u\|_{L^{2}}^{2}=-\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}(u\cdot\nabla u)\cdot\Lambda_{1}^{-\sigma}u\,dx.$ (4.13)
The right-hand side of (4.13) can be decomposed into two parts and estimated as
$\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}(u\cdot\nabla u)\cdot\Lambda_{1}^{-\sigma}u\,dx=$ $\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}(u_{1}\partial_{1}u)\cdot\Lambda_{1}^{-\sigma}u\,dx+\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}(u_{2}\partial_{2}u)\cdot\Lambda_{1}^{-\sigma}u\,dx$
$\leq$ $C\left(\left\|\Lambda_{1}^{-\sigma}(u_{1}\partial_{1}u)\right\|_{L^{2}}+\left\|\Lambda_{1}^{-\sigma}(u_{2}\partial_{2}u)\right\|_{L^{2}}\right)\left\|\Lambda_{1}^{-\sigma}u\right\|_{L^{2}}$
$:=$ $M_{0}.$
Integrating in time on equation (4.13), we obtain
$\|\Lambda_{1}^{-\sigma}u(t)\|_{L^{2}}^{2}+2\nu\int_{0}^{t}\|\Lambda_{1}^{s-\sigma}u(\tau)\|_{L^{2}}^{2}d\tau=\|\Lambda_{1}^{-\sigma}u_{0}\|_{L^{2}}^{2}+2\int_{0}^{t}M_{0}(\tau)d\tau.$ (4.14)
By Lemma 2.6,
$M_{0}$ $\leq C\Big(\left\|\partial_{2}u_{2}\right\|_{L^{2}}^{\frac{1}{2}}\left\|u_{2}\right\|_{L^{2}}^{\sigma}\left\|\partial_{1}u_{2}\right\|_{L^{2}}^{\frac{1}{2}-\sigma}\left\|\partial_{2}u\right\|_{L^{2}}$
$\quad+\left\|\partial_{2}u_{1}\right\|_{L^{2}}^{\frac{1}{2}}\left\|u_{1}\right\|_{L^{2}}^{\sigma}\left\|\partial_{1}u_{1}\right\|_{L^{2}}^{\frac{1}{2}-\sigma}\left\|\partial_{1}u\right\|_{L^{2}}\Big)\left\|\Lambda_{1}^{-\sigma}u\right\|_{L^{2}}.$
Using the divergence-free condition of $u$, we have
$M_{0}$ $\leq C\Big(\left\|\partial_{1}u_{1}\right\|_{L^{2}}^{\frac{1}{2}}\left\|u_{2}\right\|_{L^{2}}^{\sigma}\left\|\partial_{1}u_{2}\right\|_{L^{2}}^{\frac{1}{2}-\sigma}\left\|\partial_{2}u\right\|_{L^{2}}$
$\quad+\left\|\partial_{2}u_{1}\right\|_{L^{2}}^{\frac{1}{2}}\left\|u_{1}\right\|_{L^{2}}^{\sigma}\left\|\partial_{1}u_{1}\right\|_{L^{2}}^{\frac{1}{2}-\sigma}\left\|\partial_{1}u\right\|_{L^{2}}\Big)\left\|\Lambda_{1}^{-\sigma}u\right\|_{L^{2}}.$ (4.15)
By substituting (4.2), (4.3), (4.4), (4.5) and (4.6) into (4.15), we deduce
$M_{0}\leq CC_{0}^{3}\varepsilon^{3}\big((1+t)^{-\frac{-18\sigma^{2}+25\sigma+3}{12s}}+(1+t)^{-\frac{2\sigma+1}{2s}}\big).$
It can be verified that
$s<\tfrac{5}{12}+\sigma<\tfrac{-18\sigma^{2}+25\sigma+3}{12}.$
As a consequence, we obtain
$\int_{0}^{t}M_{0}(\tau)\,d\tau\leq CC_{0}^{3}\varepsilon^{3}.$
Combining this bound with the assumption $\|\Lambda_{1}^{-\sigma}u_{0}\|_{L^{2}}\leq\varepsilon$, it follows from (4.14) that
$\|\Lambda_{1}^{-\sigma}u(t)\|_{L^{2}}^{2}\leq\varepsilon^{2}+2CC_{0}^{3}\varepsilon^{3}.$
For $\varepsilon>0$ sufficiently small so that $2CC_{0}^{3}\varepsilon\leq\tfrac{1}{2}$, we further deduce
$\|\Lambda_{1}^{-\sigma}u(t)\|_{L^{2}}\leq\sqrt{\tfrac{3}{2}}\,\varepsilon\leq\tfrac{C_{0}}{2}\,\varepsilon.$
4.2 Estimate for $\sum_{m=1}^{2}\|\Lambda_{1}^{-\sigma}\partial_{m}^{k}u\|_{L^{2}}$
Applying $\Lambda_{1}^{-\sigma}\partial_{m}^{k}$ to the first equation in (1.1), and then taking the $L^{2}$ inner product with $\Lambda_{1}^{-\sigma}\partial_{m}^{k}u$, we have
$\frac{d}{dt}\sum_{m=1}^{2}\left\|\Lambda_{1}^{-\sigma}\partial_{m}^{k}u\right\|_{L^{2}}^{2}+2\nu\sum_{m=1}^{2}\left\|\Lambda_{1}^{-\sigma}\Lambda_{1}^{s}\partial_{m}^{k}u\right\|_{L^{2}}^{2}$
$=$ $-2\sum_{m=1}^{2}\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}\partial_{m}^{k}(u\cdot\nabla u)\cdot\Lambda_{1}^{-\sigma}\partial_{m}^{k}u~dx:=M_{1}.$ (4.16)
First, we expand the nonlinear term using the Leibniz rule and divide it into three parts.
$M_{1}$ $=-2\sum_{m,i,j=1}^{2}\sum_{\alpha=1}^{k}\binom{k}{\alpha}\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}(\partial_{m}^{\alpha}u_{j}\partial_{j}\partial_{m}^{k-\alpha}u_{i})~\Lambda_{1}^{-\sigma}\partial_{m}^{k}u_{i}\,dx$
$=M_{11}+M_{12}+M_{13},$
where the terms $M_{11}$, $M_{12}$, and $M_{13}$ correspond to different choices of the indices $(m,j)$ and are defined as follows. The term $M_{11}$ corresponds to the case $m=1$:
$M_{11}:=\sum_{i,j=1}^{2}\sum_{\alpha=1}^{k}\binom{k}{\alpha}\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}\big(\partial_{1}^{\alpha}u_{j}\,\partial_{j}\partial_{1}^{k-\alpha}u_{i}\big)\,\Lambda_{1}^{-\sigma}\partial_{1}^{k}u_{i}\,dx.$
The term $M_{12}$ corresponds to the case $m=2$ and $j=1$:
$M_{12}:=\sum_{i=1}^{2}\sum_{\alpha=1}^{k}\binom{k}{\alpha}\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}\big(\partial_{2}^{\alpha}u_{1}\,\partial_{1}\partial_{2}^{k-\alpha}u_{i}\big)\,\Lambda_{1}^{-\sigma}\partial_{2}^{k}u_{i}\,dx.$
Finally, the term $M_{13}$ corresponds to the case $m=2$ and $j=2$:
$M_{13}$ $:=\sum_{i=1}^{2}\sum_{\alpha=1}^{k}\binom{k}{\alpha}\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}\big(\partial_{2}^{\alpha}u_{2}\,\partial_{2}\partial_{2}^{k-\alpha}u_{i}\big)\,\Lambda_{1}^{-\sigma}\partial_{2}^{k}u_{i}\,dx$
$=-\sum_{i=1}^{2}\sum_{\alpha=1}^{k}\binom{k}{\alpha}\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}\big(\partial_{1}\partial_{2}^{\alpha-1}u_{1}\,\partial_{2}\partial_{2}^{k-\alpha}u_{i}\big)\,\Lambda_{1}^{-\sigma}\partial_{2}^{k}u_{i}\,dx,$
where we have used the incompressibility condition $\partial_{2}u_{2}=-\partial_{1}u_{1}$.
To estimate each component of $M_{1}$, it suffices to establish the following lemma.
Lemma 4.1.
Let $1\leq\beta\leq k$ be an integer. For $\ell\in\{1,2\}$ and $\tfrac{3}{4}\leq s<\sigma+\tfrac{1}{2}$, we have
$\sum_{\beta=1}^{k}\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}\big(\partial^{\beta}_{\ell}u\partial_{1}\partial^{k-\beta}_{\ell}u\big)\Lambda_{1}^{-\sigma}\partial^{k}_{\ell}udx\leq C\|\Lambda_{1}^{-\sigma+s}u\|_{H^{k}}^{2}\|\Lambda_{1}^{-\sigma}u\|_{H^{k}}.$
We now proceed to estimate the term $M_{1}$. The proof of Lemma 4.1 will be given later. Assuming Lemma 4.1 for the moment, we obtain
$|M_{1}|\leq|M_{11}|+|M_{12}|+|M_{13}|\leq C\|\Lambda_{1}^{-\sigma+s}u\|_{H^{k}}^{2}\|\Lambda_{1}^{-\sigma}u\|_{H^{k}}.$
Substituting this bound into (4.2) and integrating in time yield
$\sum_{m=1}^{2}\|\Lambda_{1}^{-\sigma}\partial_{m}^{k}u(t)\|_{L^{2}}^{2}+2\nu\int_{0}^{t}\sum_{m=1}^{2}\|\Lambda_{1}^{s}\Lambda_{1}^{-\sigma}\partial_{m}^{k}u(\tau)\|_{L^{2}}^{2}\,d\tau$
$\leq{}$ $\sum_{m=1}^{2}\|\Lambda_{1}^{-\sigma}\partial_{m}^{k}u_{0}\|_{L^{2}}^{2}+C\int_{0}^{t}\|\Lambda_{1}^{-\sigma+s}u(\tau)\|_{H^{k}}^{2}\|\Lambda_{1}^{-\sigma}u(\tau)\|_{H^{k}}\,d\tau.$ (4.17)
Combining (4.17) with the $L^{2}$ energy estimate (4.14), we arrive at
$\|\Lambda_{1}^{-\sigma}u(t)\|_{H^{k}}^{2}$ $+2\nu\int_{0}^{t}\|\Lambda_{1}^{s}\Lambda_{1}^{-\sigma}u(\tau)\|_{H^{k}}^{2}\,d\tau$
$\leq{}$ $\|\Lambda_{1}^{-\sigma}u_{0}\|_{H^{k}}^{2}+C\int_{0}^{t}\|\Lambda_{1}^{-\sigma+s}u(\tau)\|_{H^{k}}^{2}\|\Lambda_{1}^{-\sigma}u(\tau)\|_{H^{k}}\,d\tau.$ (4.18)
If $\|\Lambda_{1}^{-\sigma}u_{0}\|_{H^{k}}<\varepsilon$ with $\varepsilon<\nu/C$, then a standard bootstrap argument applied to (4.18) yields
$\|\Lambda_{1}^{-\sigma}u(t)\|_{H^{k}}^{2}+\nu\int_{0}^{t}\|\Lambda_{1}^{s}\Lambda_{1}^{-\sigma}u(\tau)\|_{H^{k}}^{2}\,d\tau\leq C\|\Lambda_{1}^{-\sigma}u_{0}\|_{H^{k}}^{2}\leq C\varepsilon^{2}.$
We now turn to the proof of Lemma 4.1.
Proof of Lemma 4.1. We decompose $J$ into two parts:
$J$ $=\sum_{\beta=1}^{k}\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}\big(\partial_{\ell}^{\beta}u\,\partial_{1}\partial_{\ell}^{k-\beta}u\big)\,\Lambda_{1}^{-\sigma}\partial_{\ell}^{k}u\,dx$
$=\sum_{\beta=1}^{k-1}\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}\big(\partial_{\ell}^{\beta}u\,\partial_{1}\partial_{\ell}^{k-\beta}u\big)\,\Lambda_{1}^{-\sigma}\partial_{\ell}^{k}u\,dx+\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}\big(\partial_{\ell}^{k}u\,\partial_{1}u\big)\,\Lambda_{1}^{-\sigma}\partial_{\ell}^{k}u\,dx$
$=:J_{1}+J_{2}.$
Since $s<\tfrac{1}{2}+\sigma$, we may employ Cauchy-Schwarz inequality and then apply Lemma 2.3 to estimate $J_{1}$:
$J_{1}$ $\leq\sum_{\beta=1}^{k-1}\|\Lambda_{1}^{-\sigma}(\partial_{\ell}^{\beta}u\,\partial_{1}\partial_{\ell}^{k-\beta}u)\|_{L^{2}}\|\Lambda_{1}^{-\sigma}\partial_{\ell}^{k}u\|_{L^{2}}$
$\leq\sum_{\beta=1}^{k-1}\|\Lambda_{1}^{-\sigma+s}\partial_{\ell}^{\beta}u\|_{L_{x_{1}}^{2}L_{x_{2}}^{\infty}}\,\|\Lambda_{1}^{\frac{1}{2}-s}\partial_{1}\partial_{\ell}^{k-\beta}u\|_{L^{2}}\,\|\Lambda_{1}^{-\sigma}\partial_{\ell}^{k}u\|_{L^{2}}$
$\leq\sum_{\beta=1}^{k-1}\|\Lambda_{1}^{-\sigma+s}\partial_{\ell}^{\beta}u\|_{L_{x_{1}}^{2}H_{x_{2}}^{1}}\,\|\Lambda_{1}^{-\sigma+s}\Lambda_{1}^{\frac{3}{2}-2s+\sigma}\partial_{\ell}^{k-\beta}u\|_{L^{2}}\,\|\Lambda_{1}^{-\sigma}\partial_{\ell}^{k}u\|_{L^{2}}.$
By the admissible ranges of $\beta$, $s$, and $\sigma$, we have
$\sum_{\beta=1}^{k-1}\|\Lambda_{1}^{-\sigma+s}\partial_{\ell}^{\beta}u\|_{L_{x_{1}}^{2}H_{x_{2}}^{1}}\leq\|\Lambda_{1}^{-\sigma+s}u\|_{H^{k}},\qquad\sum_{\beta=1}^{k-1}\|\Lambda_{1}^{-\sigma+s}\Lambda_{1}^{\frac{3}{2}-2s+\sigma}\partial_{\ell}^{k-\beta}u\|_{L^{2}}\leq\|\Lambda_{1}^{-\sigma+s}u\|_{H^{k}}.$
Consequently,
$J_{1}\leq C\|\Lambda_{1}^{-\sigma+s}u\|_{H^{k}}^{2}\|\Lambda_{1}^{-\sigma}u\|_{H^{k}}.$ (4.19)
The term $J_{2}$ can be estimated in a similar manner. By Hölder’s inequality and Lemma 2.3, we obtain
$J_{2}$ $\leq\|\Lambda_{1}^{-\sigma}(\partial_{\ell}^{k}u\,\partial_{1}u)\|_{L^{2}}\|\Lambda_{1}^{-\sigma}\partial_{\ell}^{k}u\|_{L^{2}}$
$\leq\|\Lambda_{1}^{-\sigma+s}\partial_{\ell}^{k}u\|_{L^{2}}\,\|\Lambda_{1}^{\frac{1}{2}-s}\partial_{1}u\|_{L_{x_{1}}^{2}L_{x_{2}}^{\infty}}\,\|\Lambda_{1}^{-\sigma}\partial_{\ell}^{k}u\|_{L^{2}}$
$\leq\|\Lambda_{1}^{-\sigma+s}\partial_{\ell}^{k}u\|_{L^{2}}\,\|\Lambda_{1}^{-\sigma+s}\Lambda_{1}^{\frac{3}{2}-2s+\sigma}u\|_{L_{x_{1}}^{2}H_{x_{2}}^{1}}\,\|\Lambda_{1}^{-\sigma}\partial_{\ell}^{k}u\|_{L^{2}}.$ (4.20)
Combining (4.19) and (4.20), we conclude that
$J\leq C\|\Lambda_{1}^{-\sigma+s}u\|_{H^{k}}^{2}\|\Lambda_{1}^{-\sigma}u\|_{H^{k}}.$
This completes the proof of Lemma 4.1.
4.3 The decay of $u_{1}$
To establish the decay estimate of $\|u\|_{L^{2}}$, let $\mathbb{P}$ be the Leray projector onto divergence-free fields. By Duhamel’s formulation, the solution can be expressed as:
$u(t)=e^{\nu\Lambda_{1}^{2s}t}u_{0}-\int_{0}^{t}e^{\nu\Lambda_{1}^{2s}(t-\tau)}\mathbb{P}(u\cdot\nabla u)(\tau)d\tau.$
For the first velocity component $u_{1}$, we take the $L^{2}$ norm and use the boundedness of the Leray projector $\mathbb{P}$ to obtain
$\|u_{1}(t)\|_{L^{2}}$ $\leq\|e^{\nu\Lambda_{1}^{2s}t}u_{0,1}\|_{L^{2}}+C\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}(u\cdot\nabla u_{1})(\tau)\|_{L^{2}}\,d\tau$
$\quad+C\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}(u\cdot\nabla u_{2})(\tau)\|_{L^{2}}\,d\tau.$
The linear term admits the decay estimate
$\|e^{\nu\Lambda_{1}^{2s}t}u_{0,1}\|_{L^{2}}$ $\leq C(1+\nu t)^{-\frac{\sigma}{2s}}\Big(\|\Lambda_{1}^{-\sigma}u_{0,1}\|_{L^{2}}+\|u_{0}\|_{L^{2}}\Big)$
$\leq C\varepsilon(1+\nu t)^{-\frac{\sigma}{2s}}\leq\frac{C_{0}}{4}\varepsilon(1+\nu t)^{-\frac{\sigma}{2s}},$
where we choose $C_{0}$ such that $C_{0}\geq 4C$.
For the nonlinear term, due to the decay of $u_{2}$ is faster than $u_{1}$, we only consider the term contain $u\cdot\nabla u_{1}$. Splitting it into the following two terms:
$\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}(u\cdot\nabla u_{1})(\tau)\|_{L^{2}}\,d\tau$ $\leq\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}(u_{1}\partial_{1}u_{1})(\tau)\|_{L^{2}}\,d\tau$
$\quad+\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}(u_{2}\partial_{2}u_{1})(\tau)\|_{L^{2}}\,d\tau$
$=:I_{1}+I_{2}.$
Applying Lemma 2.7 and then employing Lemma 2.1, using the identity $\partial_{1}u_{1}=-\partial_{2}u_{2}$, and substituting the bootstrap bounds (4.2), (4.3), (4.4), and (4.5), we obtain
$I_{1}+I_{2}\leq$ $C\int_{0}^{t}(t-\tau)^{-\frac{1}{4s}}\|u_{1}(\tau)\|_{L^{2}}^{\frac{1}{2}}\|\partial_{2}u_{1}(\tau)\|_{L^{2}}^{\frac{1}{2}}\|\partial_{1}u_{1}(\tau)\|_{L^{2}}d\tau$
$+C\int_{0}^{t}(t-\tau)^{-\frac{1}{4s}}\|u_{2}(\tau)\|_{L^{2}}^{\frac{1}{2}}\|\partial_{2}u_{2}(\tau)\|_{L^{2}}^{\frac{1}{2}}\|\partial_{2}u_{1}(\tau)\|_{L^{2}}d\tau$
$\leq$ $CC_{0}^{2}\varepsilon^{2}\int_{0}^{t}(t-\tau)^{-\frac{1}{4s}}(\tau+1)^{-\frac{2\sigma+1}{2s}}d\tau$
$+CC_{0}^{2}\varepsilon^{2}\int_{0}^{t}(t-\tau)^{-\frac{1}{4s}}(\tau+1)^{-\frac{12\sigma+5}{12s}}d\tau.$
Then applying Lemma 2.8 with $\alpha<1$, $\beta>\alpha$, we can derive
$I_{1}+I_{2}\leq CC_{0}^{2}\varepsilon^{2}(t+1)^{-\frac{\sigma}{2s}}.$
Combining the estimates above, we obtain
$\|u_{1}(t)\|_{L^{2}}\leq\left(\frac{C_{0}}{4}+4CC_{0}^{2}\varepsilon\right)\varepsilon(1+t)^{-\frac{\sigma}{2s}}.$
When $\varepsilon$ is sufficiently small (specifically, $\varepsilon\leq(16CC_{0})^{-1}$), we can deduce the bootstrap conclusion (4.8).
4.4 The decay of $u_{2}$
Recall the vorticity equation:
$\partial_{t}\omega+\nu\Lambda_{1}^{2s}\omega+\nabla\times(u\cdot\nabla u)=0.$
Since $u$ is divergence-free, there exists a stream function $\psi$ such that $u=\nabla^{\perp}\psi$, and the Biot–Savart law yields $u=\nabla^{\perp}\Delta^{-1}\omega$. Consequently, the second velocity component $u_{2}$ satisfies
$\partial_{t}u_{2}+\nu\Lambda_{1}^{2s}u_{2}=-\partial_{1}\Delta^{-1}\nabla\times(u\cdot\nabla u).$ (4.21)
and the integral form:
$u_{2}(t)=e^{\nu\Lambda_{1}^{2s}t}u_{0,2}-\int_{0}^{t}e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}\Delta^{-1}\nabla\times(u\cdot\nabla u)(\tau)d\tau.$
Taking the $L^{2}$ norm, we obtain
$\|u_{2}(t)\|_{L^{2}}\leq\|e^{\nu\Lambda_{1}^{2s}t}u_{0,2}\|_{L^{2}}+\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}\Delta^{-1}\nabla\times(u\cdot\nabla u)(\tau)\|_{L^{2}}\,d\tau.$
The linear term admits the decay estimate
$\|e^{\nu\Lambda_{1}^{2s}t}u_{0,2}\|_{L^{2}}$ $\leq C(1+\nu t)^{-\frac{\sigma+1}{2s}}\big(\|\Lambda_{1}^{-\sigma}\psi_{0}\|_{L^{2}}+\|u_{0,2}\|_{L^{2}}\big)$
$\leq C\varepsilon(1+\nu t)^{-\frac{\sigma+1}{2s}}\leq\frac{C_{0}}{4}\varepsilon(1+\nu t)^{-\frac{\sigma+\frac{2}{3}}{2s}},$ (4.22)
where we used the smallness of the initial data and chose $C_{0}$ suitable large. To estimate the nonlinear term, we rewrite
$\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}\Delta^{-1}\nabla\times(u\cdot\nabla u)(\tau)\|_{L^{2}}\,d\tau$
$=$ $\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}\Delta^{-1}\nabla\times\mbox{div}\ (u\otimes u)(\tau)\|_{L^{2}}\,d\tau$
$\leq{}$ $\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}\Delta^{-1}\nabla\times\partial_{1}(u_{1}u)(\tau)\|_{L^{2}}\,d\tau$
$+\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}\Delta^{-1}\nabla\times\partial_{2}(u_{2}u)(\tau)\|_{L^{2}}\,d\tau$
$=:{}$ $II_{1}+II_{2}.$
For $II_{1}$, by Lemma 2.7, we take out $\Lambda_{1}^{\sigma+\frac{2}{3}}$ and applying Lemma 2.3 and Lemma 2.4, we have
$II_{1}\leq$ $\int_{0}^{t}(t-\tau)^{-\frac{\sigma+\frac{2}{3}}{2s}}\|\Lambda^{-1}\Lambda_{1}^{\frac{4}{3}-\sigma}(u_{1}u)\|_{L^{2}}d\tau$
$\leq$ $\int_{0}^{t}(t-\tau)^{-\frac{\sigma+\frac{2}{3}}{2s}}\|\Lambda^{-\frac{1}{2}-\sigma}\Lambda_{1}^{\frac{5}{6}}(u_{1}u)\|_{L^{2}}d\tau$
$\leq$ $\int_{0}^{t}(t-\tau)^{-\frac{\sigma+\frac{2}{3}}{2s}}\Big(\|\Lambda^{-\frac{1}{2}-\sigma}(\Lambda_{1}^{\frac{5}{6}}u_{1})u\|_{L^{2}}+\|\Lambda^{-\frac{1}{2}-\sigma}u_{1}(\Lambda_{1}^{\frac{5}{6}}u)\|_{L^{2}}\Big)d\tau.$ (4.23)
By applying Lemma 2.3 with $s=-\frac{1}{2}-\sigma,s_{1}=\frac{1}{2}-\sigma,s_{2}=0$ and interpolation, we obtain
$\int_{0}^{t}(t-\tau)^{-\frac{\sigma+\frac{2}{3}}{2s}}\Big(\|\Lambda^{-\frac{1}{2}-\sigma}(\Lambda_{1}^{\frac{5}{6}}u_{1})u\|_{L^{2}}+\|\Lambda^{-\frac{1}{2}-\sigma}u_{1}(\Lambda_{1}^{\frac{5}{6}}u)\|_{L^{2}}\Big)d\tau$
$\leq$ $\int_{0}^{t}(t-\tau)^{-\frac{\sigma+\frac{2}{3}}{2s}}\|\Lambda^{\frac{1}{2}-\sigma}u\|_{L^{2}}\|\Lambda_{1}^{\frac{5}{6}}u\|_{L^{2}}d\tau$
$\leq$ $\int_{0}^{t}(t-\tau)^{-\frac{\sigma+\frac{2}{3}}{2s}}\|u\|_{L^{2}}^{\frac{1}{2}+\sigma}\|\nabla u\|_{L^{2}}^{\frac{1}{2}-\sigma}\|u\|_{L^{2}}^{\frac{1}{6}}\|\partial_{1}u\|_{L^{2}}^{\frac{5}{6}}d\tau$
$\leq$ $CC_{0}^{2}\varepsilon^{2}\int_{0}^{t}(t-\tau)^{-\frac{\sigma+\frac{2}{3}}{2s}}(\tau+1)^{-\frac{2\sigma+\frac{5}{6}}{2s}}d\tau$
$\leq$ $CC_{0}^{2}\varepsilon^{2}(t+1)^{-\frac{\sigma+\frac{2}{3}}{2s}}.$
It follows from the boundedness of the Riesz operator $\Delta^{-1}\nabla\times\partial_{2}$, together with Lemma 2.7 and subsequently Lemma 2.1, that
$II_{2}\leq$ $\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}\Delta^{-1}\nabla\times\partial_{2}(u_{2}u)(\tau)\|_{L^{2}}d\tau$
$\leq$ $C\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}(u_{2}u)(\tau)\|_{L^{2}}d\tau$
$\leq$ $C\int_{0}^{t}(t-\tau)^{-\frac{3}{4s}}\|(u_{2}u)(\tau)\|_{L_{x_{1}}^{1}L_{x_{2}}^{2}}d\tau$
$\leq$ $C\int_{0}^{t}(t-\tau)^{-\frac{3}{4s}}\|u_{2}\|_{L^{2}}^{\frac{1}{2}}\|\partial_{2}u_{2}\|_{L^{2}}^{\frac{1}{2}}\|u\|_{L^{2}}d\tau$
$\leq$ $CC_{0}^{2}\varepsilon^{2}\int_{0}^{t}(t-\tau)^{-\frac{3}{4s}}(\tau+1)^{-\frac{12\sigma+5}{12s}}d\tau$
$\leq$ $CC_{0}^{2}\varepsilon^{2}(t+1)^{-\frac{\sigma+\frac{2}{3}}{2s}}.$ (4.24)
Combining (4.22), (4.23) and (4.24), we have
$\|u_{2}(t)\|_{L^{2}}\leq\left(\frac{C_{0}}{4}+4CC_{0}^{2}\varepsilon\right)\varepsilon(1+t)^{-\frac{\sigma+\frac{2}{3}}{2s}}.$
By taking $\varepsilon$ sufficiently small, we can get the bootstrap conclusion (4.9).
4.5 The decay of $\partial_{2}u_{1}$
By Duhamel’s principle, we have
$\|\partial_{2}u_{1}(t)\|_{L^{2}}\leq\|e^{\nu\Lambda_{1}^{2s}t}\partial_{2}u_{0,1}\|_{L^{2}}+C\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{2}(u\cdot\nabla u)(\tau)\|_{L^{2}}d\tau.$
For the linear term involving initial data $\partial_{2}u_{0}$, we can obtain the decay estimate through Lemma 2.3.
$\|e^{\nu\Lambda_{1}^{2s}t}\partial_{2}u_{0,1}\|_{L^{2}}\leq C(1+\nu t)^{-\frac{\sigma}{2s}}\left(\|\Lambda_{1}^{-\sigma}\partial_{2}u_{0,1}\|_{L^{2}}+\|\partial_{2}u_{0,1}\|_{L^{2}}\right)\leq\frac{C_{0}}{4}\varepsilon(1+\nu t)^{-\frac{\sigma}{2s}}.$ (4.25)
For the nonlinear term, due to $u_{2}$ decay more faster than $u_{1}$, we only consider the case contain $u\cdot\nabla u_{1}$. After eliminating some terms using the divergence-free condition, it can be estimated as
$\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{2}(u\cdot\nabla u_{1})(\tau)\|_{L^{2}}d\tau$
$\leq$ $\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}u_{1}\partial_{1}\partial_{2}u_{1}(\tau)\|_{L^{2}}d\tau+\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}u_{2}\partial_{2}^{2}u_{1}(\tau)\|_{L^{2}}d\tau$
$:=$ $III_{1}+III_{2}.$
By using Lemma 2.7 first and then applying Lemma 2.1, we have
$III_{1}\leq C\int_{0}^{t}(t-\tau)^{-\frac{1}{4s}}\|u_{1}(\tau)\|_{L^{2}}^{\frac{1}{2}}\|\partial_{2}u_{1}(\tau)\|_{L^{2}}^{\frac{1}{2}}\|\partial_{1}\partial_{2}u_{1}(\tau)\|_{L^{2}}d\tau.$
It follows from interpolation with $0<\theta<\min\{\frac{1}{9},\frac{1-2\sigma}{2\sigma}\}$ that
$\|\partial_{1}\partial_{2}u_{1}(\tau)\|_{L^{2}}\leq\|\partial_{1}u_{1}\|_{L^{2}}^{1-\theta}\|\partial_{1}\partial_{2}^{\frac{1}{\theta}}u_{1}\|_{L^{2}}^{\theta}\leq C_{0}\varepsilon(1+t)^{-\frac{(\sigma+1)(1-\theta)}{2s}}.$ (4.26)
Similarly, for $0<\theta<\min\{\frac{1}{9},\frac{1-2\sigma}{2\sigma}\}$,
$\|\partial_{2}^{2}u_{1}(\tau)\|_{L^{2}}\leq C_{0}\varepsilon(1+t)^{-\frac{\sigma(1-\theta)}{2s}}.$ (4.27)
Combining (4.4) and (4.5), we deduce
$III_{1}\leq$ $C(C_{0}\varepsilon)^{2}\int_{0}^{t}(t-\tau)^{-\frac{1}{4s}}(1+\tau)^{-\frac{\sigma}{2s}}(1+\tau)^{-\frac{(\sigma+1)(1-\theta)}{2s}}d\tau.$ (4.28)
Given that $0<\theta<\frac{1}{9}$, it follows that
$s<\frac{5}{12}+\sigma\leq\frac{17}{18}\sigma+\frac{8}{18}<\Big(1-\frac{\theta}{2}\Big)\sigma+\frac{1}{2}(1-\theta),$
which implies
$-\frac{\sigma}{2s}-\frac{(\sigma+1)(1-\theta)}{2s}<-1.$
Therefore, by applying Lemma 2.8 with $\alpha<1$ and $\beta>1$, we obtain
$III_{1}$ $\leq C(C_{0}\varepsilon)^{2}(1+t)^{-\frac{1}{4s}}\leq C(C_{0}\varepsilon)^{2}(1+t)^{-\frac{\sigma}{2s}}.$ (4.29)
By Lemma 2.7, Lemma 2.1 and interpolation (4.27), we can derive
$III_{2}$ $\leq C\int_{0}^{t}(t-\tau)^{-\frac{1}{4s}}\|u_{2}(\tau)\|_{L^{2}}^{\frac{1}{2}}\|\partial_{2}u_{2}(\tau)\|_{L^{2}}^{\frac{1}{2}}\|\partial_{2}^{2}u_{1}(\tau)\|_{L^{2}}d\tau$
$\leq C(C_{0}\varepsilon)^{2}\int_{0}^{t}(t-\tau)^{-\frac{1}{4s}}(1+\tau)^{-\frac{6\sigma+5}{12s}}(1+\tau)^{-\frac{\sigma(1-\theta)}{2s}}d\tau.$ (4.30)
By direct calculation and the range of $s$, we can obtain
$III_{2}\leq$ $C(C_{0}\varepsilon)^{2}\int_{0}^{t}(t-\tau)^{-\frac{1}{4s}}(1+\tau)^{-\frac{12\sigma+5}{12s}}(1+\tau)^{\frac{\sigma\theta}{2s}}d\tau$
$\leq$ $C(C_{0}\varepsilon)^{2}\int_{0}^{t}(t-\tau)^{-\frac{1}{4s}}(1+\tau)^{-1}(1+\tau)^{\frac{\sigma\theta}{2s}}d\tau.$
By using Lemma 2.8 and combining the range of $\theta$ ($0<\theta<\frac{1-2\sigma}{2\sigma}$), we have
$III_{2}\leq$ $C(C_{0}\varepsilon)^{2}\int_{0}^{t}(t-\tau)^{-\frac{1}{4s}}(1+\tau)^{-1}(1+\tau)^{\frac{\sigma\theta}{2s}}d\tau$
$\leq$ $C(C_{0}\varepsilon)^{2}(1+t)^{\frac{2\sigma\theta-1}{4s}}\leq C(C_{0}\varepsilon)^{2}(1+t)^{-\frac{\sigma}{2s}}.$ (4.31)
As a direct consequence of (4.29) and (4.31), we obtain
$III_{1}+III_{2}\leq C(C_{0}\varepsilon)^{2}(1+t)^{-\frac{\sigma}{2s}}.$ (4.32)
Combining (4.25) with (4.32), we deduce
$\|\partial_{2}u_{1}(t)\|_{L^{2}}\leq\left(\frac{C_{0}}{4}+4CC_{0}^{2}\varepsilon\right)\varepsilon(1+t)^{-\frac{\sigma}{2s}}.$
By choosing $\varepsilon>0$ sufficiently small, the bootstrap bound (4.10) follows.
4.6 The decay of $\partial_{1}u_{1}$
By Duhamel’s principle, we have
$\|\partial_{1}u_{1}(t)\|_{L^{2}}\leq\|e^{\nu\Lambda_{1}^{2s}t}\partial_{1}u_{0,1}\|_{L^{2}}+C\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}(u\cdot\nabla u)(\tau)\|_{L^{2}}d\tau.\vskip-5.69046pt$
For the linear term involving initial data $\partial_{1}u_{0,1}$, we can obtain the decay estimate through Lemma 2.7:
$\|e^{\nu\Lambda_{1}^{2s}t}\partial_{1}u_{0,1}\|_{L^{2}}$ $\leq C(1+\nu t)^{-\frac{\sigma+1}{2s}}\left(\|\Lambda_{1}^{-\sigma}u_{0,1}\|_{L^{2}}+\|\partial_{1}u_{0,1}\|_{L^{2}}\right)$
$\leq\frac{C_{0}}{4}\varepsilon(1+\nu t)^{-\frac{\sigma+1}{2s}}.\vskip-5.69046pt$ (4.33)
Similar as the process deal with $u_{1}$, we only consider the nonlinear term contain $u\cdot\nabla u_{1}$. Lemma 2.3 yields
$\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}(u\cdot\nabla u_{1})(\tau)\|_{L^{2}}d\tau$
$\leq$ $C\int_{0}^{t}(t-\tau)^{-\frac{3}{4s}}\big\|\|(u\cdot\nabla u_{1})(\tau)\|_{L_{x_{1}}^{1}}\big\|_{L_{x_{2}}^{2}}d\tau$
$\leq$ $C\int_{0}^{t}(t-\tau)^{-\frac{3}{4s}}\|u_{1}(\tau)\|_{L^{2}}^{\frac{1}{2}}\|\partial_{2}u_{1}(\tau)\|_{L^{2}}^{\frac{1}{2}}\|\partial_{1}u_{1}(\tau)\|_{L^{2}}d\tau$
$+C\int_{0}^{t}(t-\tau)^{-\frac{3}{4s}}\|u_{2}(\tau)\|_{L^{2}}^{\frac{1}{2}}\|\partial_{2}u_{2}(\tau)\|_{L^{2}}^{\frac{1}{2}}\|\partial_{2}u_{1}(\tau)\|_{L^{2}}d\tau.\vskip-5.69046pt$
Since $\partial_{1}u_{1}=-\partial_{2}u_{2}$, and substituting (4.2), (4.3), (4.4) and (4.5), we have
$\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}(u\cdot\nabla u_{1})(\tau)\|_{L^{2}}d\tau$
$\leq$ $C\int_{0}^{t}(t-\tau)^{-\frac{3}{4s}}\|u_{1}(\tau)\|_{L^{2}}^{\frac{1}{2}}\|\partial_{2}u_{1}(\tau)\|_{L^{2}}^{\frac{1}{2}}\|\partial_{1}u_{1}(\tau)\|_{L^{2}}d\tau$
$+C\int_{0}^{t}(t-\tau)^{-\frac{3}{4s}}\|u_{2}(\tau)\|_{L^{2}}^{\frac{1}{2}}\|\partial_{1}u_{1}(\tau)\|_{L^{2}}^{\frac{1}{2}}\|\partial_{2}u_{1}(\tau)\|_{L^{2}}d\tau$
$\leq$ $CC_{0}^{2}\varepsilon^{2}\int_{0}^{t}(t-\tau)^{-\frac{3}{4s}}(\tau+1)^{-\frac{2\sigma+1}{2s}}d\tau$
$+CC_{0}^{2}\varepsilon^{2}\int_{0}^{t}(t-\tau)^{-\frac{3}{4s}}(\tau+1)^{-\frac{12\sigma+5}{12s}}d\tau.\vskip-5.69046pt$
Then applying Lemma 2.8 with $\alpha<1$, $\beta>1$, we can deduce
$\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}(u\cdot\nabla u_{1})(\tau)\|_{L^{2}}d\tau\leq CC_{0}^{2}\varepsilon^{2}(t+1)^{-\frac{\sigma+1}{2s}}.\vskip-5.69046pt$ (4.34)
Combining (4.33) and (4.34), we can derive (4.11) by taking $\varepsilon$ sufficiently small.
4.7 The decay of $\partial_{1}u_{2}$
By acting $\partial_{1}$ on (4.21), we obtain
$\partial_{t}\partial_{1}u_{2}+\nu\Lambda^{2s}\partial_{1}u_{2}=-\partial_{1}^{2}\Delta^{-1}\nabla\times\mbox{div}\ (u\otimes u).$
Using Duhamel’s principle and taking the $L^{2}$-norm, we have
$\|\partial_{1}u_{2}(t)\|_{L^{2}}\leq\|e^{\nu\Lambda_{1}^{2s}t}\partial_{1}u_{0,2}\|_{L^{2}}+\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{2}\Delta^{-1}\nabla\times(u\cdot\nabla u)(\tau)\|_{L^{2}}d\tau.$
The linear term decays as
$\|e^{\nu\Lambda_{1}^{2s}t}\partial_{1}u_{0,2}\|_{L^{2}}$ $\leq C(1+\nu t)^{-\frac{\sigma+2}{2s}}\big(\|\Lambda_{1}^{-\sigma}\psi_{0}\|_{L^{2}}+\|\partial_{1}u_{0,2}\|_{L^{2}}\big)$
$\leq C\varepsilon(1+\nu t)^{-\frac{\sigma+2}{2s}}\leq\frac{C_{0}}{4}\varepsilon(1+\nu t)^{-\frac{2\sigma}{s}}.$
We devide the nonlinear term into two parts:
$\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{2}\Delta^{-1}\nabla\times(u\cdot\nabla u)(\tau)\|_{L^{2}}d\tau$
$\leq$ $\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{2}\Delta^{-1}\nabla\times\partial_{1}(u_{1}u)(\tau)\|_{L^{2}}d\tau+\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{2}\Delta^{-1}\nabla\times\partial_{2}(u_{2}u)(\tau)\|_{L^{2}}d\tau$
$:=$ $V_{1}+V_{2}.$
We write $V_{1}$ in the form of a multiplier:
$V_{1}\leq C\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\Lambda^{-1}\Lambda_{1}^{3}(u_{1}u)(\tau)\|_{L^{2}}d\tau.$
Taking out $\Lambda_{1}^{4\sigma}$, and applying Lemma 2.7, we obtain
$V_{1}\leq$ $C\int_{0}^{t-1}(t-\tau)^{-\frac{2\sigma}{s}}\|\Lambda^{-1}\Lambda_{1}^{3-4\sigma}(u_{1}u)\|_{L^{2}}d\tau+C\int_{t-1}^{t}\|\partial_{1}^{2}(u_{1}u)\|_{L^{2}}d\tau.$
We bound $\Lambda_{1}^{2-4\sigma}$ by $\Lambda^{2-4\sigma}$, and then use Lemma 2.4 to get
$V_{1}\leq$ $C\int_{0}^{t-1}(t-\tau)^{-\frac{2\sigma}{s}}\|\Lambda^{1-4\sigma}\Lambda_{1}(u_{1}u)\|_{L^{2}}d\tau+C\int_{t-1}^{t}\|\partial_{1}^{2}(u_{1}u)\|_{L^{2}}d\tau$
$\leq$ $C\int_{0}^{t-1}(t-\tau)^{-\frac{2\sigma}{s}}\|\Lambda^{1-4\sigma}\big((\Lambda_{1}u)u\big)\|_{L^{2}}d\tau+C\int_{t-1}^{t}\|\partial_{1}^{2}(u_{1}u)\|_{L^{2}}d\tau$
$:=$ $V_{11}+V_{12}.$
Since $\frac{1}{3}<\sigma<\frac{1}{2}$, by Lemma 2.3, interpolation and Lemma 2.8, we derive
$V_{11}\leq$ $C\int_{0}^{t-1}(t-\tau)^{-\frac{2\sigma}{s}}\|\Lambda^{2-4\sigma}u\|_{L^{2}}\|\partial_{1}u\|_{L^{2}}d\tau$
$\leq$ $C\int_{0}^{t-1}(t-\tau)^{-\frac{2\sigma}{s}}\|\nabla u\|_{L^{2}}^{2-4\sigma}\|u\|_{L^{2}}^{4\sigma-1}\|\partial_{1}u\|_{L^{2}}d\tau$
$\leq$ $CC_{0}^{2}\varepsilon^{2}\int_{0}^{t-1}(t-\tau)^{-\frac{2\sigma}{s}}(1+\tau)^{-\frac{2\sigma+1}{2s}}d\tau$
$\leq$ $C(C_{0}\varepsilon)^{2}(1+t)^{-\frac{2\sigma}{s}}.$ (4.35)
It follows from Lemma 2.2 that
$V_{12}\leq$ $C\int_{t-1}^{t}\big(\|\partial_{1}u\partial_{1}u\|_{L^{2}}+\|u\partial_{1}^{2}u\|_{L^{2}}\big)d\tau$
$\leq$ $C\int_{t-1}^{t}\big(\|\partial_{1}u\|_{L^{2}}\|\partial_{1}^{2}u\|_{L^{2}}^{\frac{1}{2}}\|\partial_{1}\partial_{2}u\|_{L^{2}}^{\frac{1}{2}}+\|u\|_{L^{2}}^{\frac{1}{2}}\|\partial_{1}u\|_{L^{2}}^{\frac{1}{2}}\|\partial_{1}^{2}u\|_{L^{2}}^{\frac{1}{2}}\|\partial_{1}^{2}\partial_{2}u\|_{L^{2}}^{\frac{1}{2}}\big)d\tau.$ (4.36)
Then we use interpolation to obtain
$\|\partial_{1}^{2}u\|_{L^{2}}\|\partial_{1}\partial_{2}u\|_{L^{2}}\leq\|\partial_{1}u\|_{L^{2}}(\|\partial_{1}^{3}u\|_{L^{2}}+\|\partial_{1}\partial_{2}^{2}u\|_{L^{2}})\leq C\varepsilon(1+t)^{-\frac{\sigma+1}{2s}},$
$\|\partial_{1}^{2}u\|_{L^{2}}\|\partial_{1}^{2}\partial_{2}u\|_{L^{2}}\leq\|\partial_{1}u\|_{L^{2}}(\|\partial_{1}^{3}u\|_{L^{2}}+\|\partial_{1}^{3}\partial_{2}^{2}u\|_{L^{2}})\leq C\varepsilon(1+t)^{-\frac{\sigma+1}{2s}}.$ (4.37)
By substituting (4.2), (4.5) and (4.37) into (4.36), we have
$V_{12}\leq CC_{0}^{2}\varepsilon^{2}\int_{t-1}^{t}(1+\tau)^{-\frac{3\sigma+3}{4s}}+(1+\tau)^{-\frac{3\sigma+2}{4s}}d\tau\leq C(C_{0}\varepsilon)^{2}(1+t)^{-\frac{2\sigma}{s}}.$ (4.38)
Applying the boundness of Riesz operator $\Delta^{-1}\nabla\times\partial_{2}$ to $V_{2}$, we can obtain
$V_{2}\leq C\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{2}(u_{2}u)(\tau)\|_{L^{2}}d\tau.$
Using Lemma 2.7, we can deduce
$V_{2}\leq$ $C\int_{0}^{t-1}(t-\tau)^{-\frac{1}{s}}\|u_{2}u\|_{L^{2}}d\tau+C\int_{t-1}^{t}\|\partial_{1}^{2}(u_{2}u)\|_{L^{2}}d\tau$
$:=$ $V_{21}+V_{22}.$
Since $t-\tau\geq 1$ when $\tau$ varies in (0,t1)(0,t-1), we have
$V_{21}\leq C\int_{0}^{t-1}(t-\tau)^{-\frac{2\sigma}{s}}\|u_{2}u\|_{L^{2}}d\tau.$
By Lemma 2.2 and the decay rate of (4.2), (4.3) and (4.5), we have
$V_{21}\leq$ $C\int_{0}^{t-1}(t-\tau)^{-\frac{2\sigma}{s}}\|u_{2}\|_{L^{2}}^{\frac{1}{2}}\|\partial_{2}u_{2}\|_{L^{2}}^{\frac{1}{2}}\|u\|_{L^{2}}^{\frac{1}{2}}\|\partial_{1}u\|_{L^{2}}^{\frac{1}{2}}d\tau$
$\leq$ $CC_{0}^{2}\varepsilon^{2}\int_{0}^{t-1}(t-\tau)^{-\frac{2\sigma}{s}}(1+\tau)^{-\frac{3\sigma+2}{3s}}d\tau.$
Using Lemma 2.8 with $\alpha\geq 1$ and $\beta\geq\alpha$, we have
$V_{21}\leq CC_{0}^{2}\varepsilon^{2}\int_{0}^{t-1}(t-\tau)^{-\frac{2\sigma}{s}}(1+\tau)^{-\frac{3\sigma+2}{3s}}d\tau\leq$ $C(C_{0}\varepsilon)^{2}(1+t)^{-\frac{2\sigma}{s}}.$ (4.39)
By direct calculation,
$V_{22}\leq$ $C\int_{t-1}^{t}\big(\|u\partial_{1}^{2}u\|_{L^{2}}+\|\partial_{1}u\partial_{1}u\|_{L^{2}}\big)d\tau.\vskip-5.69046pt$
By (4.36), (4.37) and (4.38), we obtain
$V_{22}\leq C(C_{0}\varepsilon)^{2}(1+t)^{-\frac{2\sigma}{s}}.\vskip-5.69046pt$ (4.40)
Combining (4.35), (4.38), (4.39) and (4.40), we can obtain the bootstrap conclusion (4.12) by taking $\varepsilon$ sufficiently small.
5 Proof of Theorem 1.5
In this section, we complete the proof of Theorem 1.5. We again employ a bootstrap argument. Assume that there exist constants $C_{0}>0$ and $T>0$ such that, for all $t\in[0,T)$, the following estimates hold:
$\|\Lambda_{1}^{-\sigma}u(t)\|_{H^{k}}$ $\leq C_{0}\varepsilon,$ (5.1)
$\|[x_{2}]^{\frac{3\gamma+4}{7}}u(t)\|_{L^{2}}$ $\leq C_{0}\varepsilon(1+t)^{-\sigma/2s},$ (5.2)
$\|[x_{2}]^{\gamma}u_{2}(t)\|_{L^{2}}$ $\leq C_{0}\varepsilon(1+t)^{-(\sigma+1)/2s},$ (5.3)
$\|[x_{2}]^{\frac{5\gamma+2}{7}}\partial_{2}u_{1}(t)\|_{L^{2}}$ $\leq C_{0}\varepsilon(1+t)^{-\sigma/2s},$ (5.4)
$\|[x_{2}]^{\frac{3\gamma+4}{7}}\partial_{1}u(t)\|_{L^{2}}$ $\leq C_{0}\varepsilon(1+t)^{-(\sigma+1)/2s},$ (5.5)
$\|[x_{2}]^{\gamma}\partial_{1}u_{2}(t)\|_{L^{2}}$ $\leq C_{0}\varepsilon(1+t)^{-(2\sigma+1)/2s}.\vskip-8.5359pt$ (5.6)
Under the assumptions of Theorem 1.5, we will prove the following estimates in the following subsections:
$\|\Lambda_{1}^{-\sigma}u(t)\|_{H^{k}}$ $\leq 1/2C_{0}\varepsilon,$ (5.7)
$\|[x_{2}]^{\frac{3\gamma+4}{7}}u(t)\|_{L^{2}}$ $\leq 1/2C_{0}\varepsilon(1+t)^{-\sigma/2s},$ (5.8)
$\|[x_{2}]^{\gamma}u_{2}(t)\|_{L^{2}}$ $\leq 1/2C_{0}\varepsilon(1+t)^{-(\sigma+1)/2s},$ (5.9)
$\|[x_{2}]^{\frac{5\gamma+2}{7}}\partial_{2}u_{1}(t)\|_{L^{2}}$ $\leq 1/2C_{0}\varepsilon(1+t)^{-\sigma/2s},$ (5.10)
$\|[x_{2}]^{\frac{3\gamma+4}{7}}\partial_{1}u(t)\|_{L^{2}}$ $\leq 1/2C_{0}\varepsilon(1+t)^{-(\sigma+1)/2s},$ (5.11)
$\|[x_{2}]^{\gamma}\partial_{1}u_{2}(t)\|_{L^{2}}$ $\leq 1/2C_{0}\varepsilon(1+t)^{-(2\sigma+1)/2s}.$ (5.12)
Then using the standard bootstrap argument, we can conclude that $T=+\infty$ and derive the conclusions in Theorem 1.5.
5.1 Estimate for $\|\Lambda_{1}^{-\sigma}u(t)\|_{H^{k}}$
We splitting $\|\Lambda_{1}^{-\sigma}u(t)\|_{H^{k}}$ into two parts according to the equivalence ∥Λ_1^-σ u(t)∥_H^k ∼∥Λ_1^-σ u(t)∥_L^2 + ∑_m=1^2∥Λ_1^-σ ∂_m^k u∥_L^2, where the term $\sum_{m=1}^{2}\|\Lambda_{1}^{-\sigma}\partial_{m}^{k}u\|_{L^{2}}$ does not require coupling with decay estimates and has been completed in Subsection 4.2.
To complete the estimation of $\|\Lambda_{1}^{-\sigma}u\|_{L^{2}}$, let us recall (4.14) and (4.15) in Subsection 4.1, that is
$\|\Lambda_{1}^{-\sigma}u(t)\|_{L^{2}}^{2}+2\nu\int_{0}^{t}\|\Lambda_{1}^{s-\sigma}u(\tau)\|_{L^{2}}^{2}d\tau=\|\Lambda_{1}^{-\sigma}u_{0}\|_{L^{2}}^{2}+2\int_{0}^{t}M_{0}(\tau)d\tau,$
where
$M_{0}$ $\leq C\Big(\left\|\partial_{1}u_{1}\right\|_{L^{2}}^{\frac{1}{2}}\left\|u_{2}\right\|_{L^{2}}^{\sigma}\left\|\partial_{1}u_{2}\right\|_{L^{2}}^{\frac{1}{2}-\sigma}\left\|\partial_{2}u\right\|_{L^{2}}$
$\quad+\left\|\partial_{2}u_{1}\right\|_{L^{2}}^{\frac{1}{2}}\left\|u_{1}\right\|_{L^{2}}^{\sigma}\left\|\partial_{1}u_{1}\right\|_{L^{2}}^{\frac{1}{2}-\sigma}\left\|\partial_{1}u\right\|_{L^{2}}\Big)\left\|\Lambda_{1}^{-\sigma}u\right\|_{L^{2}}.$ (5.13)
By Lemma 2.9 and substituting (5.2), (5.3), (5.4), (5.5), (5.6) into (5.13), one can deduce
$|M_{0}|\leq CC_{0}^{3}\varepsilon^{3}\big((1+t)^{-\frac{-2\sigma^{2}+5\sigma+2}{4s}}+(1+t)^{-\frac{2\sigma+1}{2s}}\big).$
It can be checked that for $\frac{1}{3}<\sigma<\frac{1}{2}$, $s<\frac{1}{2}+\sigma<\frac{-2\sigma^{2}+5\sigma+2}{4}$, we have $\min\{\frac{-2\sigma^{2}+5\sigma+2}{4s},\frac{2\sigma+1}{2s}\}>1$ and
$\int_{0}^{t}M_{0}(\tau)d\tau\leq CC_{0}^{3}\varepsilon^{3}.$
Combining Lemma 2.9 and the initial condition (1.5), we derive from equation (4.14) that
$\|\Lambda_{1}^{-\sigma}u(t)\|_{L^{2}}^{2}$ $\leq\varepsilon^{2}+2CC_{0}^{3}\varepsilon^{3}.$
For $\varepsilon$ sufficiently small such that $2CC_{0}^{3}\varepsilon\leq\frac{1}{2}$, taking square roots of the above inequality, we obtain:
$\|\Lambda_{1}^{-\sigma}u(t)\|_{L^{2}}\leq\sqrt{\frac{3}{2}}\varepsilon\leq\frac{C_{0}}{2}\varepsilon.$
5.2 Weighted estimate of the pressure
The next step is to establish decay estimates for the solution. A major difficulty arises from the introduction of weighted norms: the Leray projection operator $\mathbb{P}$ is no longer directly bounded, since it involves Riesz transforms in a weighted setting. In particular, the weight $x_{2}]^{\eta}$ fails to be an $A_{2}$ when $\eta\geq\frac{1}{2}$, and therefore standard weighted boundedness results for Riesz transforms do not apply. Remarkably, the Riesz transform associated with the pressure term can still be eliminated under the weight $x_{2}]^{\eta}$. This is made possible by the special structural properties of the pressure, which allow us to bypass the lack of $A_{2}$ control. The precise formulation of this observation is given in the following lemma.
Lemma 5.1.
Let $\frac{1}{2}<\eta<\frac{3}{2}$. Assume that $p$ denotes the pressure associated with system (1.1), and let $\mathcal{E}=e^{\nu\Lambda_{1}^{2s}(t-\tau)}$. Then there exists a constant $C>0$, independent of $t$ and $\tau$, such that
$\|\mathcal{E}\,[x_{2}]^{\eta}\nabla p\|_{L^{2}}\leq C\,\|\mathcal{E}\,[x_{2}]^{\eta}(u\cdot\nabla u)\|_{L^{2}},$ (5.14)
and
$\|\mathcal{E}\,[x_{2}]^{\eta}\nabla\partial_{i}p\|_{L^{2}}\leq C\,\|\mathcal{E}\,[x_{2}]^{\eta}\partial_{i}(u\cdot\nabla u)\|_{L^{2}},\qquad i=1,2.$ (5.15)
Proof.
Let us prove (5.14) first. By integration by parts, we obtain
$\|\mathcal{E}[x_{2}]^{\eta}\nabla p\|_{L^{2}}^{2}=$ $\int_{\mathbb{R}^{2}}[x_{2}]^{2\eta}(\mathcal{E}\nabla p)\cdot(\mathcal{E}\nabla p)dx$
$=$ $\int_{\mathbb{R}^{2}}[x_{2}]^{2\eta}(\mathcal{E}p)\cdot(\mathcal{E}(-\Delta)p)dx+\eta(2\eta-1)\int_{\mathbb{R}^{2}}[x_{2}]^{2\eta-2}(\mathcal{E}p)^{2}dx.$ (5.16)
Notice that the pressure term in system (1.1) satisfies the elliptic equation
$\Delta p=\mbox{div}\ (u\cdot\nabla u).\vskip-5.69046pt$ (5.17)
Substituting (5.17) into (5.16), we obtain
$\|\mathcal{E}[x_{2}]^{\eta}\nabla p\|_{L^{2}}^{2}=\int_{\mathbb{R}^{2}}[x_{2}]^{2\eta}(\mathcal{E}p)(\mathcal{E}\mbox{div}\ (u\cdot\nabla u))dx+\eta(2\eta-1)\|[x_{2}]^{\eta-1}\mathcal{E}p\|_{L^{2}}^{2}.$ (5.18)
Applying integration by parts again, one can deduce
$\int_{\mathbb{R}^{2}}[x_{2}]^{2\eta}(\mathcal{E}p)(\mathcal{E}\mbox{div}\ (u\cdot\nabla u))dx=$ $-2\eta\int_{\mathbb{R}^{2}}[x_{2}]^{2\eta-1}(\mathcal{E}p)(\mathcal{E}(u\cdot\nabla u))dx$
$-\int_{\mathbb{R}^{2}}[x_{2}]^{2\eta}(\mathcal{E}\nabla p)\cdot(\mathcal{E}(u\cdot\nabla u))dx.$
Using Cauchy-Schwarz inequality, we get
$\left|\int_{\mathbb{R}^{2}}[x_{2}]^{2\eta}(\mathcal{E}p)(\mathcal{E}\mbox{div}\ (u\cdot\nabla u))dx\right|$
$\leq$ $2\eta\|[x_{2}]^{\eta-1}\mathcal{E}p\|_{L^{2}}\|[x_{2}]^{\eta}\mathcal{E}(u\cdot\nabla u)\|_{L^{2}}+\|[x_{2}]^{\eta}(\mathcal{E}\nabla p)\|_{L^{2}}\|[x_{2}]^{\eta}(\mathcal{E}(u\cdot\nabla u))\|_{L^{2}}$
$\leq$ $\eta\|[x_{2}]^{\eta-1}\mathcal{E}p\|_{L^{2}}^{2}+(\eta+\frac{1}{2})\|[x_{2}]^{\eta}\mathcal{E}(u\cdot\nabla u)\|_{L^{2}}^{2}+\frac{1}{2}\|[x_{2}]^{\eta}(\mathcal{E}\nabla p)\|_{L^{2}}^{2}.$ (5.19)
Combining (5.18) and (5.19), we can deduce
$\|\mathcal{E}[x_{2}]^{\eta}\nabla p\|_{L^{2}}^{2}\leq 4\eta^{2}\|[x_{2}]^{\eta-1}\mathcal{E}p\|_{L^{2}}^{2}+(2\eta+1)\|[x_{2}]^{\eta}\mathcal{E}(u\cdot\nabla u)\|_{L^{2}}^{2}.$ (5.20)
It follows from (5.17) that
$\|[x_{2}]^{\eta-1}\mathcal{E}p\|_{L^{2}}^{2}=\|[x_{2}]^{\eta-1}\mathcal{E}(-\Delta)^{-1}\mbox{div}\ \partial_{2}\int_{-\infty}^{x_{2}}(u\cdot\nabla u)(y_{2})dy_{2}\|_{L^{2}}^{2}.$
Since $\frac{1}{2}<\eta<\frac{3}{2}$, we can use Lemma 2.13 and 2.10 to get
$\|[x_{2}]^{\eta-1}\mathcal{E}p\|_{L^{2}}^{2}=$ $\|[x_{2}]^{\eta-1}\mathcal{E}(-\Delta)^{-1}\mbox{div}\ \partial_{2}\int_{-\infty}^{x_{2}}(u\cdot\nabla u)(y_{2})dy_{2}\|_{L^{2}}^{2}$
$\leq$ $C\|[x_{2}]^{\eta-1}\mathcal{E}\int_{-\infty}^{x_{2}}(u\cdot\nabla u)(y_{2})dy_{2}\|_{L^{2}}^{2}$
$\leq$ $C\|[x_{2}]^{\eta}\mathcal{E}(u\cdot\nabla u)\|_{L^{2}}^{2}.$ (5.21)
Combining (5.20) and (5.21), we complete the proof of (5.14). The proof of (5.15) follows analogously by replacing $p$ with $\partial_{i}p$ in the above argument. ∎
5.3 The decay of $u$ with weight
The weighted solution can be expressed as:
$[x_{2}]^{\frac{3\gamma+4}{7}}u(t)=$ $e^{\nu\Lambda_{1}^{2s}t}[x_{2}]^{\frac{3\gamma+4}{7}}u_{0}+\int_{0}^{t}e^{\nu\Lambda_{1}^{2s}(t-\tau)}[x_{2}]^{\frac{3\gamma+4}{7}}(u\cdot\nabla u)(\tau)d\tau$
$+\int_{0}^{t}e^{\nu\Lambda_{1}^{2s}(t-\tau)}[x_{2}]^{\frac{3\gamma+4}{7}}\nabla p~d\tau.$
For $u_{1}$, taking the $L^{2}$-norm and using Lemma 5.1:
$\|[x_{2}]^{\frac{3\gamma+4}{7}}u(t)\|_{L^{2}}\leq\|e^{\nu\Lambda_{1}^{2s}t}[x_{2}]^{\frac{3\gamma+4}{7}}u_{0}\|_{L^{2}}+C\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}[x_{2}]^{\frac{3\gamma+4}{7}}(u\cdot\nabla u)(\tau)\|_{L^{2}}d\tau.$ (5.22)
The linear term admits a straightforward decay estimate:
$\|e^{\nu\Lambda_{1}^{2s}t}[x_{2}]^{\frac{3\gamma+4}{7}}u_{0}\|_{L^{2}}$ $\leq C(1+\nu t)^{-\frac{\sigma}{2s}}\left(\|\Lambda_{1}^{-\sigma}[x_{2}]^{\frac{3\gamma+4}{7}}u_{0}\|_{L^{2}}+\|[x_{2}]^{\frac{3\gamma+4}{7}}u_{0}\|_{L^{2}}\right)$
$\leq C\varepsilon(1+\nu t)^{-\frac{\sigma}{2s}}\leq\frac{C_{0}}{4}\varepsilon(1+\nu t)^{-\frac{\sigma}{2s}},$
where we have chosen $C_{0}$ satisfying $C_{0}\geq 4C$. For the nonlinear terms, we can deduce
$\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}[x_{2}]^{\frac{3\gamma+4}{7}}(u\cdot\nabla u)(\tau)\|_{L^{2}}d\tau$
$\leq$ $\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}[x_{2}]^{\frac{3\gamma+4}{7}}(u_{1}\partial_{1}u_{1})(\tau)\|_{L^{2}}d\tau+\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}[x_{2}]^{\frac{3\gamma+4}{7}}(u_{2}\partial_{2}u_{1})(\tau)\|_{L^{2}}d\tau$
$+\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}[x_{2}]^{\frac{3\gamma+4}{7}}(u_{1}\partial_{1}u_{2})(\tau)\|_{L^{2}}d\tau+\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}[x_{2}]^{\frac{3\gamma+4}{7}}(u_{2}\partial_{2}u_{2})(\tau)\|_{L^{2}}d\tau$
$:=$ $I^{w}_{1}+I^{w}_{2}+I^{w}_{3}+I^{w}_{4}.$
We first deal with $I^{w}_{1}$, $I^{w}_{3}$ and $I^{w}_{4}$. It follows from Lemma 2.7 and divergence-free condition that
$I^{w}_{1}+I^{w}_{3}+I^{w}_{4}\leq$ $C\int_{0}^{t}(t-\tau)^{-\frac{1}{4s}}\big\|[x_{2}]^{\frac{3\gamma+4}{7}}\|(u_{1}\partial_{1}u_{1})(\tau)\|_{L^{1}_{x_{1}}}\big\|_{L^{2}_{x_{2}}}d\tau$
$+C\int_{0}^{t}(t-\tau)^{-\frac{1}{4s}}\big\|[x_{2}]^{\frac{3\gamma+4}{7}}\|(u_{1}\partial_{1}u_{2})(\tau)\|_{L^{1}_{x_{1}}}\big\|_{L^{2}_{x_{2}}}d\tau$
$+C\int_{0}^{t}(t-\tau)^{-\frac{1}{4s}}\big\|[x_{2}]^{\frac{3\gamma+4}{7}}\|(u_{2}\partial_{1}u_{1})(\tau)\|_{L^{1}_{x_{1}}}\big\|_{L^{2}_{x_{2}}}d\tau.$ (5.23)
Applying Lemma 2.1, we derive
$\|[x_{2}]^{\frac{3\gamma+4}{7}}(u_{1}\partial_{1}u_{1})(\tau)\|_{L^{1}_{x_{1}}L^{2}_{x_{2}}}\leq$ $C\|u_{1}\|_{L^{2}}^{\frac{1}{2}}\|\partial_{2}u_{1}\|_{L^{2}}^{\frac{1}{2}}\|[x_{2}]^{\frac{3\gamma+4}{7}}\partial_{1}u_{1}\|_{L^{2}}.$ (5.24)
$\|[x_{2}]^{\frac{3\gamma+4}{7}}(u_{1}\partial_{1}u_{2})(\tau)\|_{L^{1}_{x_{1}}L^{2}_{x_{2}}}\leq$ $C\|u_{1}\|_{L^{2}}^{\frac{1}{2}}\|\partial_{2}u_{1}\|_{L^{2}}^{\frac{1}{2}}\|[x_{2}]^{\frac{3\gamma+4}{7}}\partial_{1}u_{2}\|_{L^{2}}.$ (5.25)
$\|[x_{2}]^{\frac{3\gamma+4}{7}}(u_{2}\partial_{1}u_{1})(\tau)\|_{L^{1}_{x_{1}}L^{2}_{x_{2}}}\leq$ $C\|u_{2}\|_{L^{2}}^{\frac{1}{2}}\|\partial_{2}u_{2}\|_{L^{2}}^{\frac{1}{2}}\|[x_{2}]^{\frac{3\gamma+4}{7}}\partial_{1}u_{1}\|_{L^{2}}.$ (5.26)
Similarly, we apply Lemma 2.7 to $I^{w}_{2}$ and obtain
$I^{w}_{2}\leq C\int_{0}^{t}(t-\tau)^{-\frac{1}{4s}}\big\|[x_{2}]^{\frac{3\gamma+4}{7}}\|(u_{2}\partial_{2}u_{1})(\tau)\|_{L^{1}_{x_{1}}}\big\|_{L^{2}_{x_{2}}}.$ (5.27)
From the divergence-free condition of $u$, it follows that
$\|[x_{2}]^{\frac{3\gamma+4}{7}}(u_{2}\partial_{2}u_{1})(\tau)\|_{L^{1}_{x_{1}}L^{2}_{x_{2}}}\leq$ $C\|[x_{2}]^{\frac{-2\gamma+2}{7}}u_{2}\|_{L^{2}_{x_{1}}L^{\infty}_{x_{2}}}\|[x_{2}]^{\frac{5\gamma+2}{7}}(\partial_{2}u_{1})\|_{L^{2}}.$ (5.28)
By using Lemma 2.11 and the range of $\gamma$ ($0<\gamma<\frac{3}{10}$) that
$\|[x_{2}]^{\frac{-2\gamma+2}{7}}u_{2}\|_{L^{2}_{x_{1}}L^{\infty}_{x_{2}}}\leq$ $C\|[x_{2}]^{\frac{-2\gamma+2}{7}-\frac{1}{2}}u_{2}\|_{L^{2}}+C\|[x_{2}]^{\gamma}u_{2}\|_{L^{2}}^{\frac{1}{2}}\|[x_{2}]^{\frac{-11\gamma+4}{7}}\partial_{2}u_{2}\|_{L^{2}}^{\frac{1}{2}}$
$\leq$ $C\|[x_{2}]^{\gamma}u_{2}\|_{L^{2}}+C\|[x_{2}]^{\gamma}u_{2}\|_{L^{2}}^{\frac{1}{2}}\|[x_{2}]^{\frac{3\gamma+4}{7}}\partial_{1}u_{1}\|_{L^{2}}^{\frac{1}{2}}.$ (5.29)
Combining (5.23)–(5.29) and substituting the decay rates in (5.2)–(5.6), we have
$\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}[x_{2}]^{\frac{3\gamma+4}{7}}(u\cdot\nabla u)(\tau)\|_{L^{2}}d\tau\leq CC_{0}^{2}\varepsilon^{2}\int_{0}^{t}(t-\tau)^{-\frac{1}{4s}}(\tau+1)^{-\frac{2\sigma+1}{2s}}d\tau.$
Applying Lemma 2.8, we obtain
$\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}[x_{2}]^{\frac{3\gamma+4}{7}}(u\cdot\nabla u)(\tau)\|_{L^{2}}d\tau\leq CC_{0}^{2}\varepsilon^{2}(t+1)^{-\frac{\sigma}{2s}}.$ (5.30)
Combining (5.22) and (5.30), and taking $\varepsilon$ sufficiently small, one can derive the bootstrap conclusion (5.8).
5.4 The decay of $u_{2}$ with weight
By multiplying $x_{2}]^{\gamma}$ on both side of (4.21), and then applying Duhamel’s principle, we obtain
$\|[x_{2}]^{\gamma}u_{2}(t)\|_{L^{2}}\leq\|e^{\nu\Lambda_{1}^{2s}t}\partial_{1}[x_{2}]^{\gamma}\psi_{0}\|_{L^{2}}+\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}[x_{2}]^{\gamma}\Delta^{-1}\nabla\times(u\cdot\nabla u)(\tau)\|_{L^{2}}d\tau.$
The linear term decays as
$\|e^{\nu\Lambda_{1}^{2s}t}\partial_{1}[x_{2}]^{\gamma}\psi_{0}\|_{L^{2}}$ $\leq C(1+\nu t)^{-\frac{\sigma+1}{2s}}\big(\|\Lambda_{1}^{-\sigma}[x_{2}]^{\gamma}\psi_{0}\|_{L^{2}}+\|[x_{2}]^{\gamma}u_{0,2}\|_{L^{2}}\big)$
$\leq C\varepsilon(1+\nu t)^{-\frac{\sigma+1}{2s}}\leq\frac{C_{0}}{4}\varepsilon(1+\nu t)^{-\frac{\sigma+1}{2s}}.$
For the nonlinear term, we split it into four parts:
$\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}[x_{2}]^{\gamma}\Delta^{-1}\nabla\times(u\cdot\nabla u)(\tau)\|_{L^{2}}d\tau$
$=$ $\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}[x_{2}]^{\gamma}\Delta^{-1}\nabla\times\mbox{div}\ (u\otimes u)(\tau)\|_{L^{2}}d\tau$
$\leq$ $\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}[x_{2}]^{\gamma}\partial_{1}\partial_{2}\Delta^{-1}(u_{1}u_{1})(\tau)\|_{L^{2}}d\tau$
$+\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}[x_{2}]^{\gamma}\partial_{1}^{2}\Delta^{-1}(u_{1}u_{2})(\tau)\|_{L^{2}}d\tau$
$+\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}[x_{2}]^{\gamma}\Delta^{-1}\partial_{2}^{2}(u_{2}u_{1})(\tau)\|_{L^{2}}d\tau$
$+\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}[x_{2}]^{\gamma}\Delta^{-1}\partial_{1}\partial_{2}(u_{2}u_{2})(\tau)\|_{L^{2}}d\tau$
$:=$ $II_{1}^{w}+II_{2}^{w}+II_{3}^{w}+II_{4}^{w}.$ (5.31)
By fundamental theorem of calculus, we have
$\|[x_{2}]^{\gamma}e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{2}\partial_{2}\Delta^{-1}(u_{1}u_{1})\|_{L^{2}}$
$\leq$ $\|[x_{2}]^{\gamma}e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{2}\partial_{2}\Delta^{-1}\partial_{2}\int_{-\infty}^{x_{2}}(u_{1}u_{1})(y_{2})dy_{2}\|_{L^{2}}.$ (5.32)
Noting that $x_{2}]^{\gamma}~(0<\gamma<\frac{3}{10})$ is an $A_{2}$ weight (the proof can be found in Appendix A, see the proof of Lemma 2.13), then we can employ Lemma 2.13 to obtain
$\left\|[x_{2}]^{\gamma}\partial_{2}^{2}\Delta^{-1}e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{2}\int_{-\infty}^{x_{2}}(u_{1}u_{1})(y_{2})dy_{2}\right\|_{L^{2}}$
$\leq$ $C\left\|[x_{2}]^{\gamma}e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{2}\int_{-\infty}^{x_{2}}(u_{1}u_{1})(y_{2})dy_{2}\right\|_{L^{2}}.$ (5.33)
Applying Lemma 2.10, we have
$\left\|[x_{2}]^{\gamma}e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{2}\int_{-\infty}^{x_{2}}(u_{1}u_{1})(y_{2})dy_{2}\right\|_{L^{2}}\leq C\left\|[x_{2}]^{\gamma+1}e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{2}(u_{1}u_{1})\right\|_{L^{2}}.$ (5.34)
Combining (5.32), (5.33) and (5.34), and using Lemma 2.7, we have
$II_{1}^{w}\leq$ $\int_{0}^{t}(t-\tau)^{-\frac{3}{4s}}\big\|[x_{2}]^{\gamma+1}\|\partial_{1}(u_{1}u_{1})\|_{L^{1}_{x_{1}}}\big\|_{L_{x_{2}}^{2}}d\tau$
$\leq$ $C\int_{0}^{t}(t-\tau)^{-\frac{3}{4s}}\big\|[x_{2}]^{\gamma+1}\|u_{1}\partial_{1}u_{1}\|_{L^{1}_{x_{1}}}\big\|_{L_{x_{2}}^{2}}d\tau.$
By Lemma 2.1, we obtain
$\int_{0}^{t}(t-\tau)^{-\frac{3}{4s}}\big\|[x_{2}]^{\gamma+1}\|u_{1}\partial_{1}u_{1}\|_{L^{1}_{x_{1}}}\big\|_{L_{x_{2}}^{2}}d\tau$
$\leq$ $\int_{0}^{t}(t-\tau)^{-\frac{3}{4s}}\|[x_{2}]^{\frac{4\gamma+3}{7}}u_{1}\|_{L^{2}_{x_{1}}L^{\infty}_{x_{2}}}\|[x_{2}]^{\frac{3\gamma+4}{7}}\partial_{1}u_{1}\|_{L^{2}}d\tau.$ (5.35)
It follows from Lemma 2.11 that
$\|[x_{2}]^{\frac{4\gamma+3}{7}}u_{1}\|_{L^{2}_{x_{1}}L^{\infty}_{x_{2}}}\leq$ $C\|[x_{2}]^{\frac{4\gamma+3}{7}-\frac{1}{2}}u_{1}\|_{L^{2}}+C\|[x_{2}]^{\frac{3\gamma+4}{7}}u_{1}\|_{L^{2}}^{\frac{1}{2}}\|[x_{2}]^{\frac{5\gamma+2}{7}}\partial_{2}u_{1}\|_{L^{2}}^{\frac{1}{2}}$
$\leq$ $C\|[x_{2}]^{\frac{4\gamma+3}{7}}u_{1}\|_{L^{2}}+C\|[x_{2}]^{\frac{3\gamma+4}{7}}u_{1}\|_{L^{2}}^{\frac{1}{2}}\|[x_{2}]^{\frac{5\gamma+2}{7}}\partial_{2}u_{1}\|_{L^{2}}^{\frac{1}{2}}.$ (5.36)
Substitute (5.36) into (5.4), we obtain
$\int_{0}^{t}(t-\tau)^{-\frac{3}{4s}}\big\|[x_{2}]^{\gamma+1}\|u_{1}\partial_{1}u_{1}\|_{L^{1}_{x_{1}}}\big\|_{L_{x_{2}}^{2}}d\tau$
$\leq$ $C\int_{0}^{t}(t-\tau)^{-\frac{3}{4s}}\Big(\|[x_{2}]^{\frac{4\gamma+3}{7}}u_{1}\|_{L^{2}}+\|[x_{2}]^{\frac{3\gamma+4}{7}}u_{1}\|_{L^{2}}^{\frac{1}{2}}\|[x_{2}]^{\frac{5\gamma+2}{7}}\partial_{2}u_{1}\|_{L^{2}}^{\frac{1}{2}}\Big)\|[x_{2}]^{\frac{3\gamma+4}{7}}\partial_{1}u_{1}\|_{L^{2}}d\tau.$
Since $0<\gamma<\frac{3}{10}$, we can deduce $\|[x_{2}]^{\frac{4\gamma+3}{7}}u_{1}\|_{L^{2}}\leq\|[x_{2}]^{\frac{3\gamma+4}{7}}u_{1}\|_{L^{2}}$. By substituting the decay rates in (5.2), (5.3), (5.4), (5.5), (5.6), one can derive
$II_{1}^{w}\leq$ $C\int_{0}^{t}(t-\tau)^{-\frac{3}{4s}}\Big(\|[x_{2}]^{\frac{3\gamma+4}{7}}u_{1}\|_{L^{2}}+\|[x_{2}]^{\frac{3\gamma+4}{7}}u_{1}\|_{L^{2}}^{\frac{1}{2}}\|[x_{2}]^{\frac{5\gamma+2}{7}}\partial_{2}u_{1}\|_{L^{2}}^{\frac{1}{2}}\Big)\|[x_{2}]^{\frac{3\gamma+4}{7}}\partial_{1}u_{1}\|_{L^{2}}d\tau$
$\leq$ $C\int_{0}^{t}(t-\tau)^{-\frac{3}{4s}}(1+\tau)^{-\frac{2\sigma+1}{2s}}d\tau.$
Then by using Lemma 2.8, we can conclude
$II_{1}^{w}\leq CC_{0}^{2}\varepsilon^{2}(1+t)^{-\frac{3}{4s}}\leq CC_{0}^{2}\varepsilon^{2}(1+t)^{-\frac{\sigma+1}{2s}}.$
Next, we turn to the term $II_{2}^{w}$ in (5.31). First, we apply Lemma 2.13 to bound the Riesz operator $\partial_{1}^{2}\Delta^{-1}$, and yield
$II_{2}^{w}\leq C\int_{0}^{t}\|[x_{2}]^{\gamma}e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}(u_{1}u_{2})(\tau)\|_{L^{2}}d\tau.$
Then we use Lemma 2.7 to obtain
$\int_{0}^{t}\|[x_{2}]^{\gamma}e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}(u_{1}u_{2})(\tau)\|_{L^{2}}d\tau\leq C\int_{0}^{t}(t-\tau)^{-\frac{3}{4s}}\big\|[x_{2}]^{\gamma}\|(u_{1}u_{2})\|_{L_{x_{1}}^{1}}\big\|_{L_{x_{2}}^{2}}d\tau.$
Applying Lemma 2.1, we can deduce that
$\int_{0}^{t}(t-\tau)^{-\frac{3}{4s}}\big\|[x_{2}]^{\gamma}\|(u_{1}u_{2})\|_{L_{x_{1}}^{1}}\big\|_{L_{x_{2}}^{2}}d\tau\leq C\int_{0}^{t}(t-\tau)^{-\frac{3}{4s}}\|[x_{2}]^{\gamma}u_{2}\|_{L^{2}}\|u_{1}\|_{L^{2}}^{\frac{1}{2}}\|\partial_{2}u_{1}\|_{L^{2}}^{\frac{1}{2}}d\tau.$
It follows from Lemma 2.9 that
$\int_{0}^{t}(t-\tau)^{-\frac{3}{4s}}\|[x_{2}]^{\gamma}u_{2}\|_{L^{2}}\|u_{1}\|_{L^{2}}^{\frac{1}{2}}\|\partial_{2}u_{1}\|_{L^{2}}^{\frac{1}{2}}d\tau$
$\leq$ $C\int_{0}^{t}(t-\tau)^{-\frac{3}{4s}}\|[x_{2}]^{\gamma}u_{2}\|_{L^{2}}\|[x_{2}]^{\frac{3\gamma+4}{7}}u_{1}\|_{L^{2}}^{\frac{1}{2}}\|[x_{2}]^{\frac{5\gamma+2}{7}}\partial_{2}u_{1}\|_{L^{2}}^{\frac{1}{2}}d\tau.$
Substituting the decay rates given in (5.2)–(5.6) into the above expression, we obtain
$\int_{0}^{t}(t-\tau)^{-\frac{3}{4s}}\|[x_{2}]^{\gamma}u_{2}\|_{L^{2}}\|[x_{2}]^{\frac{3\gamma+4}{7}}u_{1}\|_{L^{2}}^{\frac{1}{2}}\|[x_{2}]^{\frac{5\gamma+2}{7}}\partial_{2}u_{1}\|_{L^{2}}^{\frac{1}{2}}d\tau$
$\leq$ $CC_{0}^{2}\varepsilon^{2}\int_{0}^{t}(t-\tau)^{-\frac{3}{4s}}(1+\tau)^{\frac{2\sigma+1}{2s}}d\tau.$
Summing up the estimates above and using Lemma 2.8, we conclude
$II_{2}^{w}\leq CC_{0}^{2}\varepsilon^{2}(1+t)^{-\frac{\sigma+1}{2s}}.$
Following a similar process with estimation of $II_{2}^{w}$, we can deduce
$II_{3}^{w}\leq$ $\int_{0}^{t}\|[x_{2}]^{\gamma}e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}\Delta^{-1}\partial_{2}^{2}(u_{2}u_{1})(\tau)\|_{L^{2}}d\tau$
$\leq$ $C\int_{0}^{t}\|[x_{2}]^{\gamma}e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}(u_{2}u_{1})(\tau)\|_{L^{2}}d\tau$
$\leq$ $C\int_{0}^{t}(t-\tau)^{-\frac{3}{4s}}\|[x_{2}]^{\gamma}(u_{2}u_{1})(\tau)\|_{L_{x_{1}}^{1}L_{x_{2}}^{2}}d\tau$
$\leq$ $C\int_{0}^{t}(t-\tau)^{-\frac{3}{4s}}\|[x_{2}]^{\gamma}u_{2}\|_{L^{2}}\|u_{1}\|_{L^{2}}^{\frac{1}{2}}\|\partial_{2}u_{1}\|_{L^{2}}^{\frac{1}{2}}d\tau$
$\leq$ $CC_{0}^{2}\varepsilon^{2}\int_{0}^{t}(t-\tau)^{-\frac{3}{4s}}(\tau+1)^{-\frac{2\sigma+1}{2s}}d\tau$
$\leq$ $CC_{0}^{2}\varepsilon^{2}(t+1)^{-\frac{\sigma+1}{2s}},$
and
$II_{4}^{w}\leq$ $\int_{0}^{t}\|[x_{2}]^{\gamma}e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}\Delta^{-1}\partial_{1}\partial_{2}(u_{2}u_{2})(\tau)\|_{L^{2}}d\tau$
$\leq$ $C\int_{0}^{t}\|[x_{2}]^{\gamma}e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}(u_{2}u_{2})(\tau)\|_{L^{2}}d\tau$
$\leq$ $C\int_{0}^{t}(t-\tau)^{-\frac{3}{4s}}\|[x_{2}]^{\gamma}(u_{2}u_{2})(\tau)\|_{L_{x_{1}}^{1}L_{x_{2}}^{2}}d\tau$
$\leq$ $C\int_{0}^{t}(t-\tau)^{-\frac{3}{4s}}\|[x_{2}]^{\gamma}u_{2}\|_{L^{2}}\|u_{2}\|_{L^{2}}^{\frac{1}{2}}\|\partial_{2}u_{2}\|_{L^{2}}^{\frac{1}{2}}d\tau$
$\leq$ $CC_{0}^{2}\varepsilon^{2}\int_{0}^{t}(t-\tau)^{-\frac{3}{4s}}(\tau+1)^{-\frac{2\sigma+2}{2s}}d\tau$
$\leq$ $CC_{0}^{2}\varepsilon^{2}(t+1)^{-\frac{\sigma+1}{2s}}.$
Combining the estimates above, we can obtain the bootstrap conclusion (5.9) by taking $\varepsilon$ sufficiently small.
5.5 The decay of $\partial_{2}u_{1}$ with weight
Applying Duhamel’s principle, we have
$\|[x_{2}]^{\frac{5\gamma+2}{7}}\partial_{2}u_{1}(t)\|_{L^{2}}\leq$ $\|[x_{2}]^{\frac{5\gamma+2}{7}}e^{\nu\Lambda_{1}^{2s}t}\partial_{2}u_{0,1}\|_{L^{2}}+\int_{0}^{t}\|[x_{2}]^{\frac{5\gamma+2}{7}}e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{2}(u\cdot\nabla u_{1})(\tau)\|_{L^{2}}d\tau$
$\quad+\int_{0}^{t}\|[x_{2}]^{\frac{5\gamma+2}{7}}e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}\partial_{2}p(\tau)\|_{L^{2}}d\tau.$
The range $0<\gamma<\frac{3}{10}$ ensures that $x_{2}]^{\frac{5\gamma+2}{7}}$ is an $A_{2}$ weight, which allows us to apply Lemma 2.13 to obtain the boundedness of the Riesz operator:
$\int_{0}^{t}\|[x_{2}]^{\frac{5\gamma+2}{7}}e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}\partial_{2}p(\tau)\|_{L^{2}}d\tau$
$=$ $\int_{0}^{t}\|[x_{2}]^{\frac{5\gamma+2}{7}}e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}\partial_{2}(-\Delta)^{-1}\mbox{div}\ (u\cdot\nabla u)(\tau)\|_{L^{2}}d\tau$
$\leq$ $C\int_{0}^{t}\|[x_{2}]^{\frac{5\gamma+2}{7}}e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{2}(u\cdot\nabla u)(\tau)\|_{L^{2}}d\tau.$
For the linear term involving initial data $\partial_{2}u_{0}$, we can obtain the expected decay through Lemma 2.3.
$\|[x_{2}]^{\frac{5\gamma+2}{7}}e^{\nu\Lambda_{1}^{2s}t}\partial_{2}u_{0,1}\|_{L^{2}}$ $\leq C(1+\nu t)^{-\frac{\sigma}{2s}}\left(\|[x_{2}]^{\frac{5\gamma+2}{7}}\Lambda_{1}^{-\sigma}\partial_{2}u_{0,1}\|_{L^{2}}+\|[x_{2}]^{\frac{5\gamma+2}{7}}\partial_{2}u_{0,1}\|_{L^{2}}\right)$
$\leq\frac{C_{0}}{4}\varepsilon(1+\nu t)^{-\frac{\sigma}{2s}}.$ (5.37)
Due to divergence-free condition, the nonlinear term can be expanded as
$\int_{0}^{t}\|[x_{2}]^{\frac{5\gamma+2}{7}}e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{2}(u\cdot\nabla u)(\tau)\|_{L^{2}}d\tau$
$\leq$ $\int_{0}^{t}\|[x_{2}]^{\frac{5\gamma+2}{7}}e^{\nu\Lambda_{1}^{2s}(t-\tau)}(u_{1}\partial_{1}\partial_{2}u)(\tau)\|_{L^{2}}d\tau+\int_{0}^{t}\|[x_{2}]^{\frac{5\gamma+2}{7}}e^{\nu\Lambda_{1}^{2s}(t-\tau)}(u_{2}\partial_{2}^{2}u)(\tau)\|_{L^{2}}d\tau$
$+\int_{0}^{t}\|[x_{2}]^{\frac{5\gamma+2}{7}}e^{\nu\Lambda_{1}^{2s}(t-\tau)}(\partial_{2}u_{1}\partial_{1}u_{2})(\tau)\|_{L^{2}}d\tau+\int_{0}^{t}\|[x_{2}]^{\frac{5\gamma+2}{7}}e^{\nu\Lambda_{1}^{2s}(t-\tau)}(\partial_{2}u_{2}\partial_{2}u_{2})(\tau)\|_{L^{2}}d\tau$
$:=$ $III_{1}^{w}+III_{2}^{w}+III_{3}^{w}+III_{4}^{w}.$
It follows from Lemma 2.1 and Lemma 2.7 that
$III_{1}^{w}\leq$ $C\int_{0}^{t}(t-\tau)^{-\frac{1}{4s}}\big\|[x_{2}]^{\frac{5\gamma+2}{7}}\|(u_{1}\partial_{1}\partial_{2}u)(\tau)\|_{L^{1}_{x_{1}}}\big\|_{L^{2}_{x_{2}}}d\tau$
$\leq$ $C\int_{0}^{t}(t-\tau)^{-\frac{1}{4s}}\|[x_{2}]^{\frac{5\gamma+2}{7}}u_{1}(\tau)\|_{L^{2}_{x_{1}}L^{\infty}_{x_{2}}}\|\partial_{1}\partial_{2}u(\tau)\|_{L^{2}}d\tau.$
It follows from Lemma 2.11 and $0<\gamma<\frac{3}{10}$ that
$\|[x_{2}]^{\frac{5\gamma+2}{7}}u_{1}(\tau)\|_{L^{2}_{x_{1}}L^{\infty}_{x_{2}}}\leq$ $C\|[x_{2}]^{\frac{5\gamma+2}{7}-\frac{1}{2}}u_{1}\|_{L^{2}}+C\|[x_{2}]^{\frac{5\gamma+2}{7}}u_{1}\|_{L^{2}}^{\frac{1}{2}}\|[x_{2}]^{\frac{5\gamma+2}{7}}\partial_{2}u_{1}\|_{L^{2}}^{\frac{1}{2}}$
$\leq$ $C\|[x_{2}]^{\frac{3\gamma+4}{7}}u_{1}\|_{L^{2}}+C\|[x_{2}]^{\frac{3\gamma+4}{7}}u_{1}\|_{L^{2}}^{\frac{1}{2}}\|[x_{2}]^{\frac{5\gamma+2}{7}}\partial_{2}u_{1}\|_{L^{2}}^{\frac{1}{2}}.$
By interpolation, for $0<\theta<\frac{1-2\sigma}{2\sigma+2}$, we have
$\|\partial_{1}\partial_{2}u(\tau)\|_{L^{2}}\leq\|\partial_{1}u\|_{L^{2}}^{1-\theta}\|\partial_{1}\partial_{2}^{\frac{1}{\theta}}u\|_{L^{2}}^{\theta}.$ (5.38)
Combining (5.38) and the decay rate (5.2)–(5.6), we obtain
$III^{w}_{1}\leq$ $C(C_{0}\varepsilon)^{2}\int_{0}^{t}(t-\tau)^{-\frac{1}{4s}}(1+\tau)^{-\frac{\sigma}{2s}}(1+\tau)^{-\frac{(\sigma+1)(1-\theta)}{2s}}d\tau.$ (5.39)
Similarly, we can deduce from Lemma 2.1 and Lemma 2.7 that
$III_{2}^{w}$ $\leq C\int_{0}^{t}(t-\tau)^{-\frac{1}{4s}}\|[x_{2}]^{\frac{5\gamma+2}{7}}u_{2}(\tau)\|_{L_{x_{1}}^{2}L_{x_{2}}^{\infty}}\|\partial_{2}^{2}u(\tau)\|_{L^{2}}d\tau.$
By applying Lemma 2.11 with $\zeta=\frac{5\gamma+2}{7}$, $\vartheta=\frac{2-2\gamma}{7}$, and combining with the Minkowski inequality and the divergence-free condition of $u$, we can deduce
$\big\|\|[x_{2}]^{\frac{5\gamma+2}{7}}u_{2}(\tau)\|_{L_{x_{1}}^{2}}\big\|_{L^{\infty}_{x_{2}}}\leq$ $C\|[x_{2}]^{\frac{5\gamma+2}{7}-\frac{1}{2}}u_{2}\|_{L^{2}}+C\|[x_{2}]^{\gamma}u_{2}\|_{L^{2}}^{\frac{1}{2}}\|[x_{2}]^{\frac{3\gamma+4}{7}}\partial_{2}u_{2}\|_{L^{2}}^{\frac{1}{2}}$
$\leq$ $C\|[x_{2}]^{\gamma}u_{2}\|_{L^{2}}+C\|[x_{2}]^{\gamma}u_{2}\|_{L^{2}}^{\frac{1}{2}}\|[x_{2}]^{\frac{3\gamma+4}{7}}\partial_{1}u_{1}\|_{L^{2}}^{\frac{1}{2}}.$
By employing the interpolation with $0<\theta<\frac{1-2\sigma}{2\sigma+2}$, we obtain ∥∂_2^2 u∥_L^2 ≤∥∂_2 u∥_L^2^1-θ ∥∂_2^1+1θ u∥_L^2^θ, and combining the decay rate (5.2)–(5.6), we can get
$III^{w}_{2}\leq$ $C(C_{0}\varepsilon)^{2}\int_{0}^{t}(t-\tau)^{-\frac{1}{4s}}(1+\tau)^{-\frac{\sigma+1}{2s}}(1+\tau)^{-\frac{\sigma(1-\theta)}{2s}}d\tau.$ (5.40)
Similarly, for $III_{3}^{w}$, by applying Lemma 2.7 and Hölder’s inequality, we have
$III_{3}^{w}$ $\leq C\int_{0}^{t}(t-\tau)^{-\frac{1}{4s}}\big\|[x_{2}]^{\frac{5\gamma+2}{7}}\|(\partial_{2}u_{1}\partial_{1}u_{2})(\tau)\|_{L^{1}_{x_{1}}}\big\|_{L^{2}_{x_{2}}}d\tau$
$\leq C\int_{0}^{t}(t-\tau)^{-\frac{1}{4s}}\|[x_{2}]^{\frac{5\gamma+2}{7}}\partial_{2}u_{1}(\tau)\|_{L^{2}}\|\partial_{1}u_{2}(\tau)\|_{L^{2}_{x_{1}}L^{\infty}_{x_{2}}}d\tau.$
By interpolation we have
$\|\partial_{1}u_{2}(\tau)\|_{L^{2}_{x_{1}}L^{\infty}_{x_{2}}}\leq$ $\|\partial_{1}u_{2}(\tau)\|_{L^{2}}^{\frac{1}{2}}\|\partial_{1}\partial_{2}u_{2}(\tau)\|_{L^{2}}^{\frac{1}{2}}$
$\leq$ $\|\partial_{1}u_{2}(\tau)\|_{L^{2}}^{\frac{1}{2}}\|\partial_{1}u_{2}(\tau)\|_{L^{2}}^{\frac{1-\theta}{2}}\|\partial_{1}\partial_{2}^{\frac{1}{\theta}}u_{2}(\tau)\|_{L^{2}}^{\frac{\theta}{2}}.$
Then combining the decay rates (5.2)–(5.6), we arrive at
$III^{w}_{3}\leq$ $C(C_{0}\varepsilon)^{2}\int_{0}^{t}(t-\tau)^{-\frac{1}{4s}}(1+\tau)^{-\frac{4\sigma+1}{4s}}(1+\tau)^{-\frac{(2\sigma+1)(1-\theta)}{4s}}d\tau.$ (5.41)
For the term $III_{4}^{w}$, we again employ Lemma 2.1 and Lemma 2.7 to get
$III_{4}^{w}$ $\leq C\int_{0}^{t}(t-\tau)^{-\frac{1}{4s}}\|[x_{2}]^{\frac{5\gamma+2}{7}}\partial_{2}u_{2}(\tau)\|_{L^{2}}\|\partial_{2}u_{2}(\tau)\|_{L^{2}}^{\frac{1}{2}}\|\partial_{2}^{2}u_{2}(\tau)\|_{L^{2}}^{\frac{1}{2}}d\tau.$
In view of the divergence-free condition and Lemma 2.9, we can deduce that
$III_{4}^{w}$ $\leq C\int_{0}^{t}(t-\tau)^{-\frac{1}{4s}}\|[x_{2}]^{\frac{3\gamma+4}{7}}\partial_{1}u_{1}(\tau)\|_{L^{2}}^{\frac{3}{2}}\|\partial_{1}\partial_{2}u_{1}(\tau)\|_{L^{2}}^{\frac{1}{2}}d\tau.$
Then, using the decay estimates (5.2)–(5.6), we obtain
$III^{w}_{4}\leq$ $C(C_{0}\varepsilon)^{2}\int_{0}^{t}(t-\tau)^{-\frac{1}{4s}}(1+\tau)^{-\frac{3\sigma+3}{4s}}d\tau.$ (5.42)
Comparing the decay rate of $II^{w}_{1}II$$II^{w}_{4}$, we can deduce
$III_{1}^{w}+III_{2}^{w}+III_{3}^{w}+III_{4}^{w}\leq$ $C(C_{0}\varepsilon)^{2}\int_{0}^{t}(t-\tau)^{-\frac{1}{4s}}(1+\tau)^{-\frac{\sigma}{2s}}(1+\tau)^{-\frac{(\sigma+1)(1-\theta)}{2s}}d\tau$
$\leq$ $C(C_{0}\varepsilon)^{2}\int_{0}^{t}(t-\tau)^{-\frac{1}{4s}}(1+\tau)^{-\frac{2\sigma+1}{2s}}(1+\tau)^{\frac{\theta(\sigma+1)}{2s}}d\tau$
$\leq$ $C(C_{0}\varepsilon)^{2}\int_{0}^{t}(t-\tau)^{-\frac{1}{4s}}(1+\tau)^{-1}(1+\tau)^{\frac{\theta(\sigma+1)}{2s}}d\tau.$
By applying Lemma 2.8 and leveraging the condition $\theta<\frac{1-2\sigma}{2\sigma+2}$, we obtain
$\int_{0}^{t}(t-\tau)^{-\frac{1}{4s}}(1+\tau)^{-1}(1+\tau)^{\frac{\theta(\sigma+1)}{2s}}d\tau\leq(1+t)^{\frac{2\theta(\sigma+1)-1}{4s}}\leq(1+t)^{-\frac{\sigma}{2s}}.$
Then combining with (5.37), we can deduce the bootstrap conclusion (5.10).
5.6 The decay of $\partial_{1}u$ with weight
By Duhamel’s principle and Lemma 5.1, we have
$\|[x_{2}]^{\frac{3\gamma+4}{7}}\partial_{1}u(t)\|_{L^{2}}\leq\|[x_{2}]^{\frac{3\gamma+4}{7}}e^{\nu\Lambda_{1}^{2s}t}\partial_{1}u_{0}\|_{L^{2}}+C\int_{0}^{t}\|[x_{2}]^{\frac{3\gamma+4}{7}}e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}(u\cdot\nabla u)(\tau)\|_{L^{2}}d\tau.$
For the linear term involving initial data $\partial_{1}u_{0}$, we can obtain the expected decay through Lemma 2.7:
$\|[x_{2}]^{\frac{3\gamma+4}{7}}e^{\nu\Lambda_{1}^{2s}t}\partial_{1}u_{0}\|_{L^{2}}$ $\leq C(1+\nu t)^{-\frac{\sigma+1}{2s}}\left(\|[x_{2}]^{\frac{3\gamma+4}{7}}\Lambda_{1}^{-\sigma}u_{0}\|_{L^{2}}+\|[x_{2}]^{\frac{3\gamma+4}{7}}\partial_{1}u_{0}\|_{L^{2}}\right)$
$\leq\frac{C_{0}}{4}\varepsilon(1+\nu t)^{-\frac{\sigma+1}{2s}}.$ (5.43)
Applying Lemma 2.1, Lemma 2.2 and Lemma 2.7,
$\int_{0}^{t}\|[x_{2}]^{\frac{3\gamma+4}{7}}e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}(u\cdot\nabla u)(\tau)\|_{L^{2}}d\tau$
$\leq$ $C\int_{0}^{t}(t-\tau)^{-\frac{3}{4s}}\big\|[x_{2}]^{\frac{3\gamma+4}{7}}\|(u\cdot\nabla u)(\tau)\|_{L_{x_{1}}^{1}}\big\|_{L_{x_{2}}^{2}}d\tau$
$\leq$ $C\int_{0}^{t}(t-\tau)^{-\frac{3}{4s}}\|u_{1}(\tau)\|_{L^{2}}^{\frac{1}{2}}\|\partial_{2}u_{1}(\tau)\|_{L^{2}}^{\frac{1}{2}}\|[x_{2}]^{\frac{3\gamma+4}{7}}\partial_{1}u_{1}(\tau)\|_{L^{2}}d\tau$
$+C\int_{0}^{t}(t-\tau)^{-\frac{3}{4s}}\|[x_{2}]^{\frac{-2\gamma+2}{7}}u_{2}\|_{L^{2}_{x_{1}}L^{\infty}_{x_{2}}}\|[x_{2}]^{\frac{5\gamma+2}{7}}\partial_{2}u_{1}(\tau)\|_{L^{2}}d\tau$
$+C\int_{0}^{t}(t-\tau)^{-\frac{3}{4s}}\|u_{1}(\tau)\|_{L^{2}}^{\frac{1}{2}}\|\partial_{2}u_{1}(\tau)\|_{L^{2}}^{\frac{1}{2}}\|[x_{2}]^{\frac{3\gamma+4}{7}}\partial_{1}u_{2}(\tau)\|_{L^{2}}d\tau$
$+C\int_{0}^{t}(t-\tau)^{-\frac{3}{4s}}\|u_{2}(\tau)\|_{L^{2}}^{\frac{1}{2}}\|\partial_{2}u_{2}(\tau)\|_{L^{2}}^{\frac{1}{2}}\|[x_{2}]^{\frac{3\gamma+4}{7}}\partial_{2}u_{2}(\tau)\|_{L^{2}}d\tau$
$:=$ $IV_{1}^{w}+IV_{2}^{w}+IV_{3}^{w}+IV_{4}^{w}.$
Applying Lemma 2.11 and the range of $\gamma$ ($0<\gamma<\frac{3}{10}$), we can obtain
$\|[x_{2}]^{\frac{-2\gamma+2}{7}}u_{2}\|_{L^{2}_{x_{1}}L^{\infty}_{x_{2}}}\leq$ $C\|[x_{2}]^{\frac{-2\gamma+2}{7}-\frac{1}{2}}u_{2}\|_{L^{2}}+C\|[x_{2}]^{\gamma}u_{2}\|_{L^{2}}^{\frac{1}{2}}\|[x_{2}]^{\frac{-11\gamma+4}{7}}\partial_{2}u_{2}\|_{L^{2}}^{\frac{1}{2}}$
$\leq$ $C\|[x_{2}]^{\gamma}u_{2}\|_{L^{2}}+C\|[x_{2}]^{\gamma}u_{2}\|_{L^{2}}^{\frac{1}{2}}\|[x_{2}]^{\frac{3\gamma+4}{7}}\partial_{1}u_{1}\|_{L^{2}}^{\frac{1}{2}}.$ (5.44)
Combining (5.44) and the decay rate (5.2)–(5.6), we can conclude
$IV_{1}^{w}+IV_{2}^{w}+IV_{3}^{w}+IV_{4}^{w}$
$\leq$ $CC_{0}^{2}\varepsilon^{2}\int_{0}^{t-1}(t-\tau)^{-\frac{3}{4s}}(\tau+1)^{-\frac{2\sigma+1}{2s}}d\tau\leq CC_{0}^{2}\varepsilon^{2}(t+1)^{-\frac{\sigma+1}{2s}}.$ (5.45)
Combining (5.43) and (5.45), we can obtain bootstrap conclusion (5.11).
5.7 The decay of $\partial_{1}u_{2}$ with weight
By Duhamel’s principle and taking the $L^{2}$-norm, we have
$\|[x_{2}]^{\gamma}\partial_{1}u_{2}(t)\|_{L^{2}}\leq\|[x_{2}]^{\gamma}e^{\nu\Lambda_{1}^{2s}t}\partial_{1}u_{0,2}\|_{L^{2}}+\int_{0}^{t}\|[x_{2}]^{\gamma}e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{2}\Delta^{-1}\nabla\times(u\cdot\nabla u)(\tau)\|_{L^{2}}d\tau.$
The linear term decays as
$\|[x_{2}]^{\gamma}e^{\nu\Lambda_{1}^{2s}t}\partial_{1}u_{0,2}\|_{L^{2}}$ $\leq C(1+\nu t)^{-\frac{\sigma+2}{2s}}\big(\|[x_{2}]^{\gamma}\Lambda_{1}^{-\sigma}\psi_{0}\|_{L^{2}}+\|[x_{2}]^{\gamma}\partial_{1}u_{0,2}\|_{L^{2}}\big)$
$\leq C\varepsilon(1+\nu t)^{-\frac{\sigma+2}{2s}}\leq\frac{C_{0}}{4}\varepsilon(1+\nu t)^{-\frac{2\sigma+1}{2s}}.$
Similar to (5.31), we have
$\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{2}[x_{2}]^{\gamma}\Delta^{-1}\nabla\times(u\cdot\nabla u)(\tau)\|_{L^{2}}d\tau$
$=$ $\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{2}[x_{2}]^{\gamma}\Delta^{-1}\nabla\times\mbox{div}\ (u\otimes u)(\tau)\|_{L^{2}}d\tau$
$\leq$ $\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{2}[x_{2}]^{\gamma}\partial_{1}\partial_{2}\Delta^{-1}(u_{1}u_{1})(\tau)\|_{L^{2}}d\tau$
$+\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{2}[x_{2}]^{\gamma}\partial_{1}^{2}\Delta^{-1}(u_{1}u_{2})(\tau)\|_{L^{2}}d\tau$
$+\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{2}[x_{2}]^{\gamma}\Delta^{-1}\partial_{2}^{2}(u_{2}u_{1})(\tau)\|_{L^{2}}d\tau$
$+\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{2}[x_{2}]^{\gamma}\Delta^{-1}\partial_{1}\partial_{2}(u_{2}u_{2})(\tau)\|_{L^{2}}d\tau$
$:=$ $V_{1}^{w}+V_{2}^{w}+V_{3}^{w}+V_{4}^{w}.$
The term $V_{1}^{w}$ can be split into:
$V_{1}^{w}=$ $\int_{0}^{t-1}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{2}[x_{2}]^{\gamma}\partial_{1}\partial_{2}\Delta^{-1}(u_{1}u_{1})(\tau)\|_{L^{2}}d\tau$
$+\int_{t-1}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{2}[x_{2}]^{\gamma}\partial_{1}\partial_{2}\Delta^{-1}(u_{1}u_{1})(\tau)\|_{L^{2}}d\tau.$ (5.46)
And the first part in right-hand side of (5.46) can be written as
$\int_{0}^{t-1}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{2}[x_{2}]^{\gamma}\partial_{1}\partial_{2}\Delta^{-1}(u_{1}u_{1})(\tau)\|_{L^{2}}d\tau$
$=$ $\int_{0}^{t-1}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{2}[x_{2}]^{\gamma}\partial_{1}\partial_{2}^{2}\Delta^{-1}\int_{-\infty}^{x_{2}}(u_{1}u_{1})(y_{2})dy_{2}(\tau)\|_{L^{2}}d\tau.$
Similar to (5.33) and (5.34), we then use Lemma 2.13 and subsequently Lemma 2.10 to get
$\int_{0}^{t-1}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{2}[x_{2}]^{\gamma}\partial_{1}\partial_{2}^{2}\Delta^{-1}\int_{-\infty}^{x_{2}}(u_{1}u_{1})(y_{2})dy_{2}(\tau)\|_{L^{2}}d\tau$
$\leq$ $C\int_{0}^{t-1}\|[x_{2}]^{\gamma}e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{3}\int_{-\infty}^{x_{2}}(u_{1}u_{1})(y_{2})dy_{2}\|_{L^{2}}d\tau$
$\leq$ $C\int_{0}^{t-1}\|[x_{2}]^{\gamma+1}e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{3}(u_{1}u_{1})\|_{L^{2}}d\tau.$ (5.47)
For the second part of (5.46), by using Lemma 2.13, we can get
$\int_{t-1}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{2}[x_{2}]^{\gamma}\partial_{1}\partial_{2}\Delta^{-1}(u_{1}u_{1})(\tau)\|_{L^{2}}d\tau\leq C\int_{t-1}^{t}\|\partial_{1}^{2}[x_{2}]^{\gamma}(u_{1}u_{1})(\tau)\|_{L^{2}}d\tau.$ (5.48)
Applying Lemma 2.7 to (5.47) and substituting the resulting estimate together with (5.48) into (5.48), we obtain
$V_{1}^{w}\leq$ $C\int_{0}^{t-1}(t-\tau)^{-\frac{5}{4s}}\big\|[x_{2}]^{\gamma+1}\|\partial_{1}(u_{1}u_{1})\|_{L^{1}_{x_{1}}}\big\|_{L_{x_{2}}^{2}}d\tau+C\int_{t-1}^{t}\|[x_{2}]^{\gamma}\partial_{1}^{2}(u_{1}u_{1})\|_{L^{2}}d\tau$
$:=$ $V_{11}^{w}+V_{12}^{w}.$
Using Hölder inequality in $x_{1}$ and $x_{2}$ direction, we can obtain
$\big\|[x_{2}]^{\gamma+1}\|\partial_{1}(u_{1}u_{1})\|_{L^{1}_{x_{1}}}\big\|_{L_{x_{2}}^{2}}\leq$ $C\|[x_{2}]^{\frac{4\gamma+3}{7}}u_{1}\|_{L^{2}_{x_{1}}L^{\infty}_{x_{2}}}\|[x_{2}]^{\frac{3\gamma+4}{7}}\partial_{1}u_{1}\|_{L^{2}}.$
It follows from Lemma 2.11 and the range of $\gamma$ ($0<\gamma<\frac{3}{10}$) that
$\|[x_{2}]^{\frac{4\gamma+3}{7}}u_{1}\|_{L^{2}_{x_{1}}L^{\infty}_{x_{2}}}\leq$ $C\|[x_{2}]^{\frac{4\gamma+3}{7}-\frac{1}{2}}u_{1}\|_{L^{2}}+C\|[x_{2}]^{\frac{3\gamma+4}{7}}u_{1}\|_{L^{2}}^{\frac{1}{2}}\|[x_{2}]^{\frac{5\gamma+2}{7}}\partial_{2}u_{1}\|_{L^{2}}^{\frac{1}{2}}$
$\leq$ $C\|[x_{2}]^{\frac{3\gamma+4}{7}}u_{1}\|_{L^{2}}+C\|[x_{2}]^{\frac{3\gamma+4}{7}}u_{1}\|_{L^{2}}^{\frac{1}{2}}\|[x_{2}]^{\frac{5\gamma+2}{7}}\partial_{2}u_{1}\|_{L^{2}}^{\frac{1}{2}}.$ (5.49)
Substituting (5.2), (5.6), we can handle $V_{11}^{w}$ as:
$V_{11}^{w}\leq CC_{0}^{2}\varepsilon^{2}\int_{0}^{t-1}(t-\tau)^{-\frac{5}{4s}}(1+\tau)^{-\frac{2\sigma+1}{2s}}d\tau.$
By Lemma 2.8, we have
$V_{11}^{w}\leq CC_{0}^{2}\varepsilon^{2}(1+t)^{-\frac{2\sigma+1}{2s}}.$
For $V_{12}^{w}$, by Hölder’s inequality and the range of $\gamma$, we can deduce
$\|[x_{2}]^{\gamma}\partial_{1}^{2}(u_{1}u_{1})\|_{L^{2}}\leq$ $C\|[x_{2}]^{\gamma}(\partial_{1}^{2}u_{1})u_{1}\|_{L^{2}}+C\|[x_{2}]^{\gamma}(\partial_{1}u_{1})^{2}\|_{L^{2}}$
$\leq$ $C\|[x_{2}]^{\gamma}u_{1}\|_{L_{x_{1}}^{\infty}L_{x_{2}}^{2}}\|\partial_{1}^{2}u_{1}\|_{L_{x_{1}}^{2}L_{x_{2}}^{\infty}}+C\|[x_{2}]^{\gamma}\partial_{1}u_{1}\|_{L^{2}}\|\partial_{1}u_{1}\|_{L^{\infty}}$
$\leq$ $C\|[x_{2}]^{\gamma}u_{1}\|_{L^{2}}^{\frac{1}{2}}\|[x_{2}]^{\gamma}\partial_{1}u_{1}\|_{L^{2}}^{\frac{1}{2}}\|\partial_{1}^{2}u_{1}\|_{L^{2}}^{\frac{1}{2}}\|\partial_{1}^{2}\partial_{2}u_{1}\|_{L^{2}}^{\frac{1}{2}}$
$+C\|[x_{2}]^{\gamma}\partial_{1}u_{1}\|_{L^{2}}\|\partial_{1}u_{1}\|_{L^{2}}^{\frac{1}{2}}\|\partial_{1}\Delta u_{1}\|_{L^{2}}^{\frac{1}{2}}$
$\leq$ $C\|[x_{2}]^{\frac{3\gamma+4}{7}}u_{1}\|_{L^{2}}^{\frac{1}{2}}\|[x_{2}]^{\frac{3\gamma+4}{7}}\partial_{1}u_{1}\|_{L^{2}}^{\frac{1}{2}}\|\partial_{1}^{2}u_{1}\|_{L^{2}}^{\frac{1}{2}}\|\partial_{1}^{2}\partial_{2}u_{1}\|_{L^{2}}^{\frac{1}{2}}$
$+C\|[x_{2}]^{\frac{3\gamma+4}{7}}\partial_{1}u_{1}\|_{L^{2}}\|\partial_{1}u_{1}\|_{L^{2}}^{\frac{1}{2}}\|\partial_{1}\Delta u_{1}\|_{L^{2}}^{\frac{1}{2}}.$ (5.50)
By interpolation, we have
$\|\partial_{1}^{2}u_{1}\|_{L^{2}}\|\partial_{1}^{2}\partial_{2}u_{1}\|_{L^{2}}\leq\|\partial_{1}u_{1}\|_{L^{2}}(\|\partial_{1}^{3}u_{1}\|_{L^{2}}+\|\partial_{1}^{3}\partial_{2}^{2}u_{1}\|_{L^{2}})\leq C_{0}^{2}\varepsilon^{2}(1+t)^{-\frac{\sigma+1}{2s}}.$ (5.51)
Substituting (5.51) into (5.50), and combining (5.1), (5.2) and (5.5), we have
$V_{12}^{w}\leq CC_{0}^{2}\varepsilon^{2}(1+t)^{-\frac{2\sigma+1}{2s}}.$
Since the Riesz operators are bounded in weighted $L^{2}$ space (Lemma 2.13), we can derive the estimates as follows:
$V^{w}_{2}+V^{w}_{3}+V^{w}_{4}\leq$ $C\int_{0}^{t}\|[x_{2}]^{\gamma}e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{2}(u_{1}u_{2})(\tau)\|_{L^{2}}d\tau$
$+C\int_{0}^{t}\|[x_{2}]^{\gamma}e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{2}(u_{2}u_{2})(\tau)\|_{L^{2}}d\tau.$
By using Lemma 2.7, we have
$V^{w}_{2}+V^{w}_{3}+V^{w}_{4}\leq$ $C\int_{0}^{t-1}(t-\tau)^{-\frac{5}{4s}}\big\|[x_{2}]^{\gamma}\|(u_{1}u_{2})\|_{L_{x_{1}}^{1}}\big\|_{L_{x_{2}}^{2}}d\tau+C\int_{t-1}^{t}\|[x_{2}]^{\gamma}\partial_{1}^{2}(u_{1}u_{2})\|_{L^{2}}d\tau$
$+C\int_{0}^{t-1}(t-\tau)^{-\frac{5}{4s}}\big\|[x_{2}]^{\gamma}\|(u_{2}u_{2})\|_{L_{x_{1}}^{1}}\big\|_{L_{x_{2}}^{2}}d\tau+C\int_{t-1}^{t}\|[x_{2}]^{\gamma}\partial_{1}^{2}(u_{2}u_{2})\|_{L^{2}}d\tau$
$:=$ $V^{w}_{a1}+V^{w}_{a2}+V^{w}_{b1}+V^{w}_{b2}.$
The first two terms can be handled by Lemma 2.2 and Lemma 2.8,
$V^{w}_{a1}\leq$ $C\int_{0}^{t-1}(t-\tau)^{-\frac{5}{4s}}\|[x_{2}]^{\gamma}u_{2}\|_{L^{2}}\|u_{1}\|_{L^{2}}^{\frac{1}{2}}\|\partial_{2}u_{1}\|_{L^{2}}^{\frac{1}{2}}d\tau$
$\leq$ $CC_{0}^{2}\varepsilon^{2}\int_{0}^{t-1}(t-\tau)^{-\frac{5}{4s}}(\tau+1)^{-\frac{2\sigma+1}{2s}}d\tau$
$\leq$ $CC_{0}^{2}\varepsilon^{2}(t+1)^{-\frac{2\sigma+1}{2s}},$
and
$V^{w}_{b1}\leq$ $C\int_{0}^{t-1}(t-\tau)^{-\frac{5}{4s}}\|[x_{2}]^{\gamma}u_{2}\|_{L^{2}}\|u_{2}\|_{L^{2}}^{\frac{1}{2}}\|\partial_{2}u_{2}\|_{L^{2}}^{\frac{1}{2}}d\tau$
$\leq$ $CC_{0}^{2}\varepsilon^{2}\int_{0}^{t-1}(t-\tau)^{-\frac{5}{4s}}(\tau+1)^{-\frac{2\sigma+2}{2s}}d\tau$
$\leq$ $CC_{0}^{2}\varepsilon^{2}(t+1)^{-\frac{2\sigma+1}{2s}}.$
The estimate of $V^{w}_{a2}$ and $V^{w}_{b2}$ are similar to the estimate of $V_{12}$ in (4.36), (4.37) and (4.38), which can be bounded by $C(C_{0}\varepsilon)^{2}(1+t)^{-\frac{2\sigma+1}{2s}}$.
Combining the estimates above, we can deduce the bootstrap conclusion (5.12) by taking $\varepsilon$ sufficiently small.
Appendix A Proof of Lemmas in Section 2
In this section, we will provide detailed proofs for all the lemmas introduced in Section 2.
Proof of Lemma 2.4:
By Plancherel’s theorem, we have
$\|\Lambda^{m}(fg)\|_{L^{2}}^{2}=\int_{\mathbb{R}^{d}}|\xi|^{2m}|\widehat{fg}(\xi)|^{2}d\xi=\int_{\mathbb{R}^{d}}|\xi|^{2m}\left|\int\hat{f}(\eta)\hat{g}(\xi-\eta)d\eta\right|^{2}d\xi$
By Peetre’s inequality, there exists a constant $C$ depending on $2m$ such that
$|\xi|^{2m}\leq C(|\eta|^{2m}+|\xi-\eta|^{2m}).$
Then, applying this inequality in the frequency domain and splitting the integral accordingly, we obtain the following estimate
$\|\Lambda^{m}(fg)\|_{L^{2}}^{2}\leq$ $C\int_{\mathbb{R}^{d}}\left(\int_{\mathbb{R}^{d}}|\eta|^{m}|\hat{f}(\eta)||\hat{g}(\xi-\eta)|d\eta\right)^{2}d\xi$
$+C\int_{\mathbb{R}^{d}}\left(\int_{\mathbb{R}^{d}}|\hat{f}(\eta)||\xi-\eta|^{m}|\hat{g}(\xi-\eta)|d\eta\right)^{2}d\xi$
$\leq$ $C(\left\|(\Lambda^{m}f)g\right\|_{L^{2}}^{2}+\left\|f(\Lambda^{m}g)\right\|_{L^{2}}^{2}).$
Thus, the desired inequality (2.1) follows. The derivation of (2.2) follows a similar rationale and is thus omitted here. This completes the proof of Lemma 2.4.
Proof of Lemma 2.8:
To prove the inequality (2.6), we cut the time interval of integration at the midpoint (t1)/2(t-1)/2. On the first subinterval, we bound the factor $(t-\tau)^{-\alpha}$, on the second subinterval, we instead bound the factor $(1+\tau)^{-\beta}$, this leads to
$\int_{0}^{t-1}(t-\tau)^{-\alpha}(1+\tau)^{-\beta}\>d\tau$
$\leq$ $\int_{0}^{(t-1)/2}(t-\tau)^{-\alpha}(1+\tau)^{-\beta}\>d\tau+\int_{(t-1)/2}^{t-1}(t-\tau)^{-\alpha}(1+\tau)^{-\beta}\>d\tau$
$\leq$ $2^{\alpha}(t+1)^{-\alpha}\int_{0}^{(t-1)/2}(1+\tau)^{-\beta}d\tau+2^{\beta}(1+t)^{-\beta}\int_{1}^{(t+1)/2}\tau^{-\alpha}d\tau.$
By direct calculation, we have
$\int_{1}^{(t+1)/2}\tau^{-\alpha}d\tau\leq\begin{cases}\ln((t+1)/2),\ &\alpha=1\\ -\frac{1}{1-\alpha},\ &\alpha>1.\end{cases}$
When $\beta>\alpha\geq 1$, we have
$\int_{0}^{t-1}(t-\tau)^{-\alpha}(1+\tau)^{-\beta}\>d\tau\leq$ $C(1+t)^{-\alpha}.$
If $\beta>1$ and $\alpha>\beta$, we have
$\int_{0}^{t-1}(t-\tau)^{-\alpha}(1+\tau)^{-\beta}\>d\tau\leq$ $C(1+t)^{-\beta}.$
To prove (2.7), similar argument shows that
$\int_{0}^{t}(t-\tau)^{-\alpha}(1+\tau)^{-\beta}\>d\tau\leq$ $\int_{0}^{t/2}(t-\tau)^{-\alpha}(1+\tau)^{-\beta}\>d\tau+\int_{t/2}^{t}(t-\tau)^{-\alpha}(1+\tau)^{-\beta}\>d\tau$
$\leq$ $2^{\alpha}t^{-\alpha}\int_{0}^{t/2}(1+\tau)^{-\beta}d\tau+(1+t/2)^{-\beta}\int_{0}^{t/2}\tau^{-\alpha}d\tau.$
By direct calculation, because of $\alpha<1$, we have
$\int_{0}^{t/2}\tau^{-\alpha}d\tau\leq C\frac{(t/2)^{1-\alpha}}{1-\alpha},\quad\int_{0}^{t/2}(1+\tau)^{-\beta}d\tau\leq\begin{cases}\frac{C}{1-\beta}(1+t)^{1-\beta},&\beta<1,\\ C\ln(1+t),&\beta=1,\\ -\frac{C}{1-\beta},&\beta>1.\end{cases}$
Then we obtain the inequality (2.7).
Proof of Lemma 2.10:
We decompose [x]^γ = —x—^γ+ s(x),where
$s(x):=\begin{cases}1-|x|^{\gamma},&\ |x|\leq 1,\\ 0,&\ |x|>1.\end{cases}$
Then
$\int_{\mathbb{R}}f^{2}[x]^{2\gamma}dx\leq$ $\int_{\mathbb{R}}f^{2}x^{2\gamma}dx+2\int_{-1}^{1}f^{2}|x|^{\gamma}(1-|x|^{\gamma})dx+\int_{-1}^{1}f^{2}(1-|x|^{\gamma})^{2}dx$
$:=$ $T_{1}+T_{2}+T_{3}.$ (A.1)
For the $T_{1}$, using integration by parts, we have
$\int_{\mathbb{R}}f^{2}x^{2\gamma}dx=$ $\int_{\mathbb{R}}\partial_{x}x(f^{2}x^{2\gamma})dx$
$=$ $-{2\gamma}\int_{\mathbb{R}}f^{2}x^{2\gamma}dx-2\int_{\mathbb{R}}x^{2\gamma+1}ff^{\prime}dx.$
Applying Cauchy-Schwarz inequality and Young’s inequality, we obtain
$(1+2\gamma)\int_{\mathbb{R}}x^{2\gamma}f^{2}dx\leq$ $2\left|\int_{\mathbb{R}}x^{2\gamma+1}ff^{\prime}dx\right|$
$\leq$ $2\left(\int_{\mathbb{R}}x^{2\gamma}f^{2}dx\right)^{\frac{1}{2}}\left(\int_{\mathbb{R}}x^{2\gamma+2}(f^{\prime})^{2}dx\right)^{\frac{1}{2}}$
$\leq$ $\int_{\mathbb{R}}x^{2\gamma}f^{2}dx+\int_{\mathbb{R}}x^{2\gamma+2}(f^{\prime})^{2}dx.$
Cancelling the common term $\int_{\mathbb{R}^{2}}x^{2}f^{2}dx$ from both sides yields
$T_{1}\leq\frac{1}{2\gamma}\|x^{\gamma+1}f^{\prime}\|_{L^{2}}^{2}\leq\frac{1}{2\gamma}\|[x]^{\gamma+1}f^{\prime}\|_{L^{2}}^{2}.$ (A.2)
By a similar argument with $T_{1}$, we can obtain
$T_{2}\leq$ $2\int_{-1}^{1}f^{2}|x|^{\gamma}dx\leq 2\int_{\mathbb{R}}f^{2}|x|^{\gamma}dx$
$\leq$ $\frac{2}{\gamma}\int_{\mathbb{R}}{f^{\prime}}^{2}|x|^{\gamma+2}dx\leq\frac{2}{\gamma}\int_{\mathbb{R}}{f^{\prime}}^{2}[x]^{2\gamma+2}dx.$ (A.3)
By direct calculation,
$T_{3}\leq\int_{-1}^{1}f^{2}(1-|x|)^{2}dx.$ (A.4)
By Poincaré inequality for functions with compact support, we have
$\int_{-1}^{1}f^{2}(1-|x|)^{2}dx\leq$ $\left(\frac{2}{\pi}\right)^{2}\int_{-1}^{1}(f^{\prime}(1-|x|)+f(-sgn(x)))^{2}dx$
$\leq$ $\frac{8}{\pi^{2}}\int_{-1}^{1}{f^{\prime}}^{2}(1-|x|)^{2}dx+\frac{8}{\pi^{2}}\int_{-1}^{1}f^{2}dx.$ (A.5)
Combining (A), (A.2), (A), (A.4) and (A.5), we have
$\int_{\mathbb{R}}f^{2}[x]^{2\gamma}dx\leq$ $\left(\frac{1}{2\gamma}+\frac{2}{\gamma}\right)\|[x]^{\gamma+1}f^{\prime}\|_{L^{2}}^{2}+\frac{8}{\pi^{2}}\int_{-1}^{1}{f^{\prime}}^{2}(1-|x|)^{2}dx$
$+\frac{8}{\pi^{2}}\int_{-1}^{1}f^{2}dx.$ (A.6)
Note that the third term in the right-hand side of (A) can be absorbed by $\int_{\mathbb{R}}f^{2}[x]^{2\gamma}dx$. Hence we have
$\|[x]^{\gamma}f\|_{L^{2}{(\mathbb{R})}}^{2}\leq C\|[x]^{\gamma+1}f^{\prime}\|_{L^{2}{(\mathbb{R})}}^{2},$
which complete the proof of Lemma 2.10.
Proof of Lemma 2.11:
Noticing that
$[x]^{2\zeta}f^{2}=$ $2\zeta\int_{-\infty}^{x}[y]^{2\zeta-1}f^{2}dy+2\int_{-\infty}^{x}[y]^{2\zeta}ff^{\prime}dy.$
By taking $L^{\infty}$-norm and square root on both side of the resulting inequality, we can deduce from Hölder inequality that
$\|[x]^{\zeta}f\|_{L^{\infty}(\mathbb{R})}\leq 2\zeta\|[x]^{\zeta-\frac{1}{2}}f\|_{L^{2}(\mathbb{R})}+2\|[x]^{\zeta-\vartheta}f\|_{L^{2}(\mathbb{R})}^{\frac{1}{2}}\|[x]^{\zeta+\vartheta}f^{\prime}\|_{L^{2}(\mathbb{R})}^{\frac{1}{2}}.$
This completes the proof of Lemma 2.11.
Proof of Lemma 2.13:
Firstly, We claim that $x_{2}]^{\kappa}$ is an $A_{2}$ weight in $\mathbb{R}^{2}$ for $1<\kappa<1$. To establish this, noticing that $|x_{2}|^{\kappa}$ is an $A_{2}$ weight on $\mathbb{R}$ (see Chapter V, Section 6.4 of [21] for more details).
For any rectangle $Q=I_{1}\times I_{2}$ on $\mathbb{R}^{2}$, the $A_{2}$ condition requires:
$\left(\frac{1}{|Q|}\int_{Q}[x_{2}]^{\kappa}\,dx\right)\left(\frac{1}{|Q|}\int_{Q}[x_{2}]^{-\kappa}\,dx\right)$
$=$ $\left(\frac{1}{|I_{1}||I_{2}|}\int_{I_{1}}\int_{I_{2}}[x_{2}]^{\kappa}\,dx_{2}dx_{1}\right)\left(\frac{1}{|I_{1}||I_{2}|}\int_{I_{1}}\int_{I_{2}}[x_{2}]^{-\kappa}\,dx_{2}dx_{1}\right)$
$\leq$ $\left(\frac{1}{|I_{2}|}\int_{I_{2}}[x_{2}]^{\kappa}\,dx_{2}\right)\left(\frac{1}{|I_{2}|}\int_{I_{2}}[x_{2}]^{-\kappa}\,dx_{2}\right).$
This reduces the problem to the one-dimensional case.
Without loss of generality, we only provide the proof for the case $0<\kappa<1$. Noticing that $x_{2}]^{\kappa}$ and $|x_{2}|^{\kappa}$ differ by less than 1 only on regions with measure less than 1. Therefore, for any interval $I_{2}$, there exists a constant $C$ such that
$\left(\frac{1}{|I_{2}|}\int_{I_{2}}[x_{2}]^{\kappa}\,dx_{2}\right)\leq\left(\frac{1}{|I_{2}|}\int_{I_{2}}|x_{2}|^{\kappa}\,dx_{2}\right)+C.$
On the other hand, for any interval $I_{2}$,
$\left(\frac{1}{|I_{2}|}\int_{I_{2}}[x_{2}]^{-\kappa}\,dx_{2}\right)\leq\min\left\{1,\left(\frac{1}{|I_{2}|}\int_{I_{2}}|x_{2}|^{-\kappa}\,dx_{2}\right)\right\}.$
Combining the fact that $|x_{2}|^{\kappa}$ is an $A_{2}$ weight on $\mathbb{R}$, we have
$\left(\frac{1}{|Q|}\int_{Q}[x_{2}]^{\kappa}\,dx\right)\left(\frac{1}{|Q|}\int_{Q}[x_{2}]^{-\kappa}\,dx\right)$
$\leq$ $\left(\frac{1}{|I_{2}|}\int_{I_{2}}|x_{2}|^{\kappa}\,dx_{2}\right)\left(\frac{1}{|I_{2}|}\int_{I_{2}}|x_{2}|^{-\kappa}\,dx_{2}\right)+C\leq 2C.$
Hence $x_{2}]^{\kappa}$ is an $A_{2}$ weight on $\mathbb{R}^{2}$. Then it follows from the boundness of Calderón-Zygmund operator in $L^{p}(\omega_{p}dx)$ (see [21] Theorem 2 in V.4.2 for example), one can deduce the desired inequality and complete the proof of Lemma 2.13.
Appendix B The decay of $u_{2}$ for $0<s<\frac{3}{4}$
Applying Duhamel’s principle to (4.21) and taking the $L^{2}$-norm, we can obtain
$\|u_{2}(t)\|_{L^{2}}\leq\|e^{\nu\Lambda_{1}^{2s}t}u_{0,2}\|_{L^{2}}+\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}\Delta^{-1}\nabla\times(u\cdot\nabla u)(\tau)\|_{L^{2}}d\tau.$
The linear term decays as
$\|e^{\nu\Lambda_{1}^{2s}t}u_{0,2}\|_{L^{2}}$ $\leq C(1+\nu t)^{-\frac{\sigma+1}{2s}}\big(\|\Lambda_{1}^{-\sigma}\psi_{0}\|_{L^{2}}+\|u_{0,2}\|_{L^{2}}\big)$
$\leq C\varepsilon(1+\nu t)^{-\frac{\sigma+1}{2s}}\leq\frac{C_{0}}{4}\varepsilon(1+\nu t)^{-\frac{3\sigma}{2s}}.$ (B.1)
For the nonlinear term, we split it into two parts:
$\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}\Delta^{-1}\nabla\times(u\cdot\nabla u)(\tau)\|_{L^{2}}d\tau$
$=$ $\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}\Delta^{-1}\nabla\times\mbox{div}\ (u\otimes u)(\tau)\|_{L^{2}}d\tau$
$\leq$ $\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}\Delta^{-1}\nabla\times\partial_{1}(u_{1}u)(\tau)\|_{L^{2}}d\tau$
$+\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}\Delta^{-1}\nabla\times\partial_{2}(u_{2}u)(\tau)\|_{L^{2}}d\tau$
$:=$ $II_{1}+II_{2}.$
For $II_{1}$, by Lemma 2.7, we take out $\Lambda_{1}^{1+\sigma}$ and obtain
$II_{1}\leq\int_{0}^{t-1}(t-\tau)^{-\frac{\sigma+1}{2s}}\|\Lambda_{1}^{1-\sigma}\Delta^{-1}\nabla\times(u_{1}u)\|_{L^{2}}d\tau+\int_{t-1}^{t}\|\partial_{1}(u_{1}u)(\tau)\|_{L^{2}}d\tau:=II_{11}+II_{12}.\vskip-5.69046pt$
Using Plancherel’s theorem, we can deduce that
$II_{11}\leq$ $\int_{0}^{t-1}(t-\tau)^{-\frac{\sigma+1}{2s}}\|\Lambda^{-1}\Lambda^{\frac{1}{2}-\sigma}\Lambda^{\sigma-\frac{1}{2}}\Lambda_{1}^{1-\sigma}(u_{1}u)\|_{L^{2}}d\tau$
$\leq$ $C\int_{0}^{t-1}(t-\tau)^{-\frac{\sigma+1}{2s}}\|\Lambda^{-\frac{1}{2}-\sigma}\Lambda_{1}^{\frac{1}{2}}(u_{1}u)\|_{L^{2}}d\tau.$
Then applying Lemma 2.4, we obtain
$\int_{0}^{t-1}(t-\tau)^{-\frac{\sigma+1}{2s}}\|\Lambda^{-\frac{1}{2}-\sigma}\Lambda_{1}^{\frac{1}{2}}(u_{1}u)\|_{L^{2}}d\tau$
$\leq$ $C\int_{0}^{t-1}(t-\tau)^{-\frac{\sigma+1}{2s}}\Big(\|\Lambda^{-\frac{1}{2}-\sigma}(\Lambda_{1}^{\frac{1}{2}}u_{1})u\|_{L^{2}}+\|\Lambda^{-\frac{1}{2}-\sigma}u_{1}(\Lambda_{1}^{\frac{1}{2}}u)\|_{L^{2}}\Big)d\tau.$
Using Lemma 2.3, we deduce
$\int_{0}^{t-1}(t-\tau)^{-\frac{\sigma+1}{2s}}\Big(\|\Lambda^{-\frac{1}{2}-\sigma}(\Lambda_{1}^{\frac{1}{2}}u_{1})u\|_{L^{2}}+\|\Lambda^{-\frac{1}{2}-\sigma}u_{1}(\Lambda_{1}^{\frac{1}{2}}u)\|_{L^{2}}\Big)d\tau$
$\leq$ $C\int_{0}^{t-1}(t-\tau)^{-\frac{\sigma+1}{2s}}\|\Lambda^{\frac{1}{2}-\sigma}u\|_{L^{2}}\|\Lambda_{1}^{\frac{1}{2}}u\|_{L^{2}}d\tau.$
Then by interpolation, we arrive
$II_{11}\leq$ $C\int_{0}^{t-1}(t-\tau)^{-\frac{\sigma+1}{2s}}\|\Lambda^{\frac{1}{2}-\sigma}u\|_{L^{2}}\|\Lambda_{1}^{\frac{1}{2}}u\|_{L^{2}}d\tau$
$\leq$ $C\int_{0}^{t-1}(t-\tau)^{-\frac{\sigma+1}{2s}}\|u\|_{L^{2}}^{\frac{1}{2}+\sigma}\|\nabla u\|_{L^{2}}^{\frac{1}{2}-\sigma}\|u\|_{L^{2}}^{\frac{1}{2}}\|\partial_{1}u\|_{L^{2}}^{\frac{1}{2}}d\tau$
$\leq$ $CC_{0}^{2}\varepsilon^{2}\int_{0}^{t-1}(t-\tau)^{-\frac{\sigma+1}{2s}}(\tau+1)^{-\frac{4\sigma+1}{4s}}d\tau.$ (B.2)
By Lemma 2.2, we have
$II_{12}\leq$ $C\int_{t-1}^{t}\|\partial_{1}u\|_{L^{2}}^{\frac{1}{2}}\|\partial_{1}\partial_{2}u\|_{L^{2}}^{\frac{1}{2}}\|u\|_{L^{2}}^{\frac{1}{2}}\|\partial_{1}u\|_{L^{2}}^{\frac{1}{2}}d\tau.$
Then by (4.1), (4.2) and (4.5), we have
$II_{12}\leq$ $CC_{0}^{2}\varepsilon^{2}\int_{t-1}^{t}(1+\tau)^{-\frac{3\sigma+2}{4s}}d\tau\leq CC_{0}^{2}\varepsilon^{2}(1+t)^{-\frac{\sigma+1}{2s}}.$ (B.3)
Then we estimate estimate $II_{2}$,
$II_{2}\leq$ $\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}(u_{2}u)(\tau)\|_{L^{2}}d\tau$
$\leq$ $\int_{0}^{t-1}(t-\tau)^{-\frac{3}{4s}}\|(u_{2}u)(\tau)\|_{L_{x_{1}}^{1}L_{x_{2}}^{2}}d\tau+\int_{t-1}^{t}\|\partial_{1}(u_{2}u)(\tau)\|_{L^{2}}d\tau$
$:=$ $II_{21}+II_{22}.$ (B.4)
By Lemma 2.1, and combining (4.2), (4.3), and (4.5), we have
$II_{21}\leq$ $C\int_{0}^{t-1}(t-\tau)^{-\frac{3}{4s}}\|u_{2}\|_{L^{2}}^{\frac{1}{2}}\|\partial_{2}u_{2}\|_{L^{2}}^{\frac{1}{2}}\|u\|_{L^{2}}d\tau$
$\leq$ $CC_{0}^{2}\varepsilon^{2}\int_{0}^{t-1}(t-\tau)^{-\frac{3}{4s}}(\tau+1)^{-\frac{10\sigma+3}{8s}}d\tau.$ (B.5)
Since $II_{22}\leq C\int_{t-1}^{t}\|(\partial_{1}u)u\|_{L^{2}}d\tau$, by a similar estimate as (B.4), we can conclude
$II_{22}\leq$ $CC_{0}^{2}\varepsilon^{2}\int_{t-1}^{t}(1+\tau)^{-\frac{3\sigma+2}{4s}}d\tau\leq CC_{0}^{2}\varepsilon^{2}(1+t)^{-\frac{\sigma+1}{2s}}.$ (B.6)
Combining (B.2), (B.3), (B.5) and (B.6), we have
$II\leq$ $CC_{0}^{2}\varepsilon^{2}\int_{0}^{t-1}(t-\tau)^{-\frac{\sigma+1}{2s}}(\tau+1)^{-\frac{4\sigma+1}{4s}}d\tau+CC_{0}^{2}\varepsilon^{2}\int_{0}^{t-1}(t-\tau)^{-\frac{3}{4s}}(\tau+1)^{-\frac{10\sigma+3}{8s}}d\tau$
$+CC_{0}^{2}\varepsilon^{2}(1+t)^{-\frac{\sigma+1}{2s}}.$
By applying Lemma 2.8, we have
$II\leq CC_{0}^{2}\varepsilon^{2}(1+t)^{-\frac{4\sigma+1}{4s}}+CC_{0}^{2}\varepsilon^{2}(1+t)^{-\frac{3}{4s}}+CC_{0}^{2}\varepsilon^{2}(1+t)^{-\frac{\sigma+1}{2s}}.$ (B.7)
Combining (B.1) and (B.7), we have
$\|u_{2}(t)\|_{L^{2}}\leq\left(\frac{C_{0}}{4}+4CC_{0}^{2}\varepsilon\right)\varepsilon(1+t)^{-\frac{4\sigma+1}{4s}}.$
By taking $\varepsilon$ sufficiently small, we can deduce (1.5).
Acknowledgments
J. Wu was partially supported by the National Science Foundation of the United States (Grant No. DMS 2104682 and DMS 2309748). N. Zhu was partially supported by the National Natural Science Foundation of China (Grant No. 12301285, No. 12171010) and project ZR2023QA002 supported by Shandong Provincial Natural Science Foundation.

### v2

@@
 The stability problem for the 2D Navier-Stokes equations with dissipation in only one direction on $\mathbb{R}^{2}$ is not fully understood.
-This dissipation is in the intermediate regime between the fully dissipative Navier-Stokes and the inviscid Euler.
+This dissipation is in the intermediate regime between the fully dissipative Navier-Stokes and the invisicd Euler.
 Navier-Stokes solutions in $\mathbb{R}^{2}$ decay algebraically in time while Euler solutions can grow rather rapidly in time.
@@
 is divergence-free. Suppose that $u_{0}$ admits a stream function $\psi_{0}$ such that
-$u_{0}=(u_{0,1},u_{0,2})=\nabla^{\perp}\psi_{0}:=(-\partial_{2}\psi_{0},\partial_{1}\psi_{0}),$
+$u_{0}=(u_{1,0},u_{2,0})=\nabla^{\perp}\psi_{0}:=(-\partial_{2}\psi_{0},\partial_{1}\psi_{0}),$
 and that there exists a sufficiently small constant $\varepsilon>0$ such that
-$\|u_{0}\|_{H^{k}}+\|\Lambda_{1}^{-\sigma}\psi_{0}\|_{L^{2}}+\|\Lambda_{1}^{-\sigma}u_{0,1}\|_{L^{2}}+\|\Lambda_{1}^{-\sigma}\partial_{2}u_{0,1}\|_{L^{2}}\leq\varepsilon.$ (1.4)
+$\|u_{0}\|_{H^{k}}+\|\Lambda_{1}^{-\sigma}\psi_{0}\|_{L^{2}}+\|\Lambda_{1}^{-\sigma}u_{1,0}\|_{L^{2}}+\|\Lambda_{1}^{-\sigma}\partial_{2}u_{1,0}\|_{L^{2}}\leq\varepsilon.$ (1.4)
 Then the system admits a unique global-in-time solution
@@
 satisfying
-$\|\Lambda_{1}^{-\sigma}u(t)\|_{H^{k}}+\nu^{\frac{1}{2}}\left(\int_{0}^{t}\|\Lambda_{1}^{s-\sigma}u(\tau)\|_{H^{k}}^{2}\,d\tau\right)^{\frac{1}{2}}\leq C\varepsilon,\qquad\forall\,t\geq 0,$
+$\|u(t)\|_{H^{k}}+\nu^{\frac{1}{2}}\left(\int_{0}^{t}\|\Lambda_{1}^{s}u(\tau)\|_{H^{k}}^{2}\,d\tau\right)^{\frac{1}{2}}\leq C\varepsilon,\quad\|\Lambda_{1}^{-\sigma}u\|_{L^{2}}\leq C\varepsilon,\qquad\forall\,t\geq 0,$
 together with the decay estimates
@@
 The enhanced decay of $u_{2}$ plays a crucial role in closing the nonlinear estimates.
-Moreover, since the Navier-Stokes equations do not propagate negative Sobolev regularity of the form $\Lambda^{-\sigma}u$ for $\sigma\geq\tfrac{1}{2}$, the restriction $\sigma<\tfrac{1}{2}$ is necessary.
+Moreover, since the Navier-Stokes equations do not propagate negative Sobolev regularity of the form $\Lambda_{1}^{-\sigma}u$ for $\sigma\geq\tfrac{1}{2}$, the restriction $\sigma<\tfrac{1}{2}$ is necessary.
 Remark 1.3.
@@
 $\|\partial_{1}u_{2}(t)\|_{L^{2}}$ $\leq C_{0}\varepsilon(1+t)^{-(2\sigma+1)/2s}.$
+Once again, these estimates shows that the second velocity component $u_{2}$ exhibits a faster decay rate than $u_{1}$, reflecting the underlying anisotropic structure of the system.
 The derivation of the above estimates parallels that of Theorem 1.2, we only present the estimates for $u_{2}$, which differs slightly, in Appendix B.
-Once again, these estimates shows that the second velocity component $u_{2}$ exhibits a faster decay rate than $u_{1}$, reflecting the underlying anisotropic structure of the system.
 The decay rate stated in Theorem 1.2 is not sufficiently fast to treat the range $s>\tfrac{11}{12}$.
@@
 satisfying
-$\|\Lambda_{1}^{-\sigma}u(t)\|_{H^{k}}+\nu^{1/2}\left(\int_{0}^{t}\|\Lambda_{1}^{s-\sigma}u(\tau)\|_{H^{k}}^{2}\,d\tau\right)^{1/2}<\varepsilon,\qquad\forall\,t\geq 0.$
+$\|u(t)\|_{H^{k}}+\nu^{1/2}\left(\int_{0}^{t}\|\Lambda_{1}^{s}u(\tau)\|_{H^{k}}^{2}\,d\tau\right)^{1/2}\leq C\varepsilon,\quad\|\Lambda_{1}^{-\sigma}u\|_{L^{2}}\leq C\epsilon,\qquad\forall\,t\geq 0.$
 Moreover, the solution enjoys the following weighted decay estimates:
@@
 Definition 2.12 ($A_{p}$ weights).
-A weight function $d\colon\mathbb{R}^{d}\to[0,\infty)$ is said to belong to the $A_{p}$ class if:
+A weight function $w\colon\mathbb{R}^{d}\to[0,\infty)$ is said to belong to the $A_{p}$ class if:
   1. 1.
-    $d\in L_{\mathrm{loc}}^{1}(\mathbb{R}^{d})$ and $d(x)>0$ almost everywhere;
+    $w\in L_{\mathrm{loc}}^{1}(\mathbb{R}^{d})$ and $w(x)>0$ almost everywhere;
   2. 2.
     there exists a constant $C>0$ such that for every cube $Q\subset\mathbb{R}^{d}$, the $A_{p}$ condition
-    $\left(\frac{1}{|Q|}\int_{Q}d(y)\,dy\right)\left(\frac{1}{|Q|}\int_{Q}d(y)^{-\frac{1}{p-1}}\,dy\right)^{p-1}\leq C,\qquad 1<p<\infty,$
+    $\left(\frac{1}{|Q|}\int_{Q}w(y)\,dy\right)\left(\frac{1}{|Q|}\int_{Q}w(y)^{-\frac{1}{p-1}}\,dy\right)^{p-1}\leq C,\qquad 1<p<\infty,$
     holds, where $C$ is independent of $Q$.
@@
 where $N_{1}$ corresponds to the case $m=1$ in (3.4),
-$N_{1}:=-2\sum_{i,j=1}^{2}\sum_{\alpha=1}^{k}\binom{k}{\alpha}\int_{\mathbb{R}^{2}}\partial_{1}^{\alpha}u_{j}\,\partial_{j}\partial_{1}^{k-\alpha}u_{i}\,\partial_{1}^{k}u_{i}\,dx,$
+$N_{1}:=-2\sum_{i,j=1}^{2}\sum_{\alpha=1}^{k}\binom{k}{\alpha}\int_{\mathbb{R}^{2}}\partial_{1}^{\alpha}u_{j}\,\partial_{j}\partial_{1}^{k-\alpha}u_{i}\,\partial_{1}^{k}u_{i}\,dx,$ (3.5)
 $N_{2}$ corresponds to the case $m=2$ and $j=1$,
-$N_{2}:=-2\sum_{i=1}^{2}\sum_{\alpha=1}^{k}\binom{k}{\alpha}\int_{\mathbb{R}^{2}}\partial_{2}^{\alpha}u_{1}\,\partial_{1}\partial_{2}^{k-\alpha}u_{i}\,\partial_{2}^{k}u_{i}\,dx,$
+$N_{2}:=-2\sum_{i=1}^{2}\sum_{\alpha=1}^{k}\binom{k}{\alpha}\int_{\mathbb{R}^{2}}\partial_{2}^{\alpha}u_{1}\,\partial_{1}\partial_{2}^{k-\alpha}u_{i}\,\partial_{2}^{k}u_{i}\,dx,$ (3.6)
 and $N_{3}$ corresponds to the case $m=2$ and $j=2$,
 $N_{3}:=$ $-2\sum_{i=1}^{2}\sum_{\alpha=1}^{k}\binom{k}{\alpha}\int_{\mathbb{R}^{2}}\partial_{2}^{\alpha}u_{2}\,\partial_{2}\partial_{2}^{k-\alpha}u_{i}\,\partial_{2}^{k}u_{i}\,dx$
-$=$  $\;2\sum_{i=1}^{2}\sum_{\alpha=1}^{k}\binom{k}{\alpha}\int_{\mathbb{R}^{2}}\partial_{1}\partial_{2}^{\alpha-1}u_{1}\,\partial_{2}\partial_{2}^{k-\alpha}u_{i}\,\partial_{2}^{k}u_{i}\,dx,$
+$=$  $\;2\sum_{i=1}^{2}\sum_{\alpha=1}^{k}\binom{k}{\alpha}\int_{\mathbb{R}^{2}}\partial_{1}\partial_{2}^{\alpha-1}u_{1}\,\partial_{2}\partial_{2}^{k-\alpha}u_{i}\,\partial_{2}^{k}u_{i}\,dx,$ (3.7)
 where, in the last step, we have used the incompressibility condition $\partial_{2}u_{2}=-\partial_{1}u_{1}$.
@@
 Estimate for the high-frequency part:
-For each $\ell\in\{1,2\}$, we decompose the analysis of $F^{III}_{H}$ into two parts.
+We decompose the analysis of $F^{III}_{H}$ into two parts according to the cases $1\leq\beta\leq k-1$ and $\beta=k$.
 $F^{III}_{H}=$ $\sum_{\beta=1}^{k-1}\int_{\mathbb{R}^{2}}\varphi(\xi_{1})\mathcal{F}_{x_{1}}\big(\partial^{\beta}_{\ell}f\partial^{k}_{\ell}f\big)(\xi_{1},x_{2})\mathcal{F}_{x_{1}}\big(\partial_{1}\partial^{k-\beta}_{\ell}f\big)(\xi_{1},x_{2})d\xi_{1}dx_{2}$
-$+\sum_{\beta=1}^{k-1}\int_{\mathbb{R}^{2}}\varphi(\xi_{1})\mathcal{F}_{x_{1}}\big(\partial^{k}_{\ell}f\partial^{k}_{\ell}f\big)(\xi_{1},x_{2})\mathcal{F}_{x_{1}}\big(\partial_{1}f\big)(\xi_{1},x_{2})d\xi_{1}dx_{2}$
+$+\int_{\mathbb{R}^{2}}\varphi(\xi_{1})\mathcal{F}_{x_{1}}\big(\partial^{k}_{\ell}f\partial^{k}_{\ell}f\big)(\xi_{1},x_{2})\mathcal{F}_{x_{1}}\big(\partial_{1}f\big)(\xi_{1},x_{2})d\xi_{1}dx_{2}$
 $:=$ $F^{III}_{H1}+F^{III}_{H2}.$
@@
 4 Proof of Theorem 1.2
-In this section, we prove Theorem 1.2. The argument is based on a bootstrap approach combined with refined energy estimates that yield global-in-time decay and stability.
+In this section, we prove Theorem 1.2. The argument couples two ingredients: the $H^{k}$ stability estimate for $u$ and the large-time decay estimates for $u$ and its derivatives.
+More precisely, by imposing a mild negative Sobolev regularity assumption on the initial data in the horizontal direction, we are able to derive algebraic decay rates for $u$ and $\partial_{1}u$ in $L^{2}$.
+At the same time, we also prove the propagation of the negative horizontal regularity $\|\Lambda_{1}^{-\sigma}u\|_{L^{2}}$.
+These decay rates, in turn, render the time integrals of the nonlinear terms arising in the high-order energy estimate convergent, which is the key difficulty when $s>\frac{3}{4}$.
+Indeed, for $s\leq\frac{3}{4}$ the stability of (1.1) can be established by energy estimates alone (Theorem 1.1), whereas for $s>\frac{3}{4}$ the dissipation $\Lambda_{1}^{2s}$ is no longer strong enough to directly absorb certain nonlinear terms in the $H^{k}$ estimate.
+The decay of $\|u_{1}\|_{L^{2}}$, $\|\partial_{1}u_{1}\|_{L^{2}}$, $\|\partial_{2}u_{1}\|_{L^{2}}$, $\|u_{2}\|_{L^{2}}$, and $\|\partial_{1}u_{2}\|_{L^{2}}$ is established via Duhamel’s principle, where a key structural observation is that $u_{2}$ decays strictly faster than $u_{1}$ owing to the incompressibility constraint and the stream-function representation.
+The two estimates are then closed simultaneously through a standard bootstrap argument.
 Bootstrap assumptions. We assume that there exists a constant $C_{0}>0$ such that, for all $t\in[0,T)$, the solution satisfies the following uniform bounds:
-$\|\Lambda_{1}^{-\sigma}u(t)\|_{H^{k}}$ $\leq C_{0}\varepsilon,$ (4.1)
+$\|u(t)\|_{H^{k}}$ $\leq C_{0}\varepsilon,$ (4.1)
 $\|u_{1}(t)\|_{L^{2}}$ $\leq C_{0}\varepsilon(1+t)^{-\sigma/2s},$ (4.2)
@@
 Improved estimates. Based on the initial condition (1.4) and the above a priori bounds, we will establish the improved estimates
-$\|\Lambda_{1}^{-\sigma}u(t)\|_{H^{k}}$ $\leq\tfrac{1}{2}C_{0}\varepsilon,$ (4.7)
+$\|u(t)\|_{H^{k}}$ $\leq\tfrac{1}{2}C_{0}\varepsilon,$ (4.7)
 $\|u_{1}(t)\|_{L^{2}}$ $\leq\tfrac{1}{2}C_{0}\varepsilon(1+t)^{-\sigma/2s},$ (4.8)
@@
 Once (4.7)–(4.12) are established, a standard bootstrap argument implies that the maximal existence time satisfies $T=+\infty$, and hence all decay estimates hold globally in time.
-4.1 Estimate for $\|\Lambda_{1}^{-\sigma}u\|_{L^{2}}$
-We establish the estimate for $\|\Lambda_{1}^{-\sigma}u\|_{L^{2}}$ through the following energy argument.
-Applying $\Lambda_{1}^{-\sigma}$ to the equation and taking the $L^{2}$ inner product with $\Lambda_{1}^{-\sigma}u$, we obtain
-$\frac{1}{2}\frac{d}{dt}\|\Lambda_{1}^{-\sigma}u\|_{L^{2}}^{2}+\nu\|\Lambda_{1}^{s-\sigma}u\|_{L^{2}}^{2}=-\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}(u\cdot\nabla u)\cdot\Lambda_{1}^{-\sigma}u\,dx.$ (4.13)
-The right-hand side of (4.13) can be decomposed into two parts and estimated as
-$\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}(u\cdot\nabla u)\cdot\Lambda_{1}^{-\sigma}u\,dx=$ $\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}(u_{1}\partial_{1}u)\cdot\Lambda_{1}^{-\sigma}u\,dx+\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}(u_{2}\partial_{2}u)\cdot\Lambda_{1}^{-\sigma}u\,dx$
-$\leq$ $C\left(\left\|\Lambda_{1}^{-\sigma}(u_{1}\partial_{1}u)\right\|_{L^{2}}+\left\|\Lambda_{1}^{-\sigma}(u_{2}\partial_{2}u)\right\|_{L^{2}}\right)\left\|\Lambda_{1}^{-\sigma}u\right\|_{L^{2}}$
-$:=$ $M_{0}.$
-Integrating in time on equation (4.13), we obtain
-$\|\Lambda_{1}^{-\sigma}u(t)\|_{L^{2}}^{2}+2\nu\int_{0}^{t}\|\Lambda_{1}^{s-\sigma}u(\tau)\|_{L^{2}}^{2}d\tau=\|\Lambda_{1}^{-\sigma}u_{0}\|_{L^{2}}^{2}+2\int_{0}^{t}M_{0}(\tau)d\tau.$ (4.14)
-By Lemma 2.6,
-$M_{0}$ $\leq C\Big(\left\|\partial_{2}u_{2}\right\|_{L^{2}}^{\frac{1}{2}}\left\|u_{2}\right\|_{L^{2}}^{\sigma}\left\|\partial_{1}u_{2}\right\|_{L^{2}}^{\frac{1}{2}-\sigma}\left\|\partial_{2}u\right\|_{L^{2}}$
-$\quad+\left\|\partial_{2}u_{1}\right\|_{L^{2}}^{\frac{1}{2}}\left\|u_{1}\right\|_{L^{2}}^{\sigma}\left\|\partial_{1}u_{1}\right\|_{L^{2}}^{\frac{1}{2}-\sigma}\left\|\partial_{1}u\right\|_{L^{2}}\Big)\left\|\Lambda_{1}^{-\sigma}u\right\|_{L^{2}}.$
-Using the divergence-free condition of $u$, we have
-$M_{0}$ $\leq C\Big(\left\|\partial_{1}u_{1}\right\|_{L^{2}}^{\frac{1}{2}}\left\|u_{2}\right\|_{L^{2}}^{\sigma}\left\|\partial_{1}u_{2}\right\|_{L^{2}}^{\frac{1}{2}-\sigma}\left\|\partial_{2}u\right\|_{L^{2}}$
-$\quad+\left\|\partial_{2}u_{1}\right\|_{L^{2}}^{\frac{1}{2}}\left\|u_{1}\right\|_{L^{2}}^{\sigma}\left\|\partial_{1}u_{1}\right\|_{L^{2}}^{\frac{1}{2}-\sigma}\left\|\partial_{1}u\right\|_{L^{2}}\Big)\left\|\Lambda_{1}^{-\sigma}u\right\|_{L^{2}}.$ (4.15)
-By substituting (4.2), (4.3), (4.4), (4.5) and (4.6) into (4.15), we deduce
-$M_{0}\leq CC_{0}^{3}\varepsilon^{3}\big((1+t)^{-\frac{-18\sigma^{2}+25\sigma+3}{12s}}+(1+t)^{-\frac{2\sigma+1}{2s}}\big).$
-It can be verified that
-$s<\tfrac{5}{12}+\sigma<\tfrac{-18\sigma^{2}+25\sigma+3}{12}.$
-As a consequence, we obtain
-$\int_{0}^{t}M_{0}(\tau)\,d\tau\leq CC_{0}^{3}\varepsilon^{3}.$
-Combining this bound with the assumption $\|\Lambda_{1}^{-\sigma}u_{0}\|_{L^{2}}\leq\varepsilon$, it follows from (4.14) that
-$\|\Lambda_{1}^{-\sigma}u(t)\|_{L^{2}}^{2}\leq\varepsilon^{2}+2CC_{0}^{3}\varepsilon^{3}.$
-For $\varepsilon>0$ sufficiently small so that $2CC_{0}^{3}\varepsilon\leq\tfrac{1}{2}$, we further deduce
-$\|\Lambda_{1}^{-\sigma}u(t)\|_{L^{2}}\leq\sqrt{\tfrac{3}{2}}\,\varepsilon\leq\tfrac{C_{0}}{2}\,\varepsilon.$
-4.2 Estimate for $\sum_{m=1}^{2}\|\Lambda_{1}^{-\sigma}\partial_{m}^{k}u\|_{L^{2}}$
-Applying $\Lambda_{1}^{-\sigma}\partial_{m}^{k}$ to the first equation in (1.1), and then taking the $L^{2}$ inner product with $\Lambda_{1}^{-\sigma}\partial_{m}^{k}u$, we have
-$\frac{d}{dt}\sum_{m=1}^{2}\left\|\Lambda_{1}^{-\sigma}\partial_{m}^{k}u\right\|_{L^{2}}^{2}+2\nu\sum_{m=1}^{2}\left\|\Lambda_{1}^{-\sigma}\Lambda_{1}^{s}\partial_{m}^{k}u\right\|_{L^{2}}^{2}$
-$=$ $-2\sum_{m=1}^{2}\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}\partial_{m}^{k}(u\cdot\nabla u)\cdot\Lambda_{1}^{-\sigma}\partial_{m}^{k}u~dx:=M_{1}.$ (4.16)
-First, we expand the nonlinear term using the Leibniz rule and divide it into three parts.
-$M_{1}$ $=-2\sum_{m,i,j=1}^{2}\sum_{\alpha=1}^{k}\binom{k}{\alpha}\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}(\partial_{m}^{\alpha}u_{j}\partial_{j}\partial_{m}^{k-\alpha}u_{i})~\Lambda_{1}^{-\sigma}\partial_{m}^{k}u_{i}\,dx$
-$=M_{11}+M_{12}+M_{13},$
-where the terms $M_{11}$, $M_{12}$, and $M_{13}$ correspond to different choices of the indices $(m,j)$ and are defined as follows. The term $M_{11}$ corresponds to the case $m=1$:
-$M_{11}:=\sum_{i,j=1}^{2}\sum_{\alpha=1}^{k}\binom{k}{\alpha}\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}\big(\partial_{1}^{\alpha}u_{j}\,\partial_{j}\partial_{1}^{k-\alpha}u_{i}\big)\,\Lambda_{1}^{-\sigma}\partial_{1}^{k}u_{i}\,dx.$
-The term $M_{12}$ corresponds to the case $m=2$ and $j=1$:
-$M_{12}:=\sum_{i=1}^{2}\sum_{\alpha=1}^{k}\binom{k}{\alpha}\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}\big(\partial_{2}^{\alpha}u_{1}\,\partial_{1}\partial_{2}^{k-\alpha}u_{i}\big)\,\Lambda_{1}^{-\sigma}\partial_{2}^{k}u_{i}\,dx.$
-Finally, the term $M_{13}$ corresponds to the case $m=2$ and $j=2$:
-$M_{13}$ $:=\sum_{i=1}^{2}\sum_{\alpha=1}^{k}\binom{k}{\alpha}\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}\big(\partial_{2}^{\alpha}u_{2}\,\partial_{2}\partial_{2}^{k-\alpha}u_{i}\big)\,\Lambda_{1}^{-\sigma}\partial_{2}^{k}u_{i}\,dx$
-$=-\sum_{i=1}^{2}\sum_{\alpha=1}^{k}\binom{k}{\alpha}\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}\big(\partial_{1}\partial_{2}^{\alpha-1}u_{1}\,\partial_{2}\partial_{2}^{k-\alpha}u_{i}\big)\,\Lambda_{1}^{-\sigma}\partial_{2}^{k}u_{i}\,dx,$
-where we have used the incompressibility condition $\partial_{2}u_{2}=-\partial_{1}u_{1}$.
-To estimate each component of $M_{1}$, it suffices to establish the following lemma.
+4.1 Estimate of $\|u\|_{H^{k}}$
+In this subsection, under the assumption that $u$ and $\partial_{1}u$ enjoy the prescribed time decay, we prove that the solution $u$ satisfies a closed high-order energy estimate in the range $\frac{3}{4}<s<\frac{1}{2}+\sigma$.
+The main difficulty is the treatment of the nonlinear interactions.
+By exploiting the decay of $u$ and $\partial_{1}u$, we show that these nonlinear terms can be absorbed into the energy inequality.
+As a consequence, under the additional assumption that the initial data are sufficiently small, we complete the bootstrap argument for $\|u\|_{H^{k}}$.
+We first record the core nonlinear estimate used in the proof.
 Lemma 4.1.
-Let $1\leq\beta\leq k$ be an integer. For $\ell\in\{1,2\}$ and $\tfrac{3}{4}\leq s<\sigma+\tfrac{1}{2}$, we have
-$\sum_{\beta=1}^{k}\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}\big(\partial^{\beta}_{\ell}u\partial_{1}\partial^{k-\beta}_{\ell}u\big)\Lambda_{1}^{-\sigma}\partial^{k}_{\ell}udx\leq C\|\Lambda_{1}^{-\sigma+s}u\|_{H^{k}}^{2}\|\Lambda_{1}^{-\sigma}u\|_{H^{k}}.$
-We now proceed to estimate the term $M_{1}$. The proof of Lemma 4.1 will be given later. Assuming Lemma 4.1 for the moment, we obtain
-$|M_{1}|\leq|M_{11}|+|M_{12}|+|M_{13}|\leq C\|\Lambda_{1}^{-\sigma+s}u\|_{H^{k}}^{2}\|\Lambda_{1}^{-\sigma}u\|_{H^{k}}.$
-Substituting this bound into (4.2) and integrating in time yield
-$\sum_{m=1}^{2}\|\Lambda_{1}^{-\sigma}\partial_{m}^{k}u(t)\|_{L^{2}}^{2}+2\nu\int_{0}^{t}\sum_{m=1}^{2}\|\Lambda_{1}^{s}\Lambda_{1}^{-\sigma}\partial_{m}^{k}u(\tau)\|_{L^{2}}^{2}\,d\tau$
-$\leq{}$ $\sum_{m=1}^{2}\|\Lambda_{1}^{-\sigma}\partial_{m}^{k}u_{0}\|_{L^{2}}^{2}+C\int_{0}^{t}\|\Lambda_{1}^{-\sigma+s}u(\tau)\|_{H^{k}}^{2}\|\Lambda_{1}^{-\sigma}u(\tau)\|_{H^{k}}\,d\tau.$ (4.17)
-Combining (4.17) with the $L^{2}$ energy estimate (4.14), we arrive at
-$\|\Lambda_{1}^{-\sigma}u(t)\|_{H^{k}}^{2}$ $+2\nu\int_{0}^{t}\|\Lambda_{1}^{s}\Lambda_{1}^{-\sigma}u(\tau)\|_{H^{k}}^{2}\,d\tau$
-$\leq{}$ $\|\Lambda_{1}^{-\sigma}u_{0}\|_{H^{k}}^{2}+C\int_{0}^{t}\|\Lambda_{1}^{-\sigma+s}u(\tau)\|_{H^{k}}^{2}\|\Lambda_{1}^{-\sigma}u(\tau)\|_{H^{k}}\,d\tau.$ (4.18)
-If $\|\Lambda_{1}^{-\sigma}u_{0}\|_{H^{k}}<\varepsilon$ with $\varepsilon<\nu/C$, then a standard bootstrap argument applied to (4.18) yields
-$\|\Lambda_{1}^{-\sigma}u(t)\|_{H^{k}}^{2}+\nu\int_{0}^{t}\|\Lambda_{1}^{s}\Lambda_{1}^{-\sigma}u(\tau)\|_{H^{k}}^{2}\,d\tau\leq C\|\Lambda_{1}^{-\sigma}u_{0}\|_{H^{k}}^{2}\leq C\varepsilon^{2}.$
-We now turn to the proof of Lemma 4.1.
-Proof of Lemma 4.1. We decompose $J$ into two parts:
-$J$ $=\sum_{\beta=1}^{k}\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}\big(\partial_{\ell}^{\beta}u\,\partial_{1}\partial_{\ell}^{k-\beta}u\big)\,\Lambda_{1}^{-\sigma}\partial_{\ell}^{k}u\,dx$
-$=\sum_{\beta=1}^{k-1}\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}\big(\partial_{\ell}^{\beta}u\,\partial_{1}\partial_{\ell}^{k-\beta}u\big)\,\Lambda_{1}^{-\sigma}\partial_{\ell}^{k}u\,dx+\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}\big(\partial_{\ell}^{k}u\,\partial_{1}u\big)\,\Lambda_{1}^{-\sigma}\partial_{\ell}^{k}u\,dx$
-$=:J_{1}+J_{2}.$
-Since $s<\tfrac{1}{2}+\sigma$, we may employ Cauchy-Schwarz inequality and then apply Lemma 2.3 to estimate $J_{1}$:
-$J_{1}$ $\leq\sum_{\beta=1}^{k-1}\|\Lambda_{1}^{-\sigma}(\partial_{\ell}^{\beta}u\,\partial_{1}\partial_{\ell}^{k-\beta}u)\|_{L^{2}}\|\Lambda_{1}^{-\sigma}\partial_{\ell}^{k}u\|_{L^{2}}$
-$\leq\sum_{\beta=1}^{k-1}\|\Lambda_{1}^{-\sigma+s}\partial_{\ell}^{\beta}u\|_{L_{x_{1}}^{2}L_{x_{2}}^{\infty}}\,\|\Lambda_{1}^{\frac{1}{2}-s}\partial_{1}\partial_{\ell}^{k-\beta}u\|_{L^{2}}\,\|\Lambda_{1}^{-\sigma}\partial_{\ell}^{k}u\|_{L^{2}}$
-$\leq\sum_{\beta=1}^{k-1}\|\Lambda_{1}^{-\sigma+s}\partial_{\ell}^{\beta}u\|_{L_{x_{1}}^{2}H_{x_{2}}^{1}}\,\|\Lambda_{1}^{-\sigma+s}\Lambda_{1}^{\frac{3}{2}-2s+\sigma}\partial_{\ell}^{k-\beta}u\|_{L^{2}}\,\|\Lambda_{1}^{-\sigma}\partial_{\ell}^{k}u\|_{L^{2}}.$
-By the admissible ranges of $\beta$, $s$, and $\sigma$, we have
-$\sum_{\beta=1}^{k-1}\|\Lambda_{1}^{-\sigma+s}\partial_{\ell}^{\beta}u\|_{L_{x_{1}}^{2}H_{x_{2}}^{1}}\leq\|\Lambda_{1}^{-\sigma+s}u\|_{H^{k}},\qquad\sum_{\beta=1}^{k-1}\|\Lambda_{1}^{-\sigma+s}\Lambda_{1}^{\frac{3}{2}-2s+\sigma}\partial_{\ell}^{k-\beta}u\|_{L^{2}}\leq\|\Lambda_{1}^{-\sigma+s}u\|_{H^{k}}.$
-Consequently,
-$J_{1}\leq C\|\Lambda_{1}^{-\sigma+s}u\|_{H^{k}}^{2}\|\Lambda_{1}^{-\sigma}u\|_{H^{k}}.$ (4.19)
-The term $J_{2}$ can be estimated in a similar manner. By Hölder’s inequality and Lemma 2.3, we obtain
-$J_{2}$ $\leq\|\Lambda_{1}^{-\sigma}(\partial_{\ell}^{k}u\,\partial_{1}u)\|_{L^{2}}\|\Lambda_{1}^{-\sigma}\partial_{\ell}^{k}u\|_{L^{2}}$
-$\leq\|\Lambda_{1}^{-\sigma+s}\partial_{\ell}^{k}u\|_{L^{2}}\,\|\Lambda_{1}^{\frac{1}{2}-s}\partial_{1}u\|_{L_{x_{1}}^{2}L_{x_{2}}^{\infty}}\,\|\Lambda_{1}^{-\sigma}\partial_{\ell}^{k}u\|_{L^{2}}$
-$\leq\|\Lambda_{1}^{-\sigma+s}\partial_{\ell}^{k}u\|_{L^{2}}\,\|\Lambda_{1}^{-\sigma+s}\Lambda_{1}^{\frac{3}{2}-2s+\sigma}u\|_{L_{x_{1}}^{2}H_{x_{2}}^{1}}\,\|\Lambda_{1}^{-\sigma}\partial_{\ell}^{k}u\|_{L^{2}}.$ (4.20)
-Combining (4.19) and (4.20), we conclude that
-$J\leq C\|\Lambda_{1}^{-\sigma+s}u\|_{H^{k}}^{2}\|\Lambda_{1}^{-\sigma}u\|_{H^{k}}.$
-This completes the proof of Lemma 4.1.
+Assume that $\frac{3}{4}<s<\frac{1}{2}+\sigma$, and that $u$ and $\partial_{1}u$ satisfy the decay bounds postulated in the bootstrap assumptions.
+Then there exist constants $\vartheta>2$ and $\Omega>1$ such that
+$\sum_{\beta=1}^{k}\int_{\mathbb{R}^{2}}\partial_{2}^{\beta}u\,\partial_{1}\partial_{2}^{k-\beta}u\,\partial_{2}^{k}u\,dx\leq C(C_{0}\varepsilon)^{\vartheta}(1+t)^{-\Omega}+C\|\partial_{2}^{k}u\|_{L^{2}}\,\|\Lambda_{1}^{s}\partial_{2}^{k}u\|_{L^{2}}^{2}.$
+The proof of this lemma will be given at the end of this subsection.
+The $L^{2}$ energy estimate for $u$ is standard:
+$\|u(t)\|_{L^{2}}^{2}+2\nu\int_{0}^{t}\|\Lambda_{1}^{s}u(\tau)\|_{L^{2}}^{2}\,d\tau\leq\|u_{0}\|_{L^{2}}^{2}.$ (4.13)
+For the $\dot{H}^{k}$ norm, using the same notations as (3.3)–(3.1)), we have
+$\frac{d}{dt}\sum_{m=1}^{2}\|\partial_{m}^{k}u\|_{L^{2}}^{2}+2\nu\sum_{m=1}^{2}\|\Lambda_{1}^{s}\partial_{m}^{k}u\|_{L^{2}}^{2}=N=N_{1}+N_{2}+N_{3},$ (4.14)
+For $N_{1}$, since there are enough $\partial_{1}$ derivatives, the estimate is direct. Assume $1\leq\alpha\leq k-1$. By Hölder’s inequality and Lemma 2.3, we obtain
+$|N_{1}|$ $\leq C\sum_{i,j=1}^{2}\sum_{\alpha=1}^{k-1}\binom{k}{\alpha}\|\partial_{1}^{\alpha}u_{j}\|_{L^{2}_{x_{1}}L^{\infty}_{x_{2}}}\|\partial_{j}\partial_{1}^{k-\alpha}u_{i}\|_{L^{4}_{x_{1}}L^{2}_{x_{2}}}\|\partial_{1}^{k}u_{i}\|_{L^{4}_{x_{1}}L^{2}_{x_{2}}}.$
+By Sobolev embedding in the $x_{1}$ variable,
+$\|\partial_{j}\partial_{1}^{k-\alpha}u_{i}\|_{L^{4}_{x_{1}}L^{2}_{x_{2}}}\|\partial_{1}^{k}u_{i}\|_{L^{4}_{x_{1}}L^{2}_{x_{2}}}$ $\leq\|\Lambda_{1}^{\frac{1}{4}}\partial_{j}\partial_{1}^{k-\alpha}u_{i}\|_{L^{2}}\|\Lambda_{1}^{\frac{1}{4}}\partial_{1}^{k}u_{i}\|_{L^{2}}$
+$\leq\|\Lambda_{1}^{s}\Lambda_{1}^{\frac{1}{4}-s}\partial_{j}\partial_{1}^{k-\alpha}u_{i}\|_{L^{2}}\|\Lambda_{1}^{s}\Lambda_{1}^{\frac{1}{4}-s}\partial_{1}^{k}u_{i}\|_{L^{2}}.$
+Substituting this into the above inequality yields
+$|N_{1}|$ $\leq C\sum_{i,j=1}^{2}\sum_{\alpha=1}^{k-1}\binom{k}{\alpha}\|\partial_{1}^{\alpha}u_{j}\|_{L^{2}_{x_{1}}L^{\infty}_{x_{2}}}\|\Lambda_{1}^{s}\Lambda_{1}^{\frac{1}{4}-s}\partial_{j}\partial_{1}^{k-\alpha}u_{i}\|_{L^{2}}\|\Lambda_{1}^{s}\Lambda_{1}^{\frac{1}{4}-s}\partial_{1}^{k}u_{i}\|_{L^{2}}$
+$\leq C\|u\|_{H^{k}}\|\Lambda_{1}^{s}u\|_{H^{k}}^{2}.$
+For the endpoint case $\alpha=k$, we similarly have
+$\sum_{i,j=1}^{2}\left|\int_{\mathbb{R}^{2}}\big(\partial_{1}^{k}u_{j}\,\partial_{j}u_{i}\big)\,\partial_{1}^{k}u_{i}\,dx\right|\leq C\|u\|_{H^{k}}\|\Lambda_{1}^{s}u\|_{H^{k}}^{2},$
+so the $\alpha=k$ contribution is controlled in the same way.
+The new difficulty, compared with Theorem 1.1, lies in $N_{2}$ and $N_{3}$: these terms involve $\partial_{1}$ acting on only one factor, so they cannot be controlled by the dissipation $\Lambda_{1}^{2s}$ alone when $s>\frac{3}{4}$.
+We handle them using the bootstrap decay bounds (4.2) and (4.5).
+Applying Lemma 4.1 to $N_{2}$, and using the identity $\partial_{2}u_{2}=-\partial_{1}u_{1}$, we obtain the same bound for $N_{3}$. Consequently,
+$|N_{2}|+|N_{3}|\leq C(C_{0}\varepsilon)^{\vartheta}(1+t)^{-\Omega}+C\|\partial_{2}^{k}u\|_{L^{2}}\,\|\Lambda_{1}^{s}\partial_{2}^{k}u\|_{L^{2}}^{2},$
+where $\vartheta>2$ and $\Omega>1$ are given by Lemma 4.1.
+Since $\|\partial_{2}^{k}u\|_{L^{2}}\leq\|u\|_{H^{k}}$ and $\|\Lambda_{1}^{s}\partial_{2}^{k}u\|_{L^{2}}\leq\|\Lambda_{1}^{s}u\|_{H^{k}}$, we infer that
+$|N|\leq C(C_{0}\varepsilon)^{\vartheta}(1+t)^{-\Omega}+C\|u\|_{H^{k}}\|\Lambda_{1}^{s}u\|_{H^{k}}^{2}.$ (4.15)
+Substituting (4.15) into (4.14) and integrating with respect to time, we get
+$\sum_{m=1}^{2}\|\partial_{m}^{k}u(t)\|_{L^{2}}^{2}+2\nu\int_{0}^{t}\sum_{m=1}^{2}\|\Lambda_{1}^{s}\partial_{m}^{k}u(\tau)\|_{L^{2}}^{2}\,d\tau$
+$\leq{}$ $\sum_{m=1}^{2}\|\partial_{m}^{k}u_{0}\|_{L^{2}}^{2}+C(C_{0}\varepsilon)^{\vartheta}+C\int_{0}^{t}\|\Lambda_{1}^{s}u(\tau)\|_{H^{k}}^{2}\|u(\tau)\|_{H^{k}}\,d\tau.$ (4.16)
+Combining (4.13) and (4.1), we conclude
+$\|u(t)\|_{H^{k}}^{2}+\nu\int_{0}^{t}\|\Lambda_{1}^{s}u(\tau)\|_{H^{k}}^{2}\,d\tau\leq C\|u_{0}\|_{H^{k}}^{2}+C(C_{0}\varepsilon)^{\vartheta}+C\int_{0}^{t}\|\Lambda_{1}^{s}u(\tau)\|_{H^{k}}^{2}\|u(\tau)\|_{H^{k}}\,d\tau.$ (4.17)
+If $\|u_{0}\|_{H^{k}}\leq\varepsilon$ with $\varepsilon$ sufficiently small, then a standard bootstrap argument yields
+$\|u(t)\|_{H^{k}}^{2}+\nu\int_{0}^{t}\|\Lambda_{1}^{s}u(\tau)\|_{H^{k}}^{2}\,d\tau\leq C\varepsilon^{2}.$
+Proof of Lemma 4.1.
+We again split the estimate into the cases $1\leq\beta\leq k-1$ and $\beta=k$, namely
+$\mathcal{N}_{\rm low}$ $:=\sum_{\beta=1}^{k-1}\int_{\mathbb{R}^{2}}\partial_{2}^{\beta}u\,\partial_{1}\partial_{2}^{k-\beta}u\,\partial_{2}^{k}u\,dx,$
+$\mathcal{N}_{\rm top}$ $:=\int_{\mathbb{R}^{2}}\partial_{2}^{k}u\,\partial_{1}u\,\partial_{2}^{k}u\,dx.$
+For the lower-order part $1\leq\beta\leq k-1$, by Hölder’s inequality,
+$|\mathcal{N}_{\rm low}|\leq C\sum_{\beta=1}^{k-1}\|\Lambda_{1}^{s}\partial_{2}^{k-\beta}u\|_{L^{2}}\|\Lambda_{1}^{1-s}\big((\partial_{2}^{\beta}u)(\partial_{2}^{k}u)\big)\|_{L^{2}}.$
+For the first factor, interpolation in the $x_{2}$ direction gives
+$\|\Lambda_{1}^{s}\partial_{2}^{k-\beta}u\|_{L^{2}}\leq\|\Lambda_{1}^{s}u\|_{L^{2}}^{\frac{\beta}{k}}\|\Lambda_{1}^{s}\partial_{2}^{k}u\|_{L^{2}}^{1-\frac{\beta}{k}}.$
+Using interpolation in the $x_{1}$ direction,
+$\|\Lambda_{1}^{s}u\|_{L^{2}}\leq\|u\|_{L^{2}}^{1-s}\|\partial_{1}u\|_{L^{2}}^{s},$
+hence
+$\|\Lambda_{1}^{s}\partial_{2}^{k-\beta}u\|_{L^{2}}\leq\Big(\|u\|_{L^{2}}^{1-s}\|\partial_{1}u\|_{L^{2}}^{s}\Big)^{\frac{\beta}{k}}\|\Lambda_{1}^{s}\partial_{2}^{k}u\|_{L^{2}}^{1-\frac{\beta}{k}}.$
+For the second factor, by Lemma 2.3,
+$\|\Lambda_{1}^{1-s}[(\partial_{2}^{\beta}u)(\partial_{2}^{k}u)]\|_{L^{2}}\leq\|\Lambda_{1}^{\frac{3}{4}-\frac{s}{2}}\partial_{2}^{\beta}u\|_{L^{2}_{x_{1}}L^{\infty}_{x_{2}}}\|\Lambda_{1}^{\frac{3}{4}-\frac{s}{2}}\partial_{2}^{k}u\|_{L^{2}}.$
+By Gagliardo–Nirenberg inequality,
+$\|\Lambda_{1}^{\frac{3}{4}-\frac{s}{2}}\partial_{2}^{\beta}u\|_{L^{2}_{x_{1}}L^{\infty}_{x_{2}}}\leq\|\Lambda_{1}^{\frac{3}{4}-\frac{s}{2}}\partial_{2}^{\beta}u\|_{L^{2}}^{\frac{1}{2}}\|\Lambda_{1}^{\frac{3}{4}-\frac{s}{2}}\partial_{2}^{\beta+1}u\|_{L^{2}}^{\frac{1}{2}}.$
+Interpolating the $\partial_{2}$ derivatives between order 0 and $k$, we obtain
+$\|\Lambda_{1}^{\frac{3}{4}-\frac{s}{2}}\partial_{2}^{\beta}u\|_{L^{2}}$ $\leq\|\Lambda_{1}^{\frac{3}{4}-\frac{s}{2}}u\|_{L^{2}}^{1-\frac{\beta}{k}}\|\Lambda_{1}^{\frac{3}{4}-\frac{s}{2}}\partial_{2}^{k}u\|_{L^{2}}^{\frac{\beta}{k}},$
+$\|\Lambda_{1}^{\frac{3}{4}-\frac{s}{2}}\partial_{2}^{\beta+1}u\|_{L^{2}}$ $\leq\|\Lambda_{1}^{\frac{3}{4}-\frac{s}{2}}u\|_{L^{2}}^{1-\frac{\beta+1}{k}}\|\Lambda_{1}^{\frac{3}{4}-\frac{s}{2}}\partial_{2}^{k}u\|_{L^{2}}^{\frac{\beta+1}{k}}.$
+Therefore,
+$\|\Lambda_{1}^{\frac{3}{4}-\frac{s}{2}}\partial_{2}^{\beta}u\|_{L^{2}_{x_{1}}L^{\infty}_{x_{2}}}\leq\|\Lambda_{1}^{\frac{3}{4}-\frac{s}{2}}u\|_{L^{2}}^{1-\frac{2\beta+1}{2k}}\|\Lambda_{1}^{\frac{3}{4}-\frac{s}{2}}\partial_{2}^{k}u\|_{L^{2}}^{\frac{2\beta+1}{2k}}.$
+Since $\frac{3}{4}<s<\frac{11}{12}$, we have $\frac{7}{24}<\frac{3}{4}-\frac{s}{2}<\frac{3}{8}<1$, and thus
+$\|\Lambda_{1}^{\frac{3}{4}-\frac{s}{2}}u\|_{L^{2}}\leq\|u\|_{L^{2}}^{1-(\frac{3}{4}-\frac{s}{2})}\|\partial_{1}u\|_{L^{2}}^{\frac{3}{4}-\frac{s}{2}}.$
+Also,
+$\|\Lambda_{1}^{\frac{3}{4}-\frac{s}{2}}\partial_{2}^{k}u\|_{L^{2}}\leq\|\partial_{2}^{k}u\|_{L^{2}}^{1-\theta_{1}}\|\Lambda_{1}^{s}\partial_{2}^{k}u\|_{L^{2}}^{\theta_{1}},\qquad\theta_{1}=\frac{3-2s}{4s}.$
+Collecting the above estimates, we obtain —N_low—≤C ∥ u ∥_L^2^γ_1 ∥ ∂_1 u ∥_L^2^γ_2 ∥ ∂_2^k u ∥_L^2^γ_3 ∥ Λ_1^s ∂_2^k u ∥_L^2^γ_4, where γ_1 = β(1-s)k + ( 14 + s2 ) ( 1 - 2β+12k ), γ_2 = βsk + ( 34 - s2 ) ( 1 - 2β+12k ), γ_3 = 6s-34s ( 1 + 2β+12k ), γ_4 = 1 - βk + 3-2s4s ( 1 + 2β+12k ).
+By Young’s inequality, we further get
+$|\mathcal{N}_{\rm low}|\leq{}$ $C\bigg[\Big(\|u\|_{L^{2}}^{\gamma_{1}}\|\partial_{1}u\|_{L^{2}}^{\gamma_{2}}\|\partial_{2}^{k}u\|_{L^{2}}^{\gamma_{3}-\gamma_{4}/2}\Big)^{\frac{2}{2-\gamma_{4}}}+C\|\partial_{2}^{k}u\|_{L^{2}}\,\|\Lambda_{1}^{s}\partial_{2}^{k}u\|_{L^{2}}^{2}\bigg].$
+Combining the decay rates (4.1), (4.2) and (4.5), we obtain
+$|\mathcal{N}_{\rm low}|\leq C(C_{0}\varepsilon)^{\frac{2(\gamma_{1}+\gamma_{2}+\gamma_{3}-\gamma_{4}/2)}{2-\gamma_{4}}}(1+t)^{-\Omega_{1}}+C\|\partial_{2}^{k}u\|_{L^{2}}\,\|\Lambda_{1}^{s}\partial_{2}^{k}u\|_{L^{2}}^{2},$
+where Ω_1 = γ1σ+ γ2(σ+ 1)s(2-γ4).
+By direct computation, we get 1 - Ω_1 = 4k(4s - 3) - 4σ(2k - 1)12ks + 12βs - 6k - 6β+ 2s - 3.
+When $k>2$, the denominator is positive.
+Therefore, $1-\Omega_{1}<0$ iff the numerator is negative, namely 4k(4s - 3) - 4σ(2k - 1)¡0, which is equivalent to s¡34+σ2-σ4k.
+Since k¿σ1-2σ, and $s<\frac{1}{2}+\sigma$, we obtain s¡12+σ¡34+σ2-σ4k.
+Hence $\Omega_{1}>1$.
+For the endpoint part $\beta=k$, denote it by $\mathcal{N}_{\rm top}$. By Hölder’s inequality and Sobolev embedding,
+$|\mathcal{N}_{\rm top}|$ $=\left|\int_{\mathbb{R}^{2}}\big(\partial_{2}^{k}u\,\partial_{1}u\big)\,\partial_{2}^{k}u\,dx\right|$
+$\leq C\|\partial_{1}u\|_{L^{2}_{x_{1}}L^{\infty}_{x_{2}}}\|\partial_{2}^{k}u\|_{L^{4}_{x_{1}}L^{2}_{x_{2}}}^{2}$
+$\leq C\|\partial_{1}u\|_{L^{2}}^{\frac{1}{2}}\|\partial_{1}\partial_{2}u\|_{L^{2}}^{\frac{1}{2}}\|\Lambda_{1}^{\frac{1}{4}}\partial_{2}^{k}u\|_{L^{2}}^{2}.$
+Using interpolation,
+$\|\partial_{1}\partial_{2}u\|_{L^{2}}$ $\leq\|\partial_{1}u\|_{L^{2}}^{\frac{k-1}{k}}\|\partial_{1}\partial_{2}^{k}u\|_{L^{2}}^{\frac{1}{k}},$
+$\|\Lambda_{1}^{\frac{1}{4}}\partial_{2}^{k}u\|_{L^{2}}^{2}$ $\leq C\|\partial_{2}^{k}u\|_{L^{2}}^{\frac{3}{2}}\|\partial_{1}\partial_{2}^{k}u\|_{L^{2}}^{\frac{1}{2}}.$
+Hence,
+$|\mathcal{N}_{\rm top}|\leq C\|\partial_{1}u\|_{L^{2}}^{\frac{1}{2}}\|\partial_{1}u\|_{L^{2}}^{\frac{k-1}{2k}}\|\partial_{1}\partial_{2}^{k}u\|_{L^{2}}^{\frac{1}{2k}}\|\partial_{2}^{k}u\|_{L^{2}}^{\frac{3}{2}}\|\partial_{1}\partial_{2}^{k}u\|_{L^{2}}^{\frac{1}{2}}.$
+Applying Young’s inequality, we obtain
+$|\mathcal{N}_{\rm top}|\leq{}$ $C\bigg[\left(\|\partial_{1}u\|_{L^{2}}^{\frac{2k-1}{2k}}\|\partial_{2}^{k}u\|_{L^{2}}^{\frac{5k-1}{4k}}\right)^{\frac{4k}{3k-1}}+C\|\partial_{2}^{k}u\|_{L^{2}}\|\partial_{1}\partial_{2}^{k}u\|_{L^{2}}^{2}\bigg].$
+Combining the decay and boundedness estimates (4.1), (4.2), and (4.5), we get —N_top— ≤(C C_0ε)^9k-33k-1 (1+t)^-Ω_2 + C∥∂_2^k u∥_L^2 ∥∂_1 ∂_2^k u∥_L^2^2, where Ω_2 = σ+12s ×4k-23k-1.
+By $k>\frac{1}{1-2\sigma}$ and $s<\frac{1}{2}+\sigma$, we have $s<\frac{1}{2}+\sigma<\frac{\sigma+1}{2}\times\frac{4k-2}{3k-1}$ and thus $\Omega_{2}>1.$
 4.2 The decay of $u_{1}$
@@
 For the first velocity component $u_{1}$, we take the $L^{2}$ norm and use the boundedness of the Leray projector $\mathbb{P}$ to obtain
-$\|u_{1}(t)\|_{L^{2}}$ $\leq\|e^{\nu\Lambda_{1}^{2s}t}u_{0,1}\|_{L^{2}}+C\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}(u\cdot\nabla u_{1})(\tau)\|_{L^{2}}\,d\tau$
+$\|u_{1}(t)\|_{L^{2}}$ $\leq\|e^{\nu\Lambda_{1}^{2s}t}u_{1,0}\|_{L^{2}}+C\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}(u\cdot\nabla u_{1})(\tau)\|_{L^{2}}\,d\tau$
 $\quad+C\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}(u\cdot\nabla u_{2})(\tau)\|_{L^{2}}\,d\tau.$
 The linear term admits the decay estimate
-$\|e^{\nu\Lambda_{1}^{2s}t}u_{0,1}\|_{L^{2}}$ $\leq C(1+\nu t)^{-\frac{\sigma}{2s}}\Big(\|\Lambda_{1}^{-\sigma}u_{0,1}\|_{L^{2}}+\|u_{0}\|_{L^{2}}\Big)$
+$\|e^{\nu\Lambda_{1}^{2s}t}u_{1,0}\|_{L^{2}}$ $\leq C(1+\nu t)^{-\frac{\sigma}{2s}}\Big(\|\Lambda_{1}^{-\sigma}u_{1,0}\|_{L^{2}}+\|u_{0}\|_{L^{2}}\Big)$
 $\leq C\varepsilon(1+\nu t)^{-\frac{\sigma}{2s}}\leq\frac{C_{0}}{4}\varepsilon(1+\nu t)^{-\frac{\sigma}{2s}},$
@@
 and the integral form:
-$u_{2}(t)=e^{\nu\Lambda_{1}^{2s}t}u_{0,2}-\int_{0}^{t}e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}\Delta^{-1}\nabla\times(u\cdot\nabla u)(\tau)d\tau.$
+$u_{2}(t)=e^{\nu\Lambda_{1}^{2s}t}u_{2,0}-\int_{0}^{t}e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}\Delta^{-1}\nabla\times(u\cdot\nabla u)(\tau)d\tau.$
 Taking the $L^{2}$ norm, we obtain
-$\|u_{2}(t)\|_{L^{2}}\leq\|e^{\nu\Lambda_{1}^{2s}t}u_{0,2}\|_{L^{2}}+\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}\Delta^{-1}\nabla\times(u\cdot\nabla u)(\tau)\|_{L^{2}}\,d\tau.$
+$\|u_{2}(t)\|_{L^{2}}\leq\|e^{\nu\Lambda_{1}^{2s}t}u_{2,0}\|_{L^{2}}+\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}\Delta^{-1}\nabla\times(u\cdot\nabla u)(\tau)\|_{L^{2}}\,d\tau.$
 The linear term admits the decay estimate
-$\|e^{\nu\Lambda_{1}^{2s}t}u_{0,2}\|_{L^{2}}$ $\leq C(1+\nu t)^{-\frac{\sigma+1}{2s}}\big(\|\Lambda_{1}^{-\sigma}\psi_{0}\|_{L^{2}}+\|u_{0,2}\|_{L^{2}}\big)$
+$\|e^{\nu\Lambda_{1}^{2s}t}u_{2,0}\|_{L^{2}}$ $\leq C(1+\nu t)^{-\frac{\sigma+1}{2s}}\big(\|\Lambda_{1}^{-\sigma}\psi_{0}\|_{L^{2}}+\|u_{2,0}\|_{L^{2}}\big)$
 $\leq C\varepsilon(1+\nu t)^{-\frac{\sigma+1}{2s}}\leq\frac{C_{0}}{4}\varepsilon(1+\nu t)^{-\frac{\sigma+\frac{2}{3}}{2s}},$ (4.19)
@@
 $=:{}$ $II_{1}+II_{2}.$
-For $II_{1}$, by Lemma 2.7, we take out $\Lambda_{1}^{\sigma+\frac{2}{3}}$ and applying Lemma 2.3 and Lemma 2.4, we have
+For $II_{1}$, by Lemma 2.7, we take out $\Lambda_{1}^{\sigma+\frac{2}{3}}$, and then use the fact that $\|\Lambda^{-\frac{1}{2}+\sigma}f\|_{L^{2}}\leq\|\Lambda_{1}^{-\frac{1}{2}+\sigma}f\|_{L^{2}}$ and applying Lemma 2.4, we have
 $II_{1}\leq$ $\int_{0}^{t}(t-\tau)^{-\frac{\sigma+\frac{2}{3}}{2s}}\|\Lambda^{-1}\Lambda_{1}^{\frac{4}{3}-\sigma}(u_{1}u)\|_{L^{2}}d\tau$
 $\leq$ $\int_{0}^{t}(t-\tau)^{-\frac{\sigma+\frac{2}{3}}{2s}}\|\Lambda^{-\frac{1}{2}-\sigma}\Lambda_{1}^{\frac{5}{6}}(u_{1}u)\|_{L^{2}}d\tau$
-$\leq$ $\int_{0}^{t}(t-\tau)^{-\frac{\sigma+\frac{2}{3}}{2s}}\Big(\|\Lambda^{-\frac{1}{2}-\sigma}(\Lambda_{1}^{\frac{5}{6}}u_{1})u\|_{L^{2}}+\|\Lambda^{-\frac{1}{2}-\sigma}u_{1}(\Lambda_{1}^{\frac{5}{6}}u)\|_{L^{2}}\Big)d\tau.$ (4.23)
+$\leq$ $\int_{0}^{t}(t-\tau)^{-\frac{\sigma+\frac{2}{3}}{2s}}\Big(\|\Lambda^{-\frac{1}{2}-\sigma}\Big[(\Lambda_{1}^{\frac{5}{6}}u_{1})u\Big]\|_{L^{2}}+\|\Lambda^{-\frac{1}{2}-\sigma}\Big[u_{1}(\Lambda_{1}^{\frac{5}{6}}u)\Big]\|_{L^{2}}\Big)d\tau.$ (4.20)
 By applying Lemma 2.3 with $s=-\frac{1}{2}-\sigma,s_{1}=\frac{1}{2}-\sigma,s_{2}=0$ and interpolation, we obtain
-$\int_{0}^{t}(t-\tau)^{-\frac{\sigma+\frac{2}{3}}{2s}}\Big(\|\Lambda^{-\frac{1}{2}-\sigma}(\Lambda_{1}^{\frac{5}{6}}u_{1})u\|_{L^{2}}+\|\Lambda^{-\frac{1}{2}-\sigma}u_{1}(\Lambda_{1}^{\frac{5}{6}}u)\|_{L^{2}}\Big)d\tau$
+$\int_{0}^{t}(t-\tau)^{-\frac{\sigma+\frac{2}{3}}{2s}}\Big(\|\Lambda^{-\frac{1}{2}-\sigma}\Big[(\Lambda_{1}^{\frac{5}{6}}u_{1})u\Big]\|_{L^{2}}+\|\Lambda^{-\frac{1}{2}-\sigma}\Big[u_{1}(\Lambda_{1}^{\frac{5}{6}}u)\Big]\|_{L^{2}}\Big)d\tau$
 $\leq$ $\int_{0}^{t}(t-\tau)^{-\frac{\sigma+\frac{2}{3}}{2s}}\|\Lambda^{\frac{1}{2}-\sigma}u\|_{L^{2}}\|\Lambda_{1}^{\frac{5}{6}}u\|_{L^{2}}d\tau$
@@
 By Duhamel’s principle, we have
-$\|\partial_{2}u_{1}(t)\|_{L^{2}}\leq\|e^{\nu\Lambda_{1}^{2s}t}\partial_{2}u_{0,1}\|_{L^{2}}+C\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{2}(u\cdot\nabla u)(\tau)\|_{L^{2}}d\tau.$
+$\|\partial_{2}u_{1}(t)\|_{L^{2}}\leq\|e^{\nu\Lambda_{1}^{2s}t}\partial_{2}u_{1,0}\|_{L^{2}}+C\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{2}(u\cdot\nabla u)(\tau)\|_{L^{2}}d\tau.$
 For the linear term involving initial data $\partial_{2}u_{0}$, we can obtain the decay estimate through Lemma 2.3.
-$\|e^{\nu\Lambda_{1}^{2s}t}\partial_{2}u_{0,1}\|_{L^{2}}\leq C(1+\nu t)^{-\frac{\sigma}{2s}}\left(\|\Lambda_{1}^{-\sigma}\partial_{2}u_{0,1}\|_{L^{2}}+\|\partial_{2}u_{0,1}\|_{L^{2}}\right)\leq\frac{C_{0}}{4}\varepsilon(1+\nu t)^{-\frac{\sigma}{2s}}.$ (4.25)
+$\|e^{\nu\Lambda_{1}^{2s}t}\partial_{2}u_{1,0}\|_{L^{2}}\leq C(1+\nu t)^{-\frac{\sigma}{2s}}\left(\|\Lambda_{1}^{-\sigma}\partial_{2}u_{1,0}\|_{L^{2}}+\|\partial_{2}u_{1,0}\|_{L^{2}}\right)\leq\frac{C_{0}}{4}\varepsilon(1+\nu t)^{-\frac{\sigma}{2s}}.$ (4.22)
 For the nonlinear term, due to $u_{2}$ decay more faster than $u_{1}$, we only consider the case contain $u\cdot\nabla u_{1}$.
@@
 $\|\partial_{1}\partial_{2}u_{1}(\tau)\|_{L^{2}}\leq\|\partial_{1}u_{1}\|_{L^{2}}^{1-\theta}\|\partial_{1}\partial_{2}^{\frac{1}{\theta}}u_{1}\|_{L^{2}}^{\theta}\leq C_{0}\varepsilon(1+t)^{-\frac{(\sigma+1)(1-\theta)}{2s}}.$ (4.23)
-Similarly, for $0<\theta<\min\{\frac{1}{9},\frac{1-2\sigma}{2\sigma}\}$,
+Similarly, we can also obtain a decay rate estimate of $\|\partial_{2}^{2}u_{1}(\tau)\|_{L^{2}}$ for $0<\theta<\min\{\frac{1}{9},\frac{1-2\sigma}{2\sigma}\}$ which will be used in the estimate for $III_{2}$,
 $\|\partial_{2}^{2}u_{1}(\tau)\|_{L^{2}}\leq C_{0}\varepsilon(1+t)^{-\frac{\sigma(1-\theta)}{2s}}.$ (4.24)
@@
 By Duhamel’s principle, we have
-$\|\partial_{1}u_{1}(t)\|_{L^{2}}\leq\|e^{\nu\Lambda_{1}^{2s}t}\partial_{1}u_{0,1}\|_{L^{2}}+C\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}(u\cdot\nabla u)(\tau)\|_{L^{2}}d\tau.\vskip-5.69046pt$
-For the linear term involving initial data $\partial_{1}u_{0,1}$, we can obtain the decay estimate through Lemma 2.7:
-$\|e^{\nu\Lambda_{1}^{2s}t}\partial_{1}u_{0,1}\|_{L^{2}}$ $\leq C(1+\nu t)^{-\frac{\sigma+1}{2s}}\left(\|\Lambda_{1}^{-\sigma}u_{0,1}\|_{L^{2}}+\|\partial_{1}u_{0,1}\|_{L^{2}}\right)$
+$\|\partial_{1}u_{1}(t)\|_{L^{2}}\leq\|e^{\nu\Lambda_{1}^{2s}t}\partial_{1}u_{1,0}\|_{L^{2}}+C\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}(u\cdot\nabla u)(\tau)\|_{L^{2}}d\tau.\vskip-5.69046pt$
+For the linear term involving initial data $\partial_{1}u_{1,0}$, we can obtain the decay estimate through Lemma 2.7:
+$\|e^{\nu\Lambda_{1}^{2s}t}\partial_{1}u_{1,0}\|_{L^{2}}$ $\leq C(1+\nu t)^{-\frac{\sigma+1}{2s}}\left(\|\Lambda_{1}^{-\sigma}u_{1,0}\|_{L^{2}}+\|\partial_{1}u_{1,0}\|_{L^{2}}\right)$
 $\leq\frac{C_{0}}{4}\varepsilon(1+\nu t)^{-\frac{\sigma+1}{2s}}.\vskip-5.69046pt$ (4.30)
@@
 Using Duhamel’s principle and taking the $L^{2}$-norm, we have
-$\|\partial_{1}u_{2}(t)\|_{L^{2}}\leq\|e^{\nu\Lambda_{1}^{2s}t}\partial_{1}u_{0,2}\|_{L^{2}}+\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{2}\Delta^{-1}\nabla\times(u\cdot\nabla u)(\tau)\|_{L^{2}}d\tau.$
+$\|\partial_{1}u_{2}(t)\|_{L^{2}}\leq\|e^{\nu\Lambda_{1}^{2s}t}\partial_{1}u_{2,0}\|_{L^{2}}+\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{2}\Delta^{-1}\nabla\times\mbox{div}\ (u\otimes u)(\tau)\|_{L^{2}}d\tau.$
 The linear term decays as
-$\|e^{\nu\Lambda_{1}^{2s}t}\partial_{1}u_{0,2}\|_{L^{2}}$ $\leq C(1+\nu t)^{-\frac{\sigma+2}{2s}}\big(\|\Lambda_{1}^{-\sigma}\psi_{0}\|_{L^{2}}+\|\partial_{1}u_{0,2}\|_{L^{2}}\big)$
+$\|e^{\nu\Lambda_{1}^{2s}t}\partial_{1}u_{2,0}\|_{L^{2}}$ $\leq C(1+\nu t)^{-\frac{\sigma+2}{2s}}\big(\|\Lambda_{1}^{-\sigma}\psi_{0}\|_{L^{2}}+\|\partial_{1}u_{2,0}\|_{L^{2}}\big)$
 $\leq C\varepsilon(1+\nu t)^{-\frac{\sigma+2}{2s}}\leq\frac{C_{0}}{4}\varepsilon(1+\nu t)^{-\frac{2\sigma}{s}}.$
 We devide the nonlinear term into two parts:
-$\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{2}\Delta^{-1}\nabla\times(u\cdot\nabla u)(\tau)\|_{L^{2}}d\tau$
+$\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{2}\Delta^{-1}\nabla\times\mbox{div}\ (u\otimes u)(\tau)\|_{L^{2}}d\tau$
 $\leq$ $\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{2}\Delta^{-1}\nabla\times\partial_{1}(u_{1}u)(\tau)\|_{L^{2}}d\tau+\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{2}\Delta^{-1}\nabla\times\partial_{2}(u_{2}u)(\tau)\|_{L^{2}}d\tau$
@@
 Combining (4.32), (4.35), (4.36) and (4.37), we can obtain the bootstrap conclusion (4.12) by taking $\varepsilon$ sufficiently small.
+4.7 Estimate for $\|\Lambda_{1}^{-\sigma}u\|_{L^{2}}$
+In this subsection, we show that, under the smallness assumption on $\|\Lambda_{1}^{-\sigma}u_{0}\|_{L^{2}}$, the decay rates of $u$ and its derivatives are sufficient to guarantee the propagation of $\|\Lambda_{1}^{-\sigma}u\|_{L^{2}}$.
+We establish the estimate for $\|\Lambda_{1}^{-\sigma}u\|_{L^{2}}$ through the following energy argument.
+Applying $\Lambda_{1}^{-\sigma}$ to the equation and taking the $L^{2}$ inner product with $\Lambda_{1}^{-\sigma}u$, we obtain
+$\frac{1}{2}\frac{d}{dt}\|\Lambda_{1}^{-\sigma}u\|_{L^{2}}^{2}+\nu\|\Lambda_{1}^{s-\sigma}u\|_{L^{2}}^{2}=-\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}(u\cdot\nabla u)\cdot\Lambda_{1}^{-\sigma}u\,dx.$ (4.38)
+The right-hand side of (4.38) can be decomposed into two parts and estimated as
+$\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}(u\cdot\nabla u)\cdot\Lambda_{1}^{-\sigma}u\,dx=$ $\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}(u_{1}\partial_{1}u)\cdot\Lambda_{1}^{-\sigma}u\,dx+\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}(u_{2}\partial_{2}u)\cdot\Lambda_{1}^{-\sigma}u\,dx$
+$\leq$ $C\left(\left\|\Lambda_{1}^{-\sigma}(u_{1}\partial_{1}u)\right\|_{L^{2}}+\left\|\Lambda_{1}^{-\sigma}(u_{2}\partial_{2}u)\right\|_{L^{2}}\right)\left\|\Lambda_{1}^{-\sigma}u\right\|_{L^{2}}$
+$:=$ $M_{0}.$
+Integrating in time on equation (4.38), we obtain
+$\|\Lambda_{1}^{-\sigma}u(t)\|_{L^{2}}^{2}+2\nu\int_{0}^{t}\|\Lambda_{1}^{s-\sigma}u(\tau)\|_{L^{2}}^{2}d\tau=\|\Lambda_{1}^{-\sigma}u_{0}\|_{L^{2}}^{2}+2\int_{0}^{t}M_{0}(\tau)d\tau.$ (4.39)
+By Lemma 2.6,
+$M_{0}$ $\leq C\Big(\left\|\partial_{2}u_{2}\right\|_{L^{2}}^{\frac{1}{2}}\left\|u_{2}\right\|_{L^{2}}^{\sigma}\left\|\partial_{1}u_{2}\right\|_{L^{2}}^{\frac{1}{2}-\sigma}\left\|\partial_{2}u\right\|_{L^{2}}$
+$\quad+\left\|\partial_{2}u_{1}\right\|_{L^{2}}^{\frac{1}{2}}\left\|u_{1}\right\|_{L^{2}}^{\sigma}\left\|\partial_{1}u_{1}\right\|_{L^{2}}^{\frac{1}{2}-\sigma}\left\|\partial_{1}u\right\|_{L^{2}}\Big)\left\|\Lambda_{1}^{-\sigma}u\right\|_{L^{2}}.$
+Using the divergence-free condition of $u$, we have
+$M_{0}$ $\leq C\Big(\left\|\partial_{1}u_{1}\right\|_{L^{2}}^{\frac{1}{2}}\left\|u_{2}\right\|_{L^{2}}^{\sigma}\left\|\partial_{1}u_{2}\right\|_{L^{2}}^{\frac{1}{2}-\sigma}\left\|\partial_{2}u\right\|_{L^{2}}$
+$\quad+\left\|\partial_{2}u_{1}\right\|_{L^{2}}^{\frac{1}{2}}\left\|u_{1}\right\|_{L^{2}}^{\sigma}\left\|\partial_{1}u_{1}\right\|_{L^{2}}^{\frac{1}{2}-\sigma}\left\|\partial_{1}u\right\|_{L^{2}}\Big)\left\|\Lambda_{1}^{-\sigma}u\right\|_{L^{2}}.$ (4.40)
+By substituting (4.2), (4.3), (4.4), (4.5) and (4.6) into (4.40), we deduce
+$M_{0}\leq CC_{0}^{3}\varepsilon^{3}\big((1+t)^{-\frac{-18\sigma^{2}+25\sigma+3}{12s}}+(1+t)^{-\frac{2\sigma+1}{2s}}\big).$
+It can be verified that
+$s<\tfrac{5}{12}+\sigma<\tfrac{-18\sigma^{2}+25\sigma+3}{12}.$
+As a consequence, we obtain
+$\int_{0}^{t}M_{0}(\tau)\,d\tau\leq CC_{0}^{3}\varepsilon^{3}.$
+Combining this bound with the assumption $\|\Lambda_{1}^{-\sigma}u_{0}\|_{L^{2}}\leq\varepsilon$, it follows from (4.39) that
+$\|\Lambda_{1}^{-\sigma}u(t)\|_{L^{2}}^{2}\leq\varepsilon^{2}+2CC_{0}^{3}\varepsilon^{3}.$
+For $\varepsilon>0$ sufficiently small so that $2CC_{0}^{3}\varepsilon\leq\tfrac{1}{2}$, we further deduce
+$\|\Lambda_{1}^{-\sigma}u(t)\|_{L^{2}}\leq\sqrt{\tfrac{3}{2}}\,\varepsilon\leq\tfrac{C_{0}}{2}\,\varepsilon.$
 5 Proof of Theorem 1.5
@@
 Assume that there exist constants $C_{0}>0$ and $T>0$ such that, for all $t\in[0,T)$, the following estimates hold:
-$\|\Lambda_{1}^{-\sigma}u(t)\|_{H^{k}}$ $\leq C_{0}\varepsilon,$ (5.1)
+$\|u(t)\|_{H^{k}}$ $\leq C_{0}\varepsilon,$ (5.1)
 $\|[x_{2}]^{\frac{3\gamma+4}{7}}u(t)\|_{L^{2}}$ $\leq C_{0}\varepsilon(1+t)^{-\sigma/2s},$ (5.2)
@@
 Under the assumptions of Theorem 1.5, we will prove the following estimates in the following subsections:
-$\|\Lambda_{1}^{-\sigma}u(t)\|_{H^{k}}$ $\leq 1/2C_{0}\varepsilon,$ (5.7)
+$\|u(t)\|_{H^{k}}$ $\leq 1/2C_{0}\varepsilon,$ (5.7)
 $\|[x_{2}]^{\frac{3\gamma+4}{7}}u(t)\|_{L^{2}}$ $\leq 1/2C_{0}\varepsilon(1+t)^{-\sigma/2s},$ (5.8)
@@
 Then using the standard bootstrap argument, we can conclude that $T=+\infty$ and derive the conclusions in Theorem 1.5.
-5.1 Estimate for $\|\Lambda_{1}^{-\sigma}u(t)\|_{H^{k}}$
-We splitting $\|\Lambda_{1}^{-\sigma}u(t)\|_{H^{k}}$ into two parts according to the equivalence ∥Λ_1^-σ u(t)∥_H^k ∼∥Λ_1^-σ u(t)∥_L^2 + ∑_m=1^2∥Λ_1^-σ ∂_m^k u∥_L^2, where the term $\sum_{m=1}^{2}\|\Lambda_{1}^{-\sigma}\partial_{m}^{k}u\|_{L^{2}}$ does not require coupling with decay estimates and has been completed in Subsection 4.2.
-To complete the estimation of $\|\Lambda_{1}^{-\sigma}u\|_{L^{2}}$, let us recall (4.14) and (4.15) in Subsection 4.1, that is
-$\|\Lambda_{1}^{-\sigma}u(t)\|_{L^{2}}^{2}+2\nu\int_{0}^{t}\|\Lambda_{1}^{s-\sigma}u(\tau)\|_{L^{2}}^{2}d\tau=\|\Lambda_{1}^{-\sigma}u_{0}\|_{L^{2}}^{2}+2\int_{0}^{t}M_{0}(\tau)d\tau,$
-where
-$M_{0}$ $\leq C\Big(\left\|\partial_{1}u_{1}\right\|_{L^{2}}^{\frac{1}{2}}\left\|u_{2}\right\|_{L^{2}}^{\sigma}\left\|\partial_{1}u_{2}\right\|_{L^{2}}^{\frac{1}{2}-\sigma}\left\|\partial_{2}u\right\|_{L^{2}}$
-$\quad+\left\|\partial_{2}u_{1}\right\|_{L^{2}}^{\frac{1}{2}}\left\|u_{1}\right\|_{L^{2}}^{\sigma}\left\|\partial_{1}u_{1}\right\|_{L^{2}}^{\frac{1}{2}-\sigma}\left\|\partial_{1}u\right\|_{L^{2}}\Big)\left\|\Lambda_{1}^{-\sigma}u\right\|_{L^{2}}.$ (5.13)
-By Lemma 2.9 and substituting (5.2), (5.3), (5.4), (5.5), (5.6) into (5.13), one can deduce
-$|M_{0}|\leq CC_{0}^{3}\varepsilon^{3}\big((1+t)^{-\frac{-2\sigma^{2}+5\sigma+2}{4s}}+(1+t)^{-\frac{2\sigma+1}{2s}}\big).$
-It can be checked that for $\frac{1}{3}<\sigma<\frac{1}{2}$, $s<\frac{1}{2}+\sigma<\frac{-2\sigma^{2}+5\sigma+2}{4}$, we have $\min\{\frac{-2\sigma^{2}+5\sigma+2}{4s},\frac{2\sigma+1}{2s}\}>1$ and
-$\int_{0}^{t}M_{0}(\tau)d\tau\leq CC_{0}^{3}\varepsilon^{3}.$
-Combining Lemma 2.9 and the initial condition (1.5), we derive from equation (4.14) that
-$\|\Lambda_{1}^{-\sigma}u(t)\|_{L^{2}}^{2}$ $\leq\varepsilon^{2}+2CC_{0}^{3}\varepsilon^{3}.$
-For $\varepsilon$ sufficiently small such that $2CC_{0}^{3}\varepsilon\leq\frac{1}{2}$, taking square roots of the above inequality, we obtain:
-$\|\Lambda_{1}^{-\sigma}u(t)\|_{L^{2}}\leq\sqrt{\frac{3}{2}}\varepsilon\leq\frac{C_{0}}{2}\varepsilon.$
+At the end, based on the decay rates of $u$ and its derivatives, we also prove the propagation of the negative horizontal derivative of $u$.
 5.1 Weighted estimate of the pressure
-The next step is to establish decay estimates for the solution.
+The energy estimate of $\|u\|_{H^{k}}$ is the same as the process in subsection 4.1.
+So the next step is to establish decay estimates for the solution.
 A major difficulty arises from the introduction of weighted norms: the Leray projection operator $\mathbb{P}$ is no longer directly bounded, since it involves Riesz transforms in a weighted setting.
@@
 $\|[x_{2}]^{\eta-1}\mathcal{E}p\|_{L^{2}}^{2}=\|[x_{2}]^{\eta-1}\mathcal{E}(-\Delta)^{-1}\mbox{div}\ \partial_{2}\int_{-\infty}^{x_{2}}(u\cdot\nabla u)(y_{2})dy_{2}\|_{L^{2}}^{2}.$
-Since $\frac{1}{2}<\eta<\frac{3}{2}$, we can use Lemma 2.13 and 2.10 to get
+Since $\frac{1}{2}<\eta<\frac{3}{2}$, we can use Lemma 2.13 and Lemma 2.10 to get
 $\|[x_{2}]^{\eta-1}\mathcal{E}p\|_{L^{2}}^{2}=$ $\|[x_{2}]^{\eta-1}\mathcal{E}(-\Delta)^{-1}\mbox{div}\ \partial_{2}\int_{-\infty}^{x_{2}}(u\cdot\nabla u)(y_{2})dy_{2}\|_{L^{2}}^{2}$
@@
 The linear term decays as
-$\|e^{\nu\Lambda_{1}^{2s}t}\partial_{1}[x_{2}]^{\gamma}\psi_{0}\|_{L^{2}}$ $\leq C(1+\nu t)^{-\frac{\sigma+1}{2s}}\big(\|\Lambda_{1}^{-\sigma}[x_{2}]^{\gamma}\psi_{0}\|_{L^{2}}+\|[x_{2}]^{\gamma}u_{0,2}\|_{L^{2}}\big)$
+$\|e^{\nu\Lambda_{1}^{2s}t}\partial_{1}[x_{2}]^{\gamma}\psi_{0}\|_{L^{2}}$ $\leq C(1+\nu t)^{-\frac{\sigma+1}{2s}}\big(\|\Lambda_{1}^{-\sigma}[x_{2}]^{\gamma}\psi_{0}\|_{L^{2}}+\|[x_{2}]^{\gamma}u_{2,0}\|_{L^{2}}\big)$
 $\leq C\varepsilon(1+\nu t)^{-\frac{\sigma+1}{2s}}\leq\frac{C_{0}}{4}\varepsilon(1+\nu t)^{-\frac{\sigma+1}{2s}}.$
@@
 $III^{w}_{2}\leq$ $C(C_{0}\varepsilon)^{2}\int_{0}^{t}(t-\tau)^{-\frac{1}{4s}}(1+\tau)^{-\frac{\sigma+1}{2s}}(1+\tau)^{-\frac{\sigma(1-\theta)}{2s}}d\tau.$ (5.39)
-Similarly, for $III_{3}^{w}$, by applying Lemma 2.7 and Hölder’s inequality, we have
+Similarly, for $III_{3}^{w}$, by applying Lemma 2.7 and Hölder inequality, we have
 $III_{3}^{w}$ $\leq C\int_{0}^{t}(t-\tau)^{-\frac{1}{4s}}\big\|[x_{2}]^{\frac{5\gamma+2}{7}}\|(\partial_{2}u_{1}\partial_{1}u_{2})(\tau)\|_{L^{1}_{x_{1}}}\big\|_{L^{2}_{x_{2}}}d\tau$
 $\leq C\int_{0}^{t}(t-\tau)^{-\frac{1}{4s}}\|[x_{2}]^{\frac{5\gamma+2}{7}}\partial_{2}u_{1}(\tau)\|_{L^{2}}\|\partial_{1}u_{2}(\tau)\|_{L^{2}_{x_{1}}L^{\infty}_{x_{2}}}d\tau.$
-By interpolation we have
+Applying interpolation to the second term in the integrand, we obtain
 $\|\partial_{1}u_{2}(\tau)\|_{L^{2}_{x_{1}}L^{\infty}_{x_{2}}}\leq$ $\|\partial_{1}u_{2}(\tau)\|_{L^{2}}^{\frac{1}{2}}\|\partial_{1}\partial_{2}u_{2}(\tau)\|_{L^{2}}^{\frac{1}{2}}$
@@
 $\int_{t-1}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{2}[x_{2}]^{\gamma}\partial_{1}\partial_{2}\Delta^{-1}(u_{1}u_{1})(\tau)\|_{L^{2}}d\tau\leq C\int_{t-1}^{t}\|\partial_{1}^{2}[x_{2}]^{\gamma}(u_{1}u_{1})(\tau)\|_{L^{2}}d\tau.$ (5.47)
-Applying Lemma 2.7 to (5.47) and substituting the resulting estimate together with (5.48) into (5.48), we obtain
+Applying Lemma 2.7 to (5.46) and substituting the resulting estimate in (5.47) into (5.45), we obtain
 $V_{1}^{w}\leq$ $C\int_{0}^{t-1}(t-\tau)^{-\frac{5}{4s}}\big\|[x_{2}]^{\gamma+1}\|\partial_{1}(u_{1}u_{1})\|_{L^{1}_{x_{1}}}\big\|_{L_{x_{2}}^{2}}d\tau+C\int_{t-1}^{t}\|[x_{2}]^{\gamma}\partial_{1}^{2}(u_{1}u_{1})\|_{L^{2}}d\tau$
@@
 Combining the estimates above, we can deduce the bootstrap conclusion (5.12) by taking $\varepsilon$ sufficiently small.
+5.7 Estimate for $\|\Lambda_{1}^{-\sigma}u\|_{L^{2}}$
+In this subsection, under the smallness assumption on $\|\Lambda_{1}^{-\sigma}u_{0}\|_{L^{2}}$, we prove that the decay rates of $u$ and its derivatives imply that $\|\Lambda_{1}^{-\sigma}u(t)\|_{L^{2}}$ is finite for all time.
+To complete the estimation of $\|\Lambda_{1}^{-\sigma}u\|_{L^{2}}$, let us recall (4.39) and (4.40) in Subsection 4.7, that is
+$\|\Lambda_{1}^{-\sigma}u(t)\|_{L^{2}}^{2}+2\nu\int_{0}^{t}\|\Lambda_{1}^{s-\sigma}u(\tau)\|_{L^{2}}^{2}d\tau=\|\Lambda_{1}^{-\sigma}u_{0}\|_{L^{2}}^{2}+2\int_{0}^{t}M_{0}(\tau)d\tau,$
+where
+$M_{0}$ $\leq C\Big(\left\|\partial_{1}u_{1}\right\|_{L^{2}}^{\frac{1}{2}}\left\|u_{2}\right\|_{L^{2}}^{\sigma}\left\|\partial_{1}u_{2}\right\|_{L^{2}}^{\frac{1}{2}-\sigma}\left\|\partial_{2}u\right\|_{L^{2}}$
+$\quad+\left\|\partial_{2}u_{1}\right\|_{L^{2}}^{\frac{1}{2}}\left\|u_{1}\right\|_{L^{2}}^{\sigma}\left\|\partial_{1}u_{1}\right\|_{L^{2}}^{\frac{1}{2}-\sigma}\left\|\partial_{1}u\right\|_{L^{2}}\Big)\left\|\Lambda_{1}^{-\sigma}u\right\|_{L^{2}}.$ (5.51)
+By Lemma 2.9 and substituting (5.2), (5.3), (5.4), (5.5), (5.6) into (5.51), one can deduce
+$|M_{0}|\leq CC_{0}^{3}\varepsilon^{3}\big((1+t)^{-\frac{-2\sigma^{2}+5\sigma+2}{4s}}+(1+t)^{-\frac{2\sigma+1}{2s}}\big).$
+It can be checked that for $\frac{1}{3}<\sigma<\frac{1}{2}$, $s<\frac{1}{2}+\sigma<\frac{-2\sigma^{2}+5\sigma+2}{4}$, we have $\min\{\frac{-2\sigma^{2}+5\sigma+2}{4s},\frac{2\sigma+1}{2s}\}>1$ and
+$\int_{0}^{t}M_{0}(\tau)d\tau\leq CC_{0}^{3}\varepsilon^{3}.$
+Combining Lemma 2.9 and the initial condition (1.5), we derive from equation (4.39) that
+$\|\Lambda_{1}^{-\sigma}u(t)\|_{L^{2}}^{2}$ $\leq\varepsilon^{2}+2CC_{0}^{3}\varepsilon^{3}.$
+For $\varepsilon$ sufficiently small such that $2CC_{0}^{3}\varepsilon\leq\frac{1}{2}$, taking square roots of the above inequality, we obtain:
+$\|\Lambda_{1}^{-\sigma}u(t)\|_{L^{2}}\leq\sqrt{\frac{3}{2}}\varepsilon\leq\frac{C_{0}}{2}\varepsilon.$
 Appendix A Proof of Lemmas in Section 2
@@
 Applying Duhamel’s principle to (4.18) and taking the $L^{2}$-norm, we can obtain
-$\|u_{2}(t)\|_{L^{2}}\leq\|e^{\nu\Lambda_{1}^{2s}t}u_{0,2}\|_{L^{2}}+\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}\Delta^{-1}\nabla\times(u\cdot\nabla u)(\tau)\|_{L^{2}}d\tau.$
+$\|u_{2}(t)\|_{L^{2}}\leq\|e^{\nu\Lambda_{1}^{2s}t}u_{2,0}\|_{L^{2}}+\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}\Delta^{-1}\nabla\times(u\cdot\nabla u)(\tau)\|_{L^{2}}d\tau.$
 The linear term decays as
-$\|e^{\nu\Lambda_{1}^{2s}t}u_{0,2}\|_{L^{2}}$ $\leq C(1+\nu t)^{-\frac{\sigma+1}{2s}}\big(\|\Lambda_{1}^{-\sigma}\psi_{0}\|_{L^{2}}+\|u_{0,2}\|_{L^{2}}\big)$
+$\|e^{\nu\Lambda_{1}^{2s}t}u_{2,0}\|_{L^{2}}$ $\leq C(1+\nu t)^{-\frac{\sigma+1}{2s}}\big(\|\Lambda_{1}^{-\sigma}\psi_{0}\|_{L^{2}}+\|u_{2,0}\|_{L^{2}}\big)$
 $\leq C\varepsilon(1+\nu t)^{-\frac{\sigma+1}{2s}}\leq\frac{C_{0}}{4}\varepsilon(1+\nu t)^{-\frac{3\sigma}{2s}}.$ (B.1)
@@
 $\int_{0}^{t-1}(t-\tau)^{-\frac{\sigma+1}{2s}}\|\Lambda^{-\frac{1}{2}-\sigma}\Lambda_{1}^{\frac{1}{2}}(u_{1}u)\|_{L^{2}}d\tau$
-$\leq$ $C\int_{0}^{t-1}(t-\tau)^{-\frac{\sigma+1}{2s}}\Big(\|\Lambda^{-\frac{1}{2}-\sigma}(\Lambda_{1}^{\frac{1}{2}}u_{1})u\|_{L^{2}}+\|\Lambda^{-\frac{1}{2}-\sigma}u_{1}(\Lambda_{1}^{\frac{1}{2}}u)\|_{L^{2}}\Big)d\tau.$
+$\leq$ $C\int_{0}^{t-1}(t-\tau)^{-\frac{\sigma+1}{2s}}\Big(\|\Lambda^{-\frac{1}{2}-\sigma}\Big[(\Lambda_{1}^{\frac{1}{2}}u_{1})u\Big]\|_{L^{2}}+\|\Lambda^{-\frac{1}{2}-\sigma}\Big[u_{1}(\Lambda_{1}^{\frac{1}{2}}u)\Big]\|_{L^{2}}\Big)d\tau.$
 Using Lemma 2.3, we deduce
-$\int_{0}^{t-1}(t-\tau)^{-\frac{\sigma+1}{2s}}\Big(\|\Lambda^{-\frac{1}{2}-\sigma}(\Lambda_{1}^{\frac{1}{2}}u_{1})u\|_{L^{2}}+\|\Lambda^{-\frac{1}{2}-\sigma}u_{1}(\Lambda_{1}^{\frac{1}{2}}u)\|_{L^{2}}\Big)d\tau$
+$\int_{0}^{t-1}(t-\tau)^{-\frac{\sigma+1}{2s}}\Big(\|\Lambda^{-\frac{1}{2}-\sigma}\Big[(\Lambda_{1}^{\frac{1}{2}}u_{1})u\Big]\|_{L^{2}}+\|\Lambda^{-\frac{1}{2}-\sigma}\Big[u_{1}(\Lambda_{1}^{\frac{1}{2}}u)\Big]\|_{L^{2}}\Big)d\tau$
 $\leq$ $C\int_{0}^{t-1}(t-\tau)^{-\frac{\sigma+1}{2s}}\|\Lambda^{\frac{1}{2}-\sigma}u\|_{L^{2}}\|\Lambda_{1}^{\frac{1}{2}}u\|_{L^{2}}d\tau.$